Information Sources: the Internet and Computer-Mediated Communication

John December (decemj@rpi.edu)

May 15, 1995

Release: 13 May 95; Release 3.95

Copyright \begin{matrix}c \\ ◯ \end{matrix}1994 John December (decemj@rpi.edu). You may use this document for any personal or educational purpose. For-profit distribution requires my permission.

DISCLAIMER: Provided “as is” without expressed or implied warranty.

PURPOSE: to list information describing the Internet and computer-mediated communication technologies, applications, culture, discussion forums, and bibliographies. Areas of interest include include the technical, rhetorical, social, cognitive, and psychological aspects of networked communication and information.

AUDIENCE: those getting started in understanding the Internet and CMC; for those experienced, it collects sources of information.

ASSUMPTION: you have access to and know how to use finger, ftp, gopher, telnet, email, World Wide Web, WAIS, or Usenet newsgroups.

NOTES:

1. Respect access privileges. (See Net Etiquette Guide, below.)

2. This information changes quite a bit; additions/comments welcome.

3. Use Archie, Veronica, or a Web spider to find a file if it is not at the site given here.

4. Learn to use Gopher, Veronica, WAIS, WWW, to find more information.

Contents

1 ABOUT this information 4 1.1 Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 INTERNET 5

2.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.1 Motivation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.2 Overviews . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.3 Facts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.4 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.2 Collections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2.1 NICs = Network Information Centers . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2.2 Series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2.3 Training . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3.1 Collections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3.2 Resources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3.3 Exploring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3.4 Courses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3.5 Tutorials for specific applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.4 Navigating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.4.1 Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.4.2 Guides . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.4.3 Collections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.4.4 Specialized guides or collections of general interest . . . . . . . . . . . . . . . . . . . . 11 2.4.5 Navigators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

CONTENTS 2

2.5 Searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5.1 New or Noteworthy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5.2 Resource Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.3 Subjects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.4 Keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.5.5 Spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.5.6 People . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

2.6 Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.7 Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

3 APPLICATIONS 20

3.1 Commerce . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.1.1 Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.1.2 Marketplaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

3.2 Communication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.2.1 Individual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.2.2 Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.2.3 Organizational . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.2.4 Mass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.2.5 Societal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.2.6 Surveillance / remote-control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.2.7 Scientific . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.3 Education . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.4 Entertainment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.5 Government . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

3.5.1 Initiatives for information infrastructures . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.5.2 Information supplied by governments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.5.3 Policy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

3.6 Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.6.1 Dissemination and Retrieval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.6.2 Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

3.7 Multiple . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.8 Scholarship . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

4 TECHNOLOGY 36 4.1 Computing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.2 Developing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.3 Human Interaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.4 Multimedia . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

4.4.1 Audio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.4.2 Graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.4.3 Multi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.4.4 Hypermedia . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

4.5 Virtual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.6 Networks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

4.6.1 Access and connectivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.6.2 Administration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.6.3 Networking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.6.4 Security . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.6.5 Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.6.6 Maps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.7 Telecommunications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

5 CULTURE 45 5.1 Art . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.2 Community . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

5.2.1 Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.2.2 Virtual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

5.3 Language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.4 People . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

5.4.1 Aspects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.4.2 Lists – directories, home pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

5.5 Society . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

CONTENTS 3

6 FORUMS 50 6.1 Academic resources in communication and related fields . . . . . . . . . . . . . . . . . . . . . 50 6.2 Meetings and Activities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 6.3 Discussion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 6.4 Periodicals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.5 Usenet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

7 ORGANIZATIONS 60 7.1 Academic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.2 Research . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.3 Commercial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

7.3.1 Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 7.3.2 Providers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.3.3 Consortia . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.3.4 Developers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.3.5 Products . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.3.6 Telecom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

7.4 Internet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.5 Network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.6 Non-profit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.7 Standards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

8 BIBLIOGRAPHY 70 8.1 Online . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 8.2 Special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 8.3 Books . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

8.3.1 Bibliographies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.3.2 Internet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.3.3 Networks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 8.3.4 CMC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

Information about updates of this document:

• John December: contact about updates, questions, comments, corrections (mailto:decemj@rpi.edu).

1 ABOUT THIS INFORMATION 4

1 ABOUT this information

1.1 Notes

• README: describes the Internet and Computer-Mediated Communication list (ftp://ftp.rpi.edu/pub/communications/internet-cmc.readme).

• Challenges/Web Info: Challenges for Web Information Providers, by John December; an article describing the development of the internet-cmc information and issues for Web information providers (http://sunsite.unc.edu/cmc/mag/1994/oct/webip.html).

• Segmented top: describes purpose, notes, how to use the segmented hypertext format of this information (http://www.rpi.edu/Internet/Guides/decemj/icmc/top.html).

• Tool Tips: describes tips for using access methods described in this list (ftp://ftp.rpi.edu/pub/communications/internet-tools.use).

• Use Tips: describes tips for using the Internet and Computer-Mediated Communication list (ftp://ftp.rpi.edu/pub/communications/internet-cmc.use).

1.2 Formats

These are the other formats for this information. Note that the files at the ftp site (ftp.rpi.edu) will not be as current as those on the web server.

• Database: the database from which I generate this information, warning–this is a LARGE file (ftp://ftp.rpi.edu/pub/communications/internet-cmc.dat).

• DVI: Device-independent file (ftp://ftp.rpi.edu/pub/communications/internet-cmc.dvi).

• HTML 1 BIG (!) page: HTML version of this information on one big page (http://www.rpi.edu/Internet/Guides/decemj/icmc/internet-cmc.html).

• HTML segmented: HTML version of this information in a web of pages divided by section, sub-section, and sub-sub-section (http://www.rpi.edu/Internet/Guides/decemj/icmc/top.html).

• LaTex: LaTex source file (ftp://ftp.rpi.edu/pub/communications/internet-cmc.tex).

• Postscript: postscript version of this information (ftp://ftp.rpi.edu/pub/communications/internet-cmc.ps.Z).

• Text Wide: text version of this information ¿80 columns wide (ftp://ftp.rpi.edu/pub/communications/internet-cmc).

• Text 80: text version of this information, 80 columns wide (ftp://ftp.rpi.edu/pub/communications/internet-cmc.txt).

2 INTERNET 5

2 INTERNET

2.1 Introduction

2.1.1 Motivation

• Gold in Networks!: a description of gold nuggets in the network, by J. Martin (ftp://nic.merit.edu/documents/fyi/fyi 10.txt).

2.1.2 Overviews

• GlobalCenter Tour: A tour of the Internet from Global Village Communication (http://www.globalcenter.net/gcweb/tour.html).

• Hitchikers Guide: describes the Internet (circa September 1989), by Ed Krol (ftp://nic.merit.edu/documents/rfc/rfc1118.txt).

• Surfing the Internet: a narrative of what the Internet has to offer, by Jean Armour Polly (ftp://nysernet.org/pub/resources/guides/surfing.2.0.3.txt).

• What is the Internet?: by Krol and Hoffman (ftp://nic.merit.edu/documents/fyi/fyi 20.txt).

• Xerox Video: An Overview of the Internet and World-Wide Web, by Xerox Palo Alto Research Center (http://pubweb.parc.xerox.com/hypertext/wwwvideo/wwwvideo.html).

2.1.3 Facts

• Internet Index: The Internet Index, a list of interesting facts about the Internet [Inspired by ‘Harper’s Index’], by Win Treese (http://www.openmarket.com/info/internet-index/current.html).

• New User’s Questions: answers questions commonly asked by new users, by Malkin and Marine (ftp://nic.merit.edu/documents/fyi/fyi 04.txt).

2.1.4 History

• Internet birthday: happy 25th, Internet (http://www.amdahl.com/internet/events/inet25.html).

• Internet future: proposals for the future of the Internet (mailto:Internet-drafts@nri.reston.va.us Body: help).

• Internet History/ISOC: Internet Society’s collection of Internet history (gopher://gopher.isoc.org/11/internet/history).

• Internet History: Master’s Thesis, Henry Edward Hardy (ftp://umcc.umich.edu/pub/users/seraphim/doc/nethist8.txt).

• Internet timeline: An Internet timeline highlighting some of the key events which helped shape the Internet as we know today, by Robert H’obbes’ Zakon (mailto:timeline@hobbes.mitre.org).

• Internet Timeline: events in the history of the Internet by Robert H’obbes’ Zakon (http://tig.com/IBC/Timeline.html).

• Netizen Anthology: The Netizens and the Wonderful World of the Net–an anthology, Ronda Hauben, Michael Hauben (http://www.columbia.edu/∼hauben/project book.html).

2.2 Collections

2.2.1 NICs = Network Information Centers

• Acceptable Use: directory of acceptable use policies for many networks (ftp://nic.merit.edu/acceptable.use.policies/).

• Domain Registration: (http://www.yahoo.com/Computers/Internet/Domain Registration/).

• DDN NIC: Network Information Center/Defense Data Network (ftp://nic.ddn.mil).

• InterNIC Telnet: Directory and Database Services Telnet Interface (telnet://guest@ds.internic.net).

• InterNIC Ftp: ftp archives for InterNIC (ftp://ds.internic.net/pub/InterNIC-info/internic.info).

• InternNIC email: Mail service to the InterNIC (mailto:mailserv@internic.net Body: help).

• InterNIC gopher: information server for Internet Network Information Center (gopher://rs.internic.net).

• InterNIC web: directory database (AT+T) and registration (Network Solutions) services (http://www.internic.net).

• InterNIC InfoSource gopher: (gopher://is.internic.net/11/infoguide/).

• InterNIC DS Web: Directory and Database Services Home Page (http://ds.internic.net).

• InterNIC Database Services: (ftp://ds.internic.net).

• InterNIC Reg Services: (ftp://rs.internic.net/netinfo/).

• Merit: network information center (ftp://nic.merit.edu).

• Merit Gopher: network information center (gopher://nic.merit.edu).

• SuraNet NIC: Southeastern Universities Research Association, Inc. Network Information Center (http://www.sura.net/index.html).

• SWITCHinfo: Swiss Network (ftp://nic.switch.ch/info service/).

2.2.2 Series

• EIN Web: Internet Engineering Notes (http://www.cis.ohio-state.edu/hypertext/information/ien.html).

• FYI: For Your Information-sub-series of RFCs about topics which relate to the Internet (ftp://nic.merit.edu/documents/fyi/).

• FYIs in HTML: (http://www.ftp.com/techsup/fyi/fyi-index.html).

• Internet Docs: Internet Documentation (RFC’s, FYI’s, etc.) from InterNIC (http://ds.internic.net/ds/dspg0intdoc.html).

• Internet Drafts: draft documents to be submitted ultimately to the Internet Activities Board to be as RFCs (ftp://nic.merit.edu/documents/internet-drafts/).

• Internet Drafts Search: (http://web.nexor.co.uk/idindex).

• Internet Handbook: a list of RFCs by subject and category (ftp://sri.com/netinfo/internet-technology-handbook-contents).

• STD: Internet Standards, sub-series of notes within the RFC series which document Internet standards (ftp://nic.merit.edu/documents/std/).

• RFC repositories: repositories where RFCs are located (ftp://isi.edu/in-notes/rfc-retrieval.txt).

• RFC: Request for Comments-a document series which describes the Internet suite of protocols and related experiments (ftp://nic.merit.edu/documents/rfc/).

• RFC Web: Internet Request For Comments (RFC) in html (http://www.cis.ohio-state.edu/hypertext/information/rfc.html).

• RFCs via mail: Request For Comments–documents about various issues for discussion, covering a broad range of networking issues. (mailto:rfc-info@isi.edu Body: help: ways to get rfcs).

2.3 Training

2.3.1 Collections

• Beginner’s Guides-Yahoo: Computers-Internet-Beginner’s Guides entry from Yahoo (http://www.yahoo.com/Computers/Internet/Beginner s Guides/).

• CBL: Computer-Based Learning Unit, The University of Leeds, United Kingdom (http://cbl.leeds.ac.uk).

• DELTA: Distributed ELectronic Telecommunications Archive-teaching and learning about business telecommunications and data communications (http://gozer.idbsu.edu/business/nethome.html).

• HelpNet Archives: reference base for new users to the Internet (ftp://ftp.temple.edu/pub/info/help-net).

• Introducing the Internet: Merit’s Directory of miscellaneous introductory Internet information (ftp://nic.merit.edu/introducing.the.internet/).

• IETF/TERENA: Internet Engineering Task Force (IETF) Training Materials Catalogue (http://coolabah.itd.adelaide.edu.au/TrainMat/catalogue.html).

• ITTI: Information Technology Training Initiative, a United Kingdom-wide initiative to provide training materials about or using Information Technology (http://info.mcc.ac.uk/CGU/ITTI/ITTI.html).

• Matrix-WELL: information about cyberspace networks (gopher://gopher.well.sf.ca.us/11/matrix).

• MOREnet: training materials from Missouri Research and Education Network (MOREnet) (ftp://ftp.more.net/pub/nic/training).

• NETTRAIN archive: NETTRAIN (networking training) discussion list archives (ftp://ubvm.cc.buffalo.edu/nettrain/).

• Nets and DB-MSU: Network and Database Resources (CWIS, gopher, electronic info.) (gopher://burrow.cl.msu.edu).

• NOCALL: Northern California Association of Law Libraries (NOCALL) and the Southern California Association of Law Libraries (SCALL) to encourage use of the Internet and provide training assistance to our members (ftp://ftp.netcom.com/pub/loftus/buddies/home.html).

• Schneider Bib: A bibliography of Internet Training Materials, by Karen Schneider (ftp://alexia.lis.uiuc.edu/pub/training.bib).

• Start Web: Where to Start for New Internet Users (in HTML), by James Milles (http://www.law.cornell.edu/test/newusers.html).

• Sunsite-UNC: Univ of North Carolina’s collection (ftp://sunsite.unc.edu/pub/docs/about-the-net/).

• Trainmat: Training Materials Gopher (gopher://trainmat.ncl.ac.uk).

• Trainpack: repository of training and training-related materials (ftp://ftp.ncl.ac.uk/pub/network-training/).

• Venturing-Newfoundland: Memorial University of Newfoundland Venturing into the Internet collection (gopher://cwis.ucs.mun.ca/11/Venturing%20into%20the%20Internet).

• Wash Coll of Law: Net Training Materials, from Washington College of Law (http://sray.wcl.american.edu/htm/training.htm).

• Web Training: University of Toronto Instructional and Research Computing Group (UTIRC) World Wide Web home page, information about HTTP, HTML, and networked multimedia training and technology (http://www.utirc.utoronto.ca/home.html).

• Wheeler: Internet Training handouts, by Bill Wheeler (ftp://s850.mwc.edu/nettrain/).

• Yale: Internet Help (gopher://yaleinfo.yale.edu).

2.3.2 Resources

• Dir of Trainers/consultants: from the Internet Business Journal (gopher://gopher.fonorola.net/00/Internet%20Business%20Journal/Directory of Trainers and Consultants).

• Montsarrat: training materials (basic and serious Internet)

(ftp://wilma.cs.brown.edu/pub/internet course.tar.Z).

• Net workbook: by Martin Raish (mailto:listserv@bingvmb.bitnet Body: GET NETKNOW NEOPHYTE BI-L).

• Newcastle/Tyne: Network Training Materials and Related Documentation (ftp://mailbase.ac.uk/pub/lists/itti-networks/files/tms-list.txt).

• Resources collection: guides, lists, sites, resources of interest and use to Internet Trainers, by Neil Enns (http://www.brandonu.ca/∼ennsnr/Resources/).

• Start: Where to Start for New Internet Users, by James Milles (ftp://sluaxa.slu.edu/pub/millesjg/newusers.faq).

• VAX/VMS Networking: Beginner’s Guide to Networking on the VAX/VMS, by Lucia Ruedenberg (ftp://ftp.temple.edu/pub/info/help-net/vms-mail.guide).

2.3.3 Exploring

• Beginners/Suarez: The Beginner’s Guide to the Internet for DOS and Windows, by Patrick J. Suarez (ftp://oak.oakland.edu/SimTel/msdos/info/bgi20.zip).

• DOS Internet Kit: public domain programs that enable Ethernet or serially-connected PCs to access Internet services, by Dean Pentcheff (ftp://tbone.biol.scarolina.edu/pub/kit/).

• DOS Internet Kit Web: public domain programs that enable Ethernet or serially-connected PCs to access Internet services, packaged as a kit, by Dean Pentcheff (http://tbone.biol.scarolina.edu/∼dean2/kit/kit.html).

• Internet Cruise: A Cruise of the Internet, computer-based tutorial for Internet navigators, DOS or Mac formats (ftp://nic.merit.edu/resources/).

• Internet Explorer’s Kit: a hypertext system about the Internet for DOS, by Ernest Perez (ftp://sunsite.unc.edu/pub/docs/about-the-net/libsoft/explorer.doc).

• Internet Hunt: a game for learning about Internet resources (ftp://ftp.cni.org/pub/net-guides/i-hunt/).

• Internet Hunt Gopher: Hunt Questions, Results, and Comments (gopher://gopher.cic.net/11/hunt).

• Internet Navigating: Navigating the Internet Workshop List, by Richard J. Smith (ftp://ubvm.cc.buffalo.edu/navigate/).

2.3.4 Courses

• CMC: a graduate course in CMC with some network training components (http://www.rpi.edu/Internet/Guides/decemj/course/cmc.html).

• Compass in Cyberspace: Internet Training from John S. Makulowich (http://www.clark.net/pub/journalism/brochure.html).

• Courses/info tech: courses related to social implications of information technology, list by Meng Weng Wong (http://www.seas.upenn.edu/∼mengwong/netsurf/otherclasses.html).

• Gopherin: a course for cyberspace, by Jim Gerland and Rich Smith (via ftp)

(ftp://ubvm.cc.buffalo.edu/gophern/).

• Internet Courses: Course Descriptions and Examples, by Neil Enns (http://www.brandonu.ca/∼ennsnr/Resources/courses.html).

• Internet Resource: Internet Resource Discovery, Organization, and Design, by Lou Rosenfeld and Dr. Joe Janes, School of Information and Library Studies, University of Michigan (http://http2.sils.umich.edu/∼lou/60694.html).

• Intro/Internet: one of several prototype classes and texts sponsored through the Globewide Network Academy (http://uu-gna.mit.edu:8001/uu-gna/text/internet/index.html).

• ISS 101: Internet survival skills, a course from Dr. Kenneth Hensarling (http://kawika.hcc.hawaii.edu/iss101/101mods.html).

• Gopherin course: Jim Gerland and Rich Smith’s course about gopher (via gopher) (gopher://wealaka.okgeosurvey1.gov/11/K12/GOPHERN).

• LeJeune Course: course materials and files, by Urban A. LeJeune (ftp://pilot.njin.net/pub/Internet-course/).

• Navigating: Navigating the Internet, workshop for teaching the Internet via email, by Richard J. Smith (gopher://jake.esu.edu/11/Help/net stuff/training).

• Navigating: Navigating the Internet workshop (mailto:listserv@ubvm.cc.buffalo.edu Body: subscribe navigate YOUR NAME).

• Navigating: an interactive workshop, by Richard J. Smith (ftp://ftp.sura.net/pub/nic/training/).

• Netsurf course: by Meng Weng Wong, covers social implications of information technologies (http://www.seas.upenn.edu/∼mengwong/netsurf/).

• NITEC: NYSERNet (New York Education Education and Research Network) Internet Training and Education Center (http://nysernet.org/nitec.info/).

• Milles: An Introduction to Using the Internet at St. Louis University School of Law (ftp://sluaxa.slu.edu/pub/millesjg/interlaw.txt).

• Mining: Mining the Internet course, Central Michigan University’s Computer Science 196 ($) (mailto:34mjkeq@cmuvm.csv.cmich.edu).

• Surfing: a course given at Florida State in Internet use (http://www.cs.fsu.edu/surfing.html).

2.3.5 Tutorials for specific applications

• Ackermann tutorial: Internet Services and Resources for Computer Scientists, by Ernest Ackermann, Department of Computer Science, Mary Washington College, Fredericksburg, VA (ftp://s850.mwc.edu/pub/tutorial).

• Competencies: Competencies for Electronic Information Services, by John Corbin (mailto:listserv@uhupvm1.uh.edu Body: GET CORBIN PRV4N6).

• Discussion help: The art of getting help–guidelines for seeking help in discussion groups (mailto:rre-request@weber.ucsd.edu Subject: archive send courtesy).

• HTML Developer’s: WWW + HTML Developer’s JumpStation, maintained by SingNet and hosted by OneWorld Information Services (http://oneworld.wa.com/htmldev/devpage/dev-page.html).

• HTML FAQ: hypertext markup language (http://www.umcc.umich.edu/∼ec/www/html faq.html).

• HTML info/CERN: HyperText Markup Language (HTML), Working and Background Materials, from CERN (http://www.w3.org/hypertext/WWW/MarkUp/MarkUp.html).

• Mosaic Web Index: An index to the National Center for Supercomputing Applications (NCSA) Mosaic online documentation and tutorials on the NCSA Web server (http://www.ncsa.uiuc.edu/SDG/Software/Mosaic/Docs/web-index.html).

• NIR/CMC Tools/December: a list summarizing Internet tools and computer-mediated communication forums; includes pointers to guides or information about each tool, by John December (http://www.rpi.edu/Internet/Guides/decemj/itools/top.html).

• Tutorial gateway: a filter for a CGI-compliant HTTP server that makes it slightly easier to develop tutorial style questions for Web users (http://www.civeng.carleton.ca/∼nholtz/tut/doc/doc.html).

• Unix Help: Nova University’s Unixhelp for users (http://alpha.acast.nova.edu/UNIXhelp/TOP .html).

• Unix Man Pages: HyperText interface to UNIX manpages, by Michael Fisk (http://nmt.edu/bin/man).

• Unix tutorials: a collection of Unix and computer information, from East Stroudsburg University (gopher://jake.esu.edu/11/Help/Tutorials).

• URL: Curling Up to Universal Resource Locators, by Eric S. Theise (gopher://gopher.well.sf.ca.us/00/matrix/internet/curling.up.02).

• URL guide/NCSA: A Beginner’s Guide to URLs, by Marc Andreessen (http://www.ncsa.uiuc.edu/demoweb/url-primer.html).

• Usenet and trn: An Introduction to Usenet News and the trn Newsreader, by Jon Bell (ftp://cs1.presby.edu/pub/trn-intro/).

• WWW Weavers: a collection of resources related to developing information on the World Wide Web, by Chris Beaumont (http://www.nas.nasa.gov/NAS/WebWeavers/).

• WWW talk/Maimone: Welcome to the World Wide Web! Tutorial slides by Mark Maimone, Carnegie Mellon Computer Science (http://www.cs.cmu.edu:8001/afs/cs/usr/mwm/www/tutorial/index.html).

• WWW talk/Musciano: An Introduction to the World Wide Web Chuck Musciano, Corporate Information Management, Harris Corporation, September 2, 1994 (http://melmac.corp.harris.com/www intro/).

• WWW talk/Wallach: The World Wide Web, Everything you’ve always wanted to know, but were afraid to ask, by Dan Wallach (http://www.cs.princeton.edu/grad/Dan Wallach/www-talk/talk0.html).

• WWW talk/Berners-Lee: WWW Seminar, Transparency track, by Tim Berners-Lee , Robert Cailliau, CERN (http://www.w3.org/hypertext/WWW/Talks/General/Transparencies.html).

• X Tutorial: X Window System Tutorial, by David Marshall (http://www.cm.cf.ac.uk/Dave/X lecture/X lecture.html).

2.4 Navigating

2.4.1 Tools

• GNN’s Tools: a list of tools and supporting documents, from Global Network Navigator (http://nearnet.gnn.com/gnn/helpdesk/tools/index.html).

• Internet Browsers: Sources for Internet Browsers and Client Software (http://life.anu.edu.au/links/syslib.html).

• Internet Tools/EARN HTML: European Academic Research Network Association Guide to Network Resources Tools in HTML (http://www.earn.net/gnrt/notice.html).

• Internet Tools NIR: A status report on networked information retrieval tools and groups, by Joint IETF/RARE/CNI Networked Information Retrieval Working Group (ftp://mailbase.ac.uk/pub/lists/nir/files/nir.status.report).

• Internet Tools/December HTML: HTML (segmented) version of a list summarizing Internet tools and computer-mediated communication forums, by John December (http://www.rpi.edu/Internet/Guides/decemj/itools/top.html).

• Internet Tools/December (text): a list summarizing Internet tools and computer-mediated communication forums, by John December (ftp://ftp.rpi.edu/pub/communications/internet-tools.readme).

• Internet Systems UNITE: a list of Internet tools and systems, by User Network Interface To Everything (ftp://mailbase.ac.uk/pub/lists/unite/files/systems-list.txt).

2.4.2 Guides

• AARNet Guide: Australian Network Sites and Resources Guide (ftp://aarnet.edu.au/pub/resource-guide/).

• AARNet User Guide: Australian Network User’s Guide (ftp://aarnet.edu.au/pub/user-guide/).

• Argus Guide: The Internet Tools and Resources Guide (http://argus-inc.com/Guide.html).

• De Presno Guide Gopher: full text of the Online World book by Odd de Presno (gopher://wuecon.wustl.edu:10672/11/online).

• De Presno Guide via FTP: a shareware book, ‘The Online World’ (long)

(ftp://ftp.eunet.no/pub/text/).

• De Preson Guide via WWW: (http://login.eunet.no/∼presno/index.html).

• De Presno Guide Email: (mailto:LISTSERV@vm1.nodak.edu Body: get tow where).

• EFF’s Guide: (formerly Big Dummy’s Guide) written by Adam Gaffin for the Electronic Frontier Foundation (ftp://ftp.eff.org/pub/Net info/EFF Net Guide/).

• EFF Guide Web: (formerly Big Dummy’s) Guide in HTML, from the Electronic Frontier Foundation (EFF) (http://www.eff.org/papers/eegtti/eegttitop.html).

• EFF Guide Search: search the (formerly Big Dummy’s) Guide (http://www.germany.eu.net/cgi-bin/eegfind.pl).

• Email 101: describes how to use email as well as other Internet features, by John Goodwin (ftp://mrcnext.cso.uiuc.edu/etext/etext93/email025.txt).

• Email/access: Accessing The Internet By E-Mail, by ‘Doctor Bob’ Rankin (mailto:mail-server@rtfm.mit.edu Body: send usenet/news.answers/internet-services/access-via-email).

• Email/Start: How to get initial internet info via e-mail, by David Rosen (http://www.scs.unr.edu/%7Ecbmr/net/start.html).

• Entering/WWW: Kevin Hughes’s guide to the web (http://www.eit.com/web/www.guide/).

• Entering/WWW ftp: ftp site for Kevin Hughes’s web guide (ftp://ftp.eit.com/pub/web.guide/).

• HWGUIDE: A hypertext version of Noonan’s Guide to Internet/Bitnet (ftp://sunsite.unc.edu/pub/docs/about-the-net/libsoft/hwguide.txt).

• InfoPop: a WINHELP (hypertext) guide to the Internet, CompuServe, BBS systems, and more, by Wally Grotophorst (ftp://ftp.gmu.edu/library/).

• Internet Companion(parts): A Beginner’s Guide to Global Networking by Tracey LaQuey (ftp://ftp.std.com/OBS/The.Internet.Companion/).

• Internet Guide: A guide to Internet/Bitnet, by Dana Noonan (ftp://sunsite.unc.edu/pub/docs/about-the-net/libsoft/guide1.txt).

• Internet Tour: A Short, Semi-Guided Tour of the Internet, from Kapor Enterprises (http://www.kei.com/internet-tour.html).

• Internet Web Text: an index to the Internet’s resources; links to orientation, guides, reference materials, browsing and exploring tools, subject and word-oriented searching tools, and information about connecting with people, by John December (http://www.rpi.edu/Internet/Guides/decemj/text.html).

• La toile d’araignee: French translation with European-oriented resources, by Francois Charoy, of John December’s Internet Web Text (http://www.loria.fr/∼charoy/ToileInternet/text.html).

• Library of Congress: a directory of Internet User Guides (ftp://ftp.loc.gov/pub/iug/).

• Meng’s: Holistic Guide to the Internet, by Meng Weng Wong (http://ccat.sas.upenn.edu/mengwong/guide.html).

• Neophyte: Network Knowledge for the Neophyte, Stuff You Need to Know in Order to Navigate the Electronic Village, by Martin Raish (ftp://hydra.uwo.ca/pub/libsoft/network knowledge for the neophyte.txt).

• NSF Resource Guide: National Science Foundation Guide to Internet Resources (ftp://ds.internic.net/resource-guide/overview).

• NWNet Internet Guide: NorthWestNet’s User Service Internet Resource Guide (ftp://ftphost.nwnet.net/user-docs/nusirg/README.nusirg).

• NYSERNet Internet Guide: New York State Education and Research Network Guide to the Internet (ftp://nysernet.org/pub/guides/Guide.V.2.2.text).

• SLAC: Introduction to the Internet–a narrative about the Internet from SLAC (Stanford Linear Accelerator Center), by Joan Winters (http://slacvm.slac.stanford.edu/FIND/internet.html).

• SURANet Internet Guide: updated as infoguide.MM-YR.txt (ftp://ftp.sura.net/pub/nic/).

• SURFnet Guide: SURFnet is the Dutch academic network (ftp://ftp.nic.surfnet.nl/surfnet/user-support/docs/training/).

• WWW/Mosaic: User’s Guide, by Craig I. Schlenoff (http://elib.cme.nist.gov/fasd/pubs/schlenoff94.html).

• Zen/Art of Internet: Zen and the Art of the Internet (1st edition), by Brendan Kehoe (ftp://ftp.csn.org/pub/net/zen/).

• Zen Web: Zen and the Art of the Internet (1st edition), by Brendan Kehoe, in HTML (http://sundance.cso.uiuc.edu/Publications/Other/Zen/zen-1.0 toc.html).

2.4.3 Collections

• GNN’s Internet Helpdesk: from Global Network Navigator’s (O’Reilly)

(http://www.ora.com/gnn/helpdesk/index.html).

• Gopher Jewels: Internet Cyberspace related gophers (http://galaxy.einet.net/GJ/internet.html).

• Hypermedia/Internet: Hypermedia and the Internet, by David Green (http://life.anu.edu.au/education/hypermedia.html).

• Info Deli: a collection of Net information, by Peter Kaminski (ftp://ftp.netcom.com/pub/info-deli/bookmark.html).

• Internet-Yahoo: pointers to the subject classification Computers-Internet (http://www.yahoo.com/Computers/Internet/).

• NetCruiser: NETCOM On-Line Communications Services, Inc.’s online guide (http://www.netcom.com/glee/cruiser.html).

• NOSC Page: list of Internet-related sources, by Richard Bocker (http://white.nosc.mil/internet.html).

• Naval Research Lab: NRL Network Research Navigator, strong on technical aspects of networking (http://netlab.itd.nrl.navy.mil).

• The Net: descriptions of how to access information through Mosaic/Web, by Robert Thau (http://www.ai.mit.edu/the-net/overview.html).

• Tractatus CyberNauticus: information about the Internet and the Web, by Denys Duchier (http://kaml1.csi.uottawa.ca:3000/tractatus.html).

• Yale: General Overview of Worldwide Internet Resources (http://www.cs.yale.edu/HTML/WorldWideWebTop.html).

• U of TX: Internet Information Resources, from the University of Texas (http://fiat.gslis.utexas.edu/internet/internet.html).

2.4.4 Specialized guides or collections of general interest

• Dern’s Info: Daniel Dern’s Internet Info, News and Views (gopher://gopher.internet.com:2200/11/).

• Agricultural Guide: Not Just Cows, by Wilfred Drew (gopher://snymorva.cs.snymor.edu/hhGOPHER ROOT1:[LIBRARY-DOCS.HTML]).

• DDN New User Guide: Defense Data Network guide for new users (ftp://nic.ddn.mil/netinfo/nug.doc).

• Electric Mystics Guide: A Complete Directory of Networked Electronic Documents, Online Conferences, Serials, Software, and Archives Relevant to Religious Studies, by Michael Strangelove (ftp://panda1.uottawa.ca/pub/religion/).

• FidoNet Guide: The Big Dummy’s Guide to Fidonet, by Michael Schuyler (gopher://digital.cosn.org/11/Networking%20Information/Reference/The%20BIG%20DUMMY%27S%20GUIDE%20TO%20FIDONET).

• Savetz Stuff: Kevin Savetz’s Page o’ Stuff, Internet-related resources and lists by Kevin Savetz (http://www.northcoast.com/savetz/savetz.html).

2.4.5 Navigators

• GNN: Global Network Navigator from O’Reilly (http://gnn.com).

• Hytelnet: assists users in reaching Internet libraries, Free-nets, CWISs, BBSs, and other information sites by telnet (telnet://hytelnet@access.usask.ca).

• Hytelnet web: (http://www.cc.ukans.edu/hytelnet html/START.TXT.html).

• Inter-Links: Internet resources, information, links, from Nova Southeastern University, in Ft. Lauderdale Florida (http://alpha.acast.nova.edu/start.html).

• Netlink: connects you with a wide variety of services, from Washington and Lee University (http://netlink.wlu.edu:1020).

• Washington and Lee: Explore Internet Resources (gopher://liberty.uc.wlu.edu/11/internet).

• WorldWindow: a service of the Washington University Libraries (telnet://library.wustl.edu).

2.5 Searching

2.5.1 New or Noteworthy

• Announcement services-Yahoo:

(http://www.yahoo.com/Computers/World Wide Web/Announcement Services/).

• Archives: summation of Announcements from WWW-ANNOUNCE, NET-HAPPENINGS mailing lists and COMP.INFORSYSTEMS.WWW.* news groups (http://cair-archive.kaist.ac.kr:80/Archive/Announce/).

• Best of Web: a gallery of best of Web Award Winners (http://wings.buffalo.edu/contest).

• Best of GNN: Global Network Navigator’s Best of honorees (http://gnn.com/gnn/wic/best.toc.html).

• Best of/PC Week: a collection of sites from PC Week Labs (http://www.ziff.com/∼pcweek/pcwbests.html).

• Cool/Day: Cool Site of the Day, from InfiNet (http://www.infi.net/cool.html).

• Commerce New: what is new in commercial services on the Web (http://www.directory.net/dir/whats-new.html).

• CUI W3 Catalog: changes to the various W3 Catalog sources for the past week (http://cuiwww.unige.ch/W3catalog/changes.html).

• CyberWeb New: Announcements for WWW Developers (http://www.charm.net/∼web/New.html).

• Fishnet: a weekly list of odds and ends of postings and information from the Internet (http://www.cs.washington.edu/homes/pauld/fishnet/).

• GNN’s WIC New: What’s New with The Whole Internet Catalog, from Global Network Navigator, O’Reilly and Associates (http://nearnet.gnn.com/gnn/wic/nunu.toc.html).

• InfoBank: InfoBank’s collection of new lists (http://www.clark.net/pub/global/new.html).

• Infobot: hotlist database (ftp://ftp.netcom.com/pub/ksedgwic/hotlist/hotlist.html).

• Infosystems–new: (Moderated) announcement of new information systems or resources (news:comp.infosystems.announce).

• Infosystems index: What’s New from comp.infosystems.announce, an index of articles from the newsgroup, by George Ferguson (http://www.cs.rochester.edu/users/grads/ferguson/announce/).

• Info sys announce: send announcements to this address (mailto:infosys@msu.edu).

• Info sys admin: send administrivia to this address (mailto:infosys-request@msu.edu).

• Net happenings newsgroup: news about resources, information, happenings on the network (news:comp.internet.net-happenings).

• Net Happenings archive: archive of a moderated mailing list that announces conferences, publications, newsletters, network tools updates, and network resources (http://www-iub.indiana.edu/cgi-bin/nethaps/).

• New index-Yahoo: Reference-Indices to Web Documents-What’s New (http://www.yahoo.com/Reference/Indices to Web Documents/What s New/).

• New Internet sites: collection of links to new Internet Resources services (gopher://liberty.uc.wlu.edu/11/internet/new internet).

• New Lists: new mailing lists announcements archive (gopher://vm1.nodak.edu/11/Local%20LISTSERV%20Resources/NEW-LIST%20Project).

• New sites: New Internet Resources (via Washington and Lee Law Lib)

(gopher://liberty.uc.wlu.edu/11/internet/new internet).

• Nova New: What’s New on the Internet collection (http://alpha.acast.nova.edu/misc/netnews.html).

• Scout Report: (http://rs.internic.net/scout report-index.html).

• Sites/Day-Yahoo: Reference-Indices to Web Documents-Sites of the Day entry from Yahoo (http://www.yahoo.com/Reference/Indices to Web Documents/Sites of the Day/).

• Useful/cool things: daily showcase of one pointer to something useful on the Internet, and one pointer to something cool (http://www.teleport.com/∼lynsared/useful.html).

• Web Announce: New resources about the World Wide Web, moderated newsgroup comp.infosystems.www.announce (http://www.halcyon.com/grant/Misc/charter.html).

• webNews: announcements of new Web sites, services, and software, a service of the Department of Computer and Information Sciences at the University of Alabama at Birmingham, Studies of Information Filtering Technology for Electronic Resources (SIFTER) Research Group (http://twinbrook.cis.uab.edu:70/webNews.80).

• webNews search: a database of recent articles from Usenet news about the Web (gopher://twinbrook.cis.uab.edu/7GO/webNews.70).

• What’s New/Gopher: what’s knew on known public gophers (gopher, telnet, wais) (gopher://liberty.uc.wlu.edu/11/internet/new internet).

• What’s New/Meta: a list of what’s new pages, by Meng Weng Wong (http://www.seas.upenn.edu/∼mengwong/whatsnew.list.html).

• What’s New/Web: What’s New With NCSA Mosaic; actually shows Web-wide new resources (http://www.ncsa.uiuc.edu/SDG/Software/Mosaic/Docs/whats-new.html).

• What’s Cool/Web-Netscape: a ’What’s cool’ page developed at Netscape communications (http://home.mcom.com/home/whats-cool.html).

• Yahoo New: What’s New on Yahoo Hierarchical Hotlist (http://www.yahoo.com/new/).

2.5.2 Resource Lists

• ALICE: Annotated Listings for Internet and Cyberspace Explorers (http://www.umanitoba.ca/ALICE/index.html).

• Awesome List: a list of useful resources, by John Makulowich (http://www.clark.net/pub/journalism/awesome.html).

• CMC resources: resources related to the study of Computer-Mediated Communication, from the CMC Studies Center (http://www.rpi.edu/∼decemj/cmc/resources.html).

• Creative/Internet: The Creative Internet Home Page, a collection of fun and useful resources (http://www.galcit.caltech.edu/∼ta/creative.html).

• Drakos List: Subjective Electronic Information Repository, by Nikos Drakos (http://cbl.leeds.ac.uk/nikos/doc/repository.html).

• Email Services: a list of services available via email, by David DeSimone (ftp://sunsite.unc.edu/pub/docs/about-the-net/libsoft/email services.txt).

• Exploring: Exploring the Internet, by Ed Kubaitis (http://www.cen.uiuc.edu/exploring.html).

• Freeside FAQ: Meng surfs the Internet (http://www.seas.upenn.edu/∼mengwong/fsfaq.html).

• GNN’s Internet Page: Internet information, from Global Network Navigator (http://nearnet.gnn.com/wic/internet.toc.html).

• Hot/Cool List: What’s Hot and Cool on the Web, art/music/interesting emphasis (http://kzsu.stanford.edu/uwi/reviews.html).

• InfoBank: Information Bank’s collection of Internet resources links (http://www.clark.net/pub/global/front.html).

• Internet-EIT: Enterprise Integration Technologies Web Resources (http://www.eit.com/web/web.html).

• Internet-Enns: a collection of resources that are useful to Internet trainers, as well as just about anyone else who is on the Net, by Neil Enns (http://www.brandonu.ca/∼ennsnr/Resources/).

• Internet FAQ: Internet Services Frequently Asked Questions (FAQ) list for alt.internet.services Usenet newsgroup (ftp://rtfm.mit.edu/pub/usenet/news.answers/internet-services/faq).

• Internet Meta-Index: Internet Resources Meta-Index, from the National Center for Supercomputing Applications (NCSA) (http://www.ncsa.uiuc.edu/SDG/Software/Mosaic/MetaIndex.html).

• Justin’s/Underground: Justin’s Links from the Underground (http://www.links.net).

• MaasInfo Indexes: The index of indexes of Internet online files, by Robert Elton Maas (ftp://NCTUCCCA.edu.tw/documents/Internet/MaasInfo/).

• May List: lists telnet sites/services, by CC May (ftp://aug3.augsburg.edu/files/bbs lists/nal008.txt).

• Planet Earth: planet earth home page, a list of things on the Internet, by Richard Bocker (http://white.nosc.mil/info.html).

• Power Index: from Web Communications, lists a variety of resources in many categories (http://www.webcom.com/power/index.html).

• Smith’s BigFun List: telnet, ftp, and other sites, by Jeremy Smith (ftp://owl.nstn.ns.ca/pub/netinfo/bigfun.txt).

• Spider’s Web: a lists of links to lists, web/net stuff, searchers, images, and references (http://gagme.wwa.com/∼boba/spider.html).

• ThesisNet FAQ: a summary of academic resources pertaining to cyberspace (http://www.seas.upenn.edu/∼mengwong/thesisfaq.html#ftp).

• Thousand/Sites: A Thousand Points of Sites, a random way (!) to encounter web sites (http://legendre.ucsd.edu/Research/Fisher/Home/randomjump.html).

• Tong’s Collection: Links to other sites, a miscellaneous collection of information and entertainment, by Andrew Tong (http://www.ugcs.caltech.edu/∼werdna/info.html).

• Top Tens: from Internet Training and Consulting Services, resource lists which are picks of the Net and Web in Top-level resources, art, commerce, fun, Internet training, K-12 education, library, Internet books and journals (http://www.itcs.com/itcs/topten.html).

• URouLette: a way to psuedo-randomly choose a URL to visit (http://kuhttp.cc.ukans.edu/cwis/organizations/kucia/uroulette/uroulette.html).

• Useless/WWW: ’America’s Funniest Home Hypermedia,’ a collection of strange, trivial, bizarre, unusual, and weird WWW pages, collected by Paul Phillips (http://www.primus.com/staff/paulp/useless.html).

• Yale Overview: General Overview of Worldwide Internet Resources (http://www.cs.yale.edu/HTML/WorldWideWebTop.html).

• Yanoff’s Services List (ftp): Special Internet Connections (Internet Services List), by Scott Yanoff (ftp://ftp.csd.uwm.edu/pub/inet.services.txt).

• Yanoff List HTML: Hyptertext markup language version of Scott Yanoff’s Special Internet Connections listing of resources by subject (http://www.uwm.edu/Mirror/inet.services.html).

• WWW/Internet: a collection of links to manuals and demos (http://tecfa.unige.ch/info-www.html).

2.5.3 Subjects

• CSOIRG Home Page: The Clearinghouse for Subject-Oriented Internet Resource Guides at the School of Information and Library Studies, University of Michigan (http://www.lib.umich.edu/chhome.html).

• CSOIRG ftp: subject-specific Internet Resource Guides collected by UMich (ftp://una.hh.lib.umich.edu/inetdirs/).

• CSOIRG gopher: UMich collection via gopher (gopher://una.hh.lib.umich.edu/11/inetdirs).

• CyberSight: unique subjects on the net (http://cybersight.com/cgi-bin/cs/s?main.gmml).

• CyberZine: a yellow pages user friendly guide to Cyberspace (http://cyberzine.org/index.html).

• Galaxy: a service of EINet, a collection of information, searchable via index or by topic trees (http://www.einet.net/galaxy.html).

• GNN WIC: Global Network Navigator’s Whole Internet Catalog, from O’Reilly and Associates (http://nearnet.gnn.com/wic/newrescat.toc.html).

• Gopher Jewels gopher: subject-oriented gopher list (gopher://cwis.usc.edu/11/Other Gophers and Information Resources/Gophers by Subject/Gopher Jewels).

• Gopher Jewels web: a collection of subject-oriented gophers (http://galaxy.einet.net/GJ/index.html).

• Gopher Jewels links: .links format (ftp://usc.edu/pub/gopher/).

• Gopher Jewels Info: lists information available about Gopher Jewels (mailto:listproc@einet.net Body: get gopher-lists.txt).

• Gopher Trees: a collection of subjet-oriented gopher trees (gopher://burrow.cl.msu.edu/11/internet/subject).

• Hyperdex: hyperindex over webNews and more (http://twinbrook.cis.uab.edu:70/hhyperdex.80).

• INFOMINE: resources covering a range of disciplines, by the Library of the University of California-Riverside (http://lib-www.ucr.edu).

• Internet at Large: Subject-oriented listing (http://www.sdsc.edu/1/SDSC/Geninfo/Internet).

• InterNIC Dir of Dirs: InterNIC’s Directory of Directories, subject-oriented listing of registered Internet resources and services (http://ds.internic.net/ds/dsdirofdirs.html).

• IWT Narrative: Internet Web Text’s narrative about subject-oriented searching (http://www.rpi.edu/Internet/Guides/decemj/nar-subject.html).

• Joel’s List: Joel’s Hierarchical Subject Index (http://www.cen.uiuc.edu/∼jj9544/index.html).

• LOC/resources: US Library of Congress code for classifying resources on the Net (gopher://info.anu.edu.au/11/elibrary/lc).

• LOC/VL: US Library of Congress categorization of WWW Virtual Library (http://www.w3.org/hypertext/DataSources/bySubject/LibraryOfCongress.html).

• Mother/BBS: Mother of all BBS, a project to allow users to add information to a subject tree, developed by Oliver McBryan (http://www.cs.colorado.edu/homes/mcbryan/public html/bb/summary.html).

• Planet Earth: Richard P. Bocker created this web to collect a wide variety of Internet resources (http://white.nosc.mil/info.html).

• Starting Point: (http://www.stpt.com).

• Study Carrels: Subject and Discipline-Specific Internet Resources, from North Carolina State University Library (http://dewey.lib.ncsu.edu/disciplines/index.html).

• Subject Lists: Resources Classified by Subject (LC Classification)

(gopher://info.anu.edu.au:70/11/elibrary/lc).

• UMBC Web: University of Maryland–Baltimore County, exceptional Internet based resources by subject category (http://umbc7.umbc.edu/∼jack/subject-list.html).

• USENET FAQ Index: forms-based gateway to usenet FAQs, by INTAC (http://www.intac.com/FAQ.html).

• USENET Periodic Postings: list of Usenet periodic postings (ftp://rtfm.mit.edu/pub/usenet/news.answers/periodic-postings/).

• USENET repository: collection of FAQ’s and files from Usenet newsgroups (ftp://rtfm.mit.edu/pub/usenet/).

• USENET news.answers: a hypertext presentation of the answer lists posted in the news.answers newsgroup (http://www.cis.ohio-state.edu/hypertext/faq/usenet/).

• WAIS subject tree: a list of WAIS servers separated into subject areas (http://www.ub2.lu.se/auto new/UDC.html).

• WAIS Servers: directory of wide area information servers (wais://cnidr.org:210/directory-of-servers?).

• WAIS sources: a collection of WAIS sources (gopher://liberty.uc.wlu.edu/11/internet/indexsearches/inetsearches).

• Web of Wonder: links to Web resources on various subjects, by Lance Weitzel (http://www.digimark.net/wow/).

• Yahoo: Yet Another Hierarchically Odiferous Oracle, an extendible collection of subjects (http://www.yahoo.com/%CHECKED).

• Yanoff List HTML: Hyptertext markup language version of Scott Yanoff’s Special Internet Connections listing of resources by subject (http://www.uwm.edu/Mirror/inet.services.html).

• WWW VL: World Wide Web Virtual Library, a large hypertext collection of information organized by subject (http://www.w3.org/hypertext/DataSources/bySubject/Overview.html).

• WWW VL/LOC: The World-Wide Web Virtual Library viewed as US Library of Congress Classification (http://www.w3.org/hypertext/DataSources/bySubject/LibraryOfCongress.html).

2.5.4 Keyword

• Academic lists: searchable index of academic email conferences (wais://munin.ub2.lu.se:210/academic email conf).

• ALIWEB: Archie-Like Indexing for the Web, by Martijn Koster (http://web.nexor.co.uk/aliweb/doc/aliweb.html).

• Archieplex: Archie via Web–access archie servers (search ftp sites) via the web (http://web.nexor.co.uk/archie.html).

• CUSI: SUSI), by Martijn Kosterrch interface; a forms-based interface into many indices, engines, and Web Spider databases (forerunner (http://web.nexor.co.uk/public/cusi/doc/list.html).

• CUSI-R: Customizable Unified Search Index via Radio Buttons, by David Rosen (http://www.scs.unr.edu/∼cbmr/net/search/cusi-r.html).

• Discussion groups: search a list of discussion groups, Bitnet and Internet interest groups (Dartmouth list) (http://alpha.acast.nova.edu/cgi-bin/lists).

• External info: collects some of the most useful search engines available on the WWW (http://www is.cs.utwente.nl:8080/cgi-bin/local/nph-susi1.pl).

• Gloss: Glossary-of-Servers Server, a system to find data sources that match keyword queries (http://gloss.stanford.edu).

• Gopher Jewels Search: Search the Gopher Jewels (a collection of subject-oriented gophers) (http://galaxy.einet.net/gopher/gopher.html).

• GNA Meta-Library: search the Globalwide Networking Academy (GNA) library of resources and information (http://uu-nna.mit.edu:8001/uu-nna/meta-library/index.html).

• Hytelnet search gopher: given a keyword, returns Hytelnet entries (gopher://liberty.uc.wlu.edu:3004/7).

• Hytelnet Web search: search all Hytelnet resource entries, via a web form, from Galaxy (http://galaxy.einet.net/hytelnet/HYTELNET.html).

• Infoseek: a keyword searching service ($) over WWW pages, Usenet, computer magazines, newspaper newsires, press releases, company profiles, moview reviews, technical support databases (http://www.infoseek.com).

• IWeb: Web search service, allows you to find resources matching a keyword query; allows you to contribute to the database by registering (http://sparta.lcs.mit.edu/iweb/welcome.html).

• JumpStation: referencing the information available on the World Wide Web (http://www.stir.ac.uk/jsbin/js).

• NIKOS: An Internet Resource Locator (http://www.rns.com/cgi-bin/nomad).

• Shase: Virtual Shareware Library (http://www.fagg.uni-lj.si/SHASE/).

• WAISGATE: WAIS to WWW gateway, search WAIS databases through search terms (http://www.wais.com/directory-of-servers.html).

• Web Catalog/CUI: is a collection of URL references built from a number of hand-crafted HTML lists, from Centre Universitaire d’Informatique, l’Universite de Geneve (http://cuiwww.unige.ch/w3catalog).

• Web Search Engines: a meta-index of search engines on the Web, with a forms interface, from Centre Universitaire d’Informatique, l’Universite de Geneve (http://cuiwww.unige.ch/meta-index.html).

• Web Spiders info: search the web for information about resources, collecting information into a database which can be queried; e.g., Web Crawler, Web Nomad, Web Worm, RBSE database, Lycos (Araneida, Lycosidae, Lycosa), Harvest Brokers; entry from Internet Tools Summary (http://www.rpi.edu/Internet/Guides/decemj/itools/nir-tools-spiders.html).

• Web Spiders web: Wanderers, Spiders and Robots; includes list of known robots/spiders, guidelines, standard for robot exclusion, by Martijn Koster (http://web.nexor.co.uk/mak/doc/robots/robots.html).

2.5.5 Spaces

• FTP Sites list: list of Internet sites accepting anonymous ftp, maintained by Perry Rovers (ftp://rtfm.mit.edu/pub/usenet/news.answers/ftp-list/).

• FTP Sites Web: Web interface to Perry Rover’s FTP site monster list (http://www.info.net/Public/ftp-list.html).

• Gopher Sites: list of all gophers (long) (ftp://liberty.uc.wlu.edu/pub/lawlib/all.gophers.links).

• Telnet–Hytelnet/telnet: access to the Hytelnet Service through telnet, by Peter Scott (telnet://hytelnet@access.usask.ca).

• Telnet-Hytelnet/WWW: access to Hytelnet by Peter Scott, provided by the University of Kansas (http://www.cc.ukans.edu/hytelnet html/START.TXT.html).

• WAIS Servers: list of Wide Area Information Servers (WAIS), by WAIS, Inc.

(http://www.wais.com).

• WWW Servers/list: Wandex’s list of WWW sites (http://www.netgen.com/cgi/wandex).

• WWW Servers/sites: A comprehensive list of WWW sites generated from a Web Wanderer program, by Matthew Gray (http://www.netgen.com/cgi/comprehensive).

• WWW Servers/geo: a long list of registered WWW servers listed geographically by continent and country (http://www.w3.org/hypertext/DataSources/WWW/Servers.html).

• WWW Sites(CityLink): state and city web sites (http://www.neosoft.com/citylink/).

• WWW Sites(Virtual Tourist): a geographic map to aid in locating Web sites and other resources (http://wings.buffalo.edu/world).

• Web Sites(City Net): explore and browse resources by geography (continent, region, country, city) (http://www.city.net).

• Web Sites index: index to WWW sites, by John Doyle (http://herald.usask.ca/∼scottp/home.html).

• Web Sites/Time: Timex World Time–shows Time zones with current time and WWW sites in each (http://www.timeinc.com/vibe/vibeworld/worldmap.html).

2.5.6 People

• Directory services: a collection of white pages servers to look up people (gopher://gopher.nd.edu/11/Non-Notre%20Dame%20Information%20Sources/Phone%20Books– Other%20Institutions).

• Find a friend: mainframe resources to locate people; Searches cover the entire United States. Results are E-mailed within 48 hours (http://www.ais.net/findafriend).

• Finding people: a collection of resources to help you locate a specific person on the net (gopher://yaleinfo.yale.edu:7700/11/Internet-People).

• Home Pages/directories: a collection of personal home pages lists and directories (http://www.rpi.edu/Internet/Guides/decemj/icmc/culture-people-lists.html).

• Knowbot: provides a uniform user interface to heterogeneous remote information services (Internic Point of contacts, MCImail, x500 databases, finger, nwhois, etc.) (telnet://info.cnri.reston.va.us:185).

• Netfind: a simple Internet ‘white pages’ user directory (http://www.rpi.edu/Internet/Guides/decemj/itools/nir-utilities-netfind.html).

• Netsearch(tm): a database of companies and contacts (http://www.ais.net:80/netsearch/).

• Searching for People: e-mail addresses, phone books, from Washington and Lee (gopher://liberty.uc.wlu.edu/11/internet/personsearches).

2.6 Directories

• American Universities: (http://www.clas.ufl.edu/CLAS/american-universities.html).

• CS Depts: (http://www.cs.cmu.edu:8001/Web/People/anwar/CS-departments.html).

• CWIS Web: a listing of Campus-Wide information systems, by Polly-Alida Farrington (http://www.rpi.edu/Internet/cwis.html).

• Commercial List: List of Commercial Services on the Web, from Open Market (http://www.directory.net).

• Dartmouth Merged SIGL: Special Interest Group Lists (ftp://dartcms1.dartmouth.edu/siglists/).

• Directory Services-Yahoo: (http://www.yahoo.com/Computers/Internet/Directory Services/).

• Electronic Conferences: Directory of Scholarly Electronic Conferences, by Diane K. Kovacs (ftp://ksuvxa.kent.edu/library/acadlist.readme).

• Electronic Conferences: browse or search the Directory of Scholarly Electronic Conferences, by Diane K. Kovacs (http://www.austin.unimelb.edu.au:800/1s/acad).

• Electronic Conferences: gopher presentation of the Directory of Scholarly Electronic Conferences, by Diane K. Kovacs (gopher://info.monash.edu.au/11/Other/lists).

• Electronic Journals/ARL: Association of Research Libraries Directory of Electronic Journals and Newsletters (gopher://arl.cni.org:70/11/scomm/edir).

• Electronic Journals/Strangelove: Directory of Electronic Journals and Newsletters, by Michael Strangelove (ftp://ftp.cni.org/pub/net-guides/strangelove/).

• Electronic Magazines/archive: an archive of various electronic magazines (ftp://etext.archive.umich.edu/pub/Zines/).

• Electronic Resources: lists resources (CWIS, texts, guides, WWW, WAIS, etc.) by University of Waterloo Library (gopher://watserv2.uwaterloo.ca/11/servers).

• EZines List: summary of electronic zines by John Labovitz (http://www.meer.net/∼johnl/e-zine-list/).

• Finding Lists: how to find LISTSERV lists (mailto:listserv@vm1.nodak.edu Body: get NEW-LIST wouters).

• Free-Nets Home Page: Free-net information, presented as a public service by Peter Scott, at the University of Saskatchewan Libraries (http://herald.usask.ca/∼scottp/free.html).

• Free Databases: a catalog databases that are available without payment (ftp://idiom.berkeley.ca.us/pub/free-databases).

• InterActive Yellow Pages: by Lighthouse Productions–telephone-style yellow pages international directory (mailto:tuna@netcom.com Subject: yellow pages Body: YOUR EMAIL ADDRESS).

• IYP: InterActive Yellow Pages (ftp://ftp.netcom.com/pub/tuna/).

• InfoMagnet: helps you find, join, search and participate in LISTSERV discussion groups on the Internet (http://www.clark.net/pub/listserv/imag.html).

• Internet Computer Index: lists of PC, Macintosh, and Unix-related resources (http://ici.proper.com).

• Internet Consultants: CommerceNet’s directory of Internet consultants (products and services) (http://www.commerce.net/directories/consultants/consultants.html).

• Internet Trainers/consultants: (gopher://gopher.fonorola.net/00/Internet%20Business%20Journal/Directory of Trainers and Consultants).

• Interest Groups List: The Lists of Lists, a listing of special interest group mailing lists available on the Interne (ftp://sri.com/netinfo/interest-groups.txt).

• Japanese: key information sources on Japan currently available over the Internet (http://fuji.stanford.edu/japan information/japan information guide.html).

• Library Access Script: a script to login to online card catalogs (ftp://sonoma.edu/pub/libs.tar).

• Library Guide: Accessing on-line Bibliographic Databases, by Billy Barron and Marie-Christine Mahe (ftp://ftp.utdallas.edu/pub/staff/billy/libguide).

• Library Catalogs Noonan: a list of academic and public library catalogs online, by Dana Noonan (ftp://sunsite.unc.edu/pub/docs/about-the-net/libsoft/guide2.txt).

• Library Catalogs St Geo: Internet - Accessible Library Catalogs and Databases, Art St. George and Ron Larsen (ftp://nic.cerf.net/internet/resources/library catalog/).

• NET-LETTER GUIDE: A newshound’s guide to newsy periodicals available through the Internet (mailto:higgins@dorsai.dorsai.org Body: SUBSCRIBE NET-LETTER).

• Net Orgs: Outposts on the Electronic Frontier, International, Groups Supporting the Online Community (ftp://rtfm.mit.edu/pub/usenet/news.answers/net-community/orgs-list).

• Nonprofit Center: supports and provides information about non-profit organizations (gopher://envirolink.org/00/.EnviroOrgs/.inc/.center).

• Nonprofits: Nonprofit Organizations on the Internet, by Ellen Spertus (http://www.ai.mit.edu/people/ellens/non.html).

• Online books: On-line books (http://www.cs.cmu.edu:8001/Web/books.html).

• Online bibs: On-line Bibliographies and Journal Contents (http://www.cs.cmu.edu:8001/Web/bibliographies.html).

• Online journals: On-line Journals (http://www.cs.cmu.edu:8001/Web/journals.html).

• Online libraries: (http://www.cs.cmu.edu:8001/Web/e-libraries.html).

• PAML: Publicly Available Mailing Lists, a list of mailing lists available primarily through the Internet and the UUCP network, by Stephanie da Silva (http://www.neosoft.com/internet/paml).

• Patents: patent searching by class/subclass, by Gregory Aharonian (http://sunsite.unc.edu/patents/intropat.html).

• Publisher’s Catalogs: (http://herald.usask.ca/∼scottp/publish.html).

• Reference: Reference Resources via the World Wide Web (http://vm.cfsan.fda.gov/referenc.html).

• Reference-Yahoo: Reference section from Yahoo (http://www.yahoo.com/Reference/).

• Software downloads: publicly available sites for software (http://alpha.acast.nova.edu/downloads.html).

• Technical Reports: technical reports archive (ftp://daneel.rdt.monash.edu.au/pub/techreports/).

• UNCAT: The electronic catalog of uncataloged titles (businesses, nonprofit groups, trade associations, museums and self-published authors), from Sapphire Press (http://www.sapphire.com/UNCAT/).

• Virtual Tourist: geographical directory of WWW Information Services (http://wings.buffalo.edu/world).

• WAIS sources: a brief description of the content of many WAIS sources on the Internet, grouped into relevant categories (ftp://kirk.bond.edu.au/pub/Bond Uni/doc/wais/readme).

• WAIS Directory of Servers: (wais://cnidr.org:210/directory-of-servers?).

• WAIS Databases: a hypertext list (http://kaml1.csi.uottawa.ca:3000/wais.html).

• WWW Sites: comprehensive list–a long list of WWW sites gathered through an automated program, Webwanderer, by Matthew Gray (http://www.mit.edu:8001/afs/sipb/user/mkgray/ht/compre3.html).

2.7 Services

• Anonymous post: (mailto:anonymus+ping@tygra.michigan.com Body: help).

• Anonymous post: (mailto:help@anon.penet.fi Body: help).

• Anonymous remail: (mailto:remail@tamsun.tamu.edu Subject: remail help).

• BITFTP: obtain files via anonymous ftp on Bitnet (mailto:Bitftp@PUCC.Princeton.edu Body: help).

• Correct Time/NBS: tells National Bureau of Standards correct time (telnet://india.colorado.edu:13).

• Domain Name Lookup: translate domain names to numeric IP address (telnet://lookup@130.59.1.40).

• Email FTP: get files at anonymous ftp sites via email. (mailto:ftpmail@decwrl.dec.com Body: help).

• Email Gopher: use a gopher via email (mailto:gophermail@ncc.go.jp Body: help).

• Email Gopher: use a gopher via email (mailto:gophermail@calvin.edu Body: help).

• Email WWW: www URL); URL = Uniform resource locator (body (mailto:listproc@www0.cern.ch Body: help).

• Email Usenet: post to newsgroup via email (mailto:hierarchy-group-name@cs.utexas.edu Subject: Your Subject Body: Your Contents).

• FAXNET: FAX’s via email ($) (mailto:info@awa.com Body: help).

• FAX FAQ: How can I send a fax from the Internet?, by Kevin Savetz (ftp://rtfm.mit.edu/pub/usenet/news.answers/internet-services/fax-faq).

• HTML Validation:

(http://www.cc.gatech.edu/grads/j/Kipp.Jones/HaLidation/validation-form.html).

• IP address resolver: given a site name, obtain its numerical IP address (mailto:resolve@widener.edu Body: site SITE NAME).

• Mail name server: gives email information about many email names (e.g. IP address given name) (mailto:dns@grasp.insa-lyon.fr Body: help).

• RFCs via mail: Request For Comments–documents about various issues for discussion, covering a broad range of networking issues. (mailto:rfc-info@isi.edu Body: help: ways to get rfcs).

3 APPLICATIONS 20

3 APPLICATIONS

This section lists examples and lists of examples of applications of CMC on the Internet in a variety of areas.

3.1 Commerce

3.1.1 Information

• Advertising Blacklist: a compendium of advertisers who have misused Net access, compiled by Axel Boldt (http://math-www.uni-paderborn.de/∼axel/BL/blacklist.html).

• Advertising Guide: The Internet Advertising Resource Guide, maintained by Hairong Li, The Missouri School of Journalism; links to a large number of sources of information about Internet advertising, including collections and topics, studies, storefronts, and other information (http://www.missouri.edu/internet-advertising-guide.html).

• Advertisting FAQ: Internet Advertising FAQ, by Strangelove Internet Enterprises (mailto:interBEX1@intnet.bc.ca).

• Advertising/marketing law: emphasis on infomercials, home shopping, and direct response tv, includes intellectual property and telemarketing, by Lewis Rose (http://www.webcom.com/∼lewrose/home.html).

• BizWeb: a subject-oriented directory of companies on the Web (http://www.bizweb.com).

• Businesses: List of Commercial Services on the Web, from Open Market (http://www.directory.net).

• Business/Corporations: lists of business and corporations on the Net (http://www.yahoo.com/Economy/Business/Corporations/).

• Business sites: Interesting Business Sites on the Web, by Bob O’Keefe at the School of Management, Rensselaer Polytechnic Institute (http://www.rpi.edu/∼okeefe/business.html).

• Business uses: Commercial Use (of the net) Strategies Home Page, by Andrew P. Dinsdale (http://www.dserv.com/∼andrew/business.html).

• Canadian Business: Candadian Internet Business Directory (http://cibd.com/cibd/CIBDHome.html).

• Career Mosaic: high-tech companies offering career information (http://www.careermosaic.com/cm/).

• Commercial List: a directory of many Commercial Services on the Web (http://www.directory.net).

• Commercial Services: a list of telnet connections to many commercial information services, from The World (gopher://gopher.std.com:70/11/Commercial).

• Computer+Communications: Computer and Communication Company Sites on the Web, by James E. (Jed) Donnelley (http://www-atp.llnl.gov/companies.html).

• Entrepeneurs: useful business information and services for entrepreneurs (http://sashimi.wwa.com/∼notime/eotw/EOTW.html).

• FECRS: Fairfax Electronic Commerce Resource Center; Continuous Acquisition Lifecycle Support (CALS); enterprise integration, electronic commerce and business processing, re-engineering (http://www.ecrc.gmu.edu/index.html).

• Hermes: research project on the commercial uses of the World Wide Web (http://www.umich.edu/∼sgupta/hermes.htm).

• IBC: Internet Business Center is a World-Wide Web server for information specifically related to business use of the Internet (http://www.tig.com/IBC/).

• IBD: Internet Business Directory, product/service information (http://ibd.ar.com).

• InterBEX: Business Exchange, selective content oriented business information (mailto:interBEX-index@intnet.bc.ca).

• InterQuote: continuously updating stock market information service (http://wwa.com/∼quote).

• Internet info: compiles information on the commercial activity on Internet (mailto:info@internetinfo.com).

• Inet Marketing: archives of the Internet Marketing mailing list (http://galaxy.einet.net/hypermail/inet-marketing/).

• Marketing/CMEs: Marketing in Computer-Mediated Environments Home Page, from Owen Graduate School of Management, Vanderbilt University, Nashville, TN, USA (http://colette.ogsm.vanderbilt.edu).

• MARTECH: discuss MARketing with TECHnology tools, such as marketing via the Internet (mailto:LISTSERV@cscns.com Body: subscribe MARTECH YOUR NAME).

• MESCH: The Multi-WAIS Engine for Searching Commercial Hosts, allows you to search the databases of several commercial WWW providers (http://www.ip.net/cgi-bin/mesch).

• Publications: Print Publications Related to Business Use of the Internet, from Tenagra (http://arganet.tenagra.com/Tenagra/books.html).

• Stock Quotes: QuoteCom, a service dedicated to providing financial market data to Internet users (http://www.quote.com).

• Thomas Ho: favorite Electronic Commerce WWW resources, includes information sources, links to articles, economic development, service/presence providers (http://www.engr.iupui.edu/∼ho/interests/commmenu.html).

• What’s New/Commerce: what is new in commercial services on the Web (http://www.directory.net/dir/whats-new.html).

3.1.2 Marketplaces

—– Lists of Marketplaces

• Shop D Net: a list of virtual marketplaces, from Blake and Associates, Internet Marketing Consultants (http://www.neosoft.com/citylink/blake/malls.html).

—– Selected Instances of Marketplaces

• CommerceNet: Internet-based infrastructure for electronic commerce, created and operated by a consortium of major Silicon Valley users, providers and developers under Smart Valley, Inc. (http://logic.stanford.edu/cit/commercenet.html).

• Digital’s Emall: Digital Equipment Corporation’s Electronic Shopping Mall (http://www.service.digital.com/html/emall.html).

• eMall: WWW shopping and information (http://eMall.com).

• IBC: Internet Business Connection, An electronic shopping mall and a service for companies that would like to promote their products or services on the Internet (http://www.charm.net/∼ibc).

• Internet Mall: Shopping on the Information Highway A monthly list of commercial services available via the Internet (ftp://netcom.com/pub/Guides/).

• Internet Shopping Network: products from hundres of vendors, along with query interface (http://shop.internet.net).

• Sofcom: home shopping, information providers, no-profits (http://www.sofcom.com.au).

3.2 Communication

3.2.1 Individual

• CMC-Interpersonal: list of interpersonal tools for CMC (mail, talk)

(http://www.rpi.edu/Internet/Guides/decemj/itools/cmc-interpersonal.html).

• Home Pages: a collection of personal home pages lists (http://www.rpi.edu/Internet/Guides/decemj/icmc/culture-people-lists.html).

3.2.2 Group

• CCCC: Computerized Conferencing and Communications Center at New Jersey Institute of Technology (NJIT) (http://it.njit.edu/njIT/Department/CCCC/default.html).

• CSCW Info: Computer-Supported Cooperative Work information (groupware), unOfficial Yellow Pages of CSCW (ftp://gorgon.tft.tele.no/pub/groupware/).

• CSCW Lab: Computer-Supported Cooperative Work laboratory at the Fraunhofer-Institut fur Graphische Datenverarbeitung, Germany (http://www.igd.fhg.de/cscw-lab/).

• CSCW pages: The unOfficial Yellow Pages of CSCW (http://www.tft.tele.no/cscw/).

• Collaborative Comm: Collaborative Networked Communication–MUDs as Systems Tools, by Remy Evard (http://www.ccs.neu.edu/USER/remy/documents/cncmast.html).

• Collaborative Design: Architecture and CAAD, Swiss Federal Institute of Technology, Zurich (http://www.arch.ethz.ch/∼dave/cmc-collab.html).

• CRTR-U of CO: Collaboration Technology Research Group (group user interfaces, distributed editors, workflow systems, cooperativee grown information systems), University of Colorado (http://www.cs.colorado.edu/homes/carlosm/public html/ctrg.html).

• Collaboration: Workshop on Wide-Area Collaboration and Cooperative Computing (http://www.ai.mit.edu/projects/iiip/colab/workshop.html).

• Communication Archive: Sunsite Communication archive–papers, logs, information, including April Fools Page, maintained by David Barberi (ftp://sunsite.unc.edu/pub/academic/communications/).

• Communication Archive Web: Sunsite Communication archive–papers, logs, information, including April Fools Page, maintained by David Barberi (http://sunsite.unc.edu/dbarberi/communications.html).

• Computer Network Conf: A Discussion on Computer Network Conferencing, by D. Reed (ftp://nic.merit.edu/documents/rfc/rfc1324.txt).

• Groupware Yellow Pages: information resources, from Consensus (http://www.consensus.com:8300/GWYP TOC.html).

• Internet Citizen’s Band: internet group teleconferencing program (http://www.echo.com/∼kzin/icb.html).

• Internet Relay Chat(IRC): real-time, many-many text discussion divided into channels (ftp://cs.bu.edu/irc/support/).

• IRC FAQ: Internet Relay Chat Frequently Asked Questions and Answers (http://www.kei.com/irc.html).

• ISO/IEC stds: draft of standards on standard on computer conferencing (gopher://mars.dsv.su.se:70/0/iso-mess/gc/X.acc-First CD.TXT).

• MOO Papers: Pavel Curtis’ collection of MU* papers (ftp://ftp.parc.xerox.com/pub/MOO/papers/).

• MUD info: Multiple User Dialogue/Dimension/Dungeon FAQs, Lists, Information, Collections, Servers, Archives, Newsgroups, by Lydia Leong (http://www.cis.upenn.edu/∼lwl/mudinfo.html).

• MUD papers/Cherny: gender and discourse related MUD papers, by Lynn Cherny (http://bhasha.stanford.edu/∼cherny/papers.html).

• NCW: National Center for the Workplace (NCW) Gopher (gopher://uclink.berkeley.edu:3030/1).

• Project H: an ongoing computer-mediated collaboration, on a Comserve-sponsored hotline, of more than a hundred international researchers (http://www.arch.su.edu.au/PROJECTH/index.html).

• SHARE: A Methodology and Environment for Collaborative Product Development (http://www.eit.com/projects/share/share/share-home.html).

• Tools: a list of Internet tools for group communication (http://www.rpi.edu/Internet/Guides/decemj/itools/cmc-group.html).

3.2.3 Organizational

• Client/Server: Client/Server Coffeehouse, discussion about client-server systems in organizations (http://www.onr.com/clients.html).

• CORPS: Computing, Organizations, Policy, and Society, study of computerization in organizations, Information and Computer Science, University of California, Irvine (http://www.ics.uci.edu/CORPS/).

• CWIS List: Campus-Wide Info Systems list (ftp://sunsite.unc.edu/pub/docs/about-the-net/cwis/cwis-l).

• CWIS Paper: Campus-Wide Info Systems by Judy Hallman (ftp://sunsite.unc.edu/pub/docs/about-the-net/cwis/hallman.txt).

• CWIS Web: a listing of Campus-Wide information systems, by Polly-Alida Farrington (http://www.rpi.edu/Internet/cwis.html).

• Campus Net: Creating a Campus Network Infrastructure, by Steve Griffin (ftp://gandalf.iat.unc.edu/technote/teknote4.txt).

• Campus Net Bib: Creating a Campus Network Infrastructure–Bibliography (ftp://gandalf.iat.unc.edu/guides/irg-15.txt).

• IOS: Interorganizational Systems (IOS) WorldNet (http://www-iwi.unisg.ch/iswnet/index.html).

• OMT: Organization and Management Theory Division the Academy of Management, to foster communication and to serve as an information resource to Organization scholars worldwide (http://cwaves.stfx.ca/Subjects/Business/omt.html).

3.2.4 Mass

—– Lists

• Campus Newspapers: Campus Newspapers on the Internet, by Jonathan Bell (http://beacon-www.asa.utk.edu/resources/papers.html).

• Internet news: Internet accessible news, Newspapers, tv, radio, press services, and publications, by C. Sam Sternberg (ftp://ftp.shell.portal.com/pub/jshunter/news.html).

• Commercial News/Web: a list of commercial news services on the WWW, by Gary Ritzenthaler (http://www.jou.ufl.edu/commres/webjou.htm).

• Communication Resources: Communications and Mass Media Resources, links, reviews, and resources for scholars and those interested in Journalism, Mass Media, and Communications Research, by Steve Brown (http://www.jou.ufl.edu/commres/commhome.htm).

• CyberNews: student-run media connectivity, newsire on the Internet (http://www.hmc.edu/www/people/teverett/cybernews/Home.html).

• Daily business: Daily sources of business and economic news (http://www.helsinki.fi/∼lsaarine/news.html).

• Electronic newspapers: by Steve Outing (mailto:majordomo@marketplace.com Body: get online-news online-newspapers.list).

• ENewsstand gopher: information provided by U.S.magazine publishers (gopher://gopher.internet.com).

• GSN: Global Student News, a collaboration of student journalists, by Gary Ritzenthaler (http://www.jou.ufl.edu/forums/gsn/).

• Journalism/Comm schools: World Wide Web sites of interest to those in the fields of journalism and mass communications (http://www.jou.ufl.edu/commres/jouwww.htm).

• Journalism: Journalism List of Internet Resources, by John S. Makulowich (ftp://rtfm.mit.edu/pub/usenet/news.answers/journalism-net-resources).

• Journalism Directory: Internet Resources, by John S. Makulowich (ftp://ftp.clark.net/pub/journalism/).

• Media BBS: Media BBS List, broadcasting/journalism BBS list, by Mark S. Leff (mailto:listserv@ulkyvm.louisville.edu Body: GET MEDIA BBSLIST).

• Media List ftp: e-mail addresses of media outlets, by Adam Gaffin (ftp://ftp.std.com/customers/periodicals/Middlesex-News/medialist).

• Media List list: e-mail addresses of media outlets, by Adam Gaffin (mailto:majordomo@world.std.com Body: subscribe medialist).

• Movies and TV: Nova’s Movies and TV information (http://alpha.acast.nova.edu/movies.html).

• Online newspapers: a service of MediaInfo Interactive(TM)

(http://marketplace.com/e-papers.list.www/e-papers.home.page.html).

• Periodicals: electronic periodicals and journals (gopher://gopher.cic.net/11/e-serials).

• SPJ: Society of Professional Journalists (ftp://ftp.netcom.com/pub/spj/html/spj.html).

• TV-film-video: Vortex’s collection of materials, including Cable Regulation Digest (gopher://vortex.com/11/tv-film-video).

• TV Networks-Yahoo: Business-Corporations-Media-Television-Networks, entry from Yahoo (http://www.yahoo.com/Business/Corporations/Media/Television/Networks/).

• TV Schedules: the schedules of some television channels/shows (http://white.nosc.mil/television.html).

• Student Newspapers: high school and college newspapers online (gopher://blick.journ.latech.edu).

• Tools: a list of Internet tools for mass communication

(http://www.rpi.edu/Internet/Guides/decemj/itools/cmc-mass.html). —– Application Areas • Amateur Radio: the World of Internet Amateur (Ham) Radio

(http://www.acs.ncsu.edu/HamRadio/).

• Cable TV: Cable Television Laboratories, Inc. (http://www.cablelabs.com).

• Intelsat: the leading provider of satellite telecommunications services (http://www.intelsat.int:8080).

• Live TV: demonstration consists of live video and audio using the vsdemo VuSystem application that uses a live television source on the VuNet (http://tns-www.lcs.mit.edu/cgi-bin/vs/vsdemo).

• Satellite: The Satellite TV Page (http://itre.uncecs.edu/misc/sat.html). —– Outlets

• AP: Associated Press (http://www.trib.com/news/AP.html).

• BBC Networking Club: (http://www.bbcnc.org.uk).

• C-SPAN: Cable-Satellite Public Affairs Network (gopher://c-span.org).

• CBC Radio: (http://radioworks.cbc.ca).

• CBS: (http://www.cbs.com).

• ClariNet: complete electronic newspaper delivered via Net (mailto:info@clarinet.com).

• gina: Internet news agency (http://www.earthlink.net/gina/).

• ITR: Internet Talk Radio (http://www.ncsa.uiuc.edu/radio/radio.html).

• Cyberspace Report: a public affairs radio show aired on KUCI, 88.9 FM in Irvine, California (http://www.ics.uci.edu/∼ejw/csr/cyber.html).

• PBS: United States Public Broadcasting Service (gopher://gopher.pbs.org).

• Palo Alto Weekly: (http://www.service.com/PAW/home.html).

• TV Net: Web Homes for Television and Cable Stations throughout the U.S. and the World (http://cinenet.net/TVnet.html).

• TV nets: Business-Corporations-Television-Networks, from Yahoo (http://www.yahoo.com/Business/Corporations/Television/Networks/).

• VOA: Voice of America (USA) (ftp://ftp.voa.gov).

• VOA gopher: Voice of America (USA) (gopher://gopher.voa.gov).

• Wired: (mailto:editor@wired.com).

3.2.5 Societal

• Cyberpunk: texts for understanding technology and culture, from the English Server (http://english-server.hss.cmu.edu/Cyber.html).

• Friends + Partners: USA and Russia information and communication (http://solar.rtd.utk.edu/friends/home.html).

• Net Revisited: The Network Nation Revisited, a bachelor’s thesis examining Hiltz and Turoff’s predictions for CMC on societal change, by David Belson (http://www.stevens-tech.edu/∼dbelson/thesis/thesis.html).

3.2.6 Surveillance / remote-control

• Anthony’s List: Internet-accessible machines for vending, temperature sensors, tele-operated machinery, equipment status, cameras/video, and others, by Anthony Anderberg (http://www.dsu.edu/∼anderbea/machines.html).

• bsy’s List: Internet Accessible (non-Coke) Machines, connections to many surveillance and remote sensing examples (http://www.cs.cmu.edu:8001/afs/cs.cmu.edu/user/bsy/www/iam.html).

• Camera in office: Spy on Dennis Gannon, Research Director of Center for Innovative Computer Applications (http://www.cica.indiana.edu/htbin/camera).

• Interesting Devices: Computers-Internet-Interesting Devices Connected to the Net, from Yahoo (http://www.yahoo.com/Computers/Internet/Interesting Devices Connected to the Net/).

• Robot: Robotic Tele-Excavation at the University of Southern California, Mercury Project, users tele-operate a robot arm moving over a terrain filled with buried artifacts (http://www.usc.edu/dept/raiders/).

• Temperature: Temperature on CU/Boulder Engineering Center Roof (http://www.cs.colorado.edu/htbin/temp).

3.2.7 Scientific

• CISTI: Canada Institute for Scientific and Technical Information–worldwide scientific, technical and medical information (http://www.cisti.nrc.ca/cisti/cisti.html).

• CAS: The Chicago Academy of Sciences (http://www.mcs.com/∼cas/home.html).

• CS Tech Reports: Computer Science Technical Reports Archive Sites (http://www.rdt.monash.edu.au/tr/siteslist.html).

• DoD IAC: Department of Defense (DoD) Information Analysis Center (IAC) Hub Page (http://www.dtic.dla.mil/iac/).

• EnviroWeb: environmental information source on the Internet (http://envirolink.org).

• NAS ftp: National Academy of Sciences (ftp://ftp.nas.edu).

• NAS web: National Academy of Sciences (http://www.nas.edu).

• NSF-USA: The National Science Foundation (NSF), United States (http://www.nsf.gov).

• Science-Yahoo: (http://www.yahoo.com/Science/).

• SETN: The Science and Engineering Television Network, Inc., a non-profit consortium of scientific societies, universities, laboratories and corporations organized to foster the development of scientific communication in the medium of television (http://www.service.com/stv/setncall.html).

• SPIE Web: services from The International Society for Optical Engineering (SPIE)

(http://www.spie.org).

• WATERS: Wide Area Technical Report Service, a WAIS-based technical reports repository (http://www.cs.odu.edu/WATERS/WATERS-GS.html).

3.3 Education

—– Major Tree Collections

• Education-Galaxy: EINet Galaxy’s collection of resources about education (http://galaxy.einet.net/galaxy/Social-Sciences/Education.html).

• Education-WWW VL: WWW Virtual Library entry for education (http://www.w3.org/hypertext/DataSources/bySubject/Education/Overview.html).

• Education-Yahoo: Education section from Yahoo (http://www.yahoo.com/Education/). —– Other Collections

• AskERIC Web: Educational Resources Information Center, Clearinghouse on Information and Technology, federally-funded collection of education-related resources (http://eryx.syr.edu/Main.html).

• Best for ed: a selection of the best of the Internet for educators (gopher://info.asu.edu/11/asu-cwis/education/other).

• BBN NSN: BBN’s National School Network Testbed (gopher://copernicus.bbn.com).

• CALICO: Computer Assisted Language Instruction Consortium (mailto:CALICO@Dukemvs.ac.duke.edu).

• CAUSE gopher: association for managing and using information technology in higher education (gopher://cause-gopher.Colorado.edu).

• Center/Excellence: Center for Excellence in Education, to help our best students and teachers keep the United States competitive in science and technology, and to nurture international understanding among potential leaders of many countries (http://rsi.cee.org).

• Chronicle web: ACADEME THIS WEEK, offers guide to contents of The Chronicle of Higher Education and all Positions Available ads, calendar of events, deadlines for grants, papers, fellowships, put out every Tuesday at noon EST (http://chronicle.merit.edu).

• Cisco: education archive, lists a variey of resources about school connectivity to the Internet as well as resources on the WWW (http://sunsite.unc.edu/cisco/edu-arch.html).

• COL: Commonwealth of Learning, distance education techniques and associated communications technologies (http://www.col.org).

• CoVis: Learning Through Collaborative Visualization, Northwestern University (http://www.covis.nwu.edu).

• Cyberion City: virtual environment built by schoolchildren (telnet://guest@michael.ai.mit.edu).

• CTC: Cornell Theory Center Math and Science Gateway, resources in mathematics and science for educators and students in grades 9-12 (http://www.tc.cornell.edu/Edu/MathSciGateway/index.html).

• DeweyWeb: facilitating communication between students from all over the world (http://ics.soe.umich.edu).

• Distance Ed: (ftp://una.hh.lib.umich.edu/inetdirsstacks/disted:ellsworth).

• Distance Learning: directory of information about Distance Learning (http://www.mitn.msu.edu/distance.htm).

• Dr. E’s Compendium: Dr. E’s Eclectic Compendium of Electronic Resources for Adult/Distance Education, by J. H. Ellsworth (ftp://ftp.std.com/pub/je/dre-list.txt).

• NDLC: National (USA) Distance Learning Center (telnet://ndlc@ndlc.occ.uky.edu).

• Distance Ed DB: A database from the International Centre for Distance Learning at the UK Open University (mailto:n.ismail@open.ac.uk).

• Distance Ed DB: contains full text of Mindweave , edited by Mason and Kaye (telnet://icdl@acsvax.open.ac.uk).

• Diversity U: Diversity University, an experiment in interactive learning through internet (telnet://moo.du.org:8888).

• EdWeb-CPB: Corporation for Public Broadcasting, an on-line tutorial on education, technology, school reform and the Information Highway (http://k12.cnidr.org:90).

• Education Gopher: Florida Tech’s collection of educational resources (gopher://sci-ed.fit.edu).

• EOS web: Educational Online Sources, pointers for education (http://www.netspace.org/eos/).

• Educational Technology: The World-Wide Web Virtual Library–Educational Technology (http://tecfa.unige.ch/info-edu-comp.html).

• Education Technology Initiatives: Higher Education Funding Councils for England, Wales, Scotland and Northern Ireland now support three major initiatives in the UK which are aimed at improving universities’ awareness and use of technological innovations in teaching and learning (gopher://gopher.csv.warwick.ac.uk/11/remote/other-remote/edu-tech).

• Educator’s Email: An Educator’s Guide to E-Mail Lists (ftp://nic.umass.edu/pub/ednet/educatrs.lst).

• Educator’s USENET: Ednet Guide to Usenet Newsgroups (ftp://nic.umass.edu/pub/ednet/edusenet.gde).

• Empire Schoolhouse: Empire Internet Schoolhouse, a selection of K-12 resources, projects and discussion groups (gopher://nysernet.org:3000/11/).

• EnviroWeb: a project of the EnviroLink Network, sponsors many on-line environmental education initiatives (http://envirolink.org).

• ETB/NLM: Educational Technology Branch (ETB), part of the Lister Hill National Center for Biomedical Communications (LHNCBC) at the National Library of Medicine (NLM) (http://wwwetb.nlm.nih.gov).

• Exploratorium: a collage of 650 interactive exhibits in the areas of science, art, and human perception (http://www.exploratorium.edu).

• GENII: Group Exploring the National Information Infrastructure (http://www.deakin.edu.au/edu/MSEE/GENII/GENII-Home-Page.html).

• Global SchoolNet Foundation: a non-profit organization dedicated to the instructional applications of telecommunication (http://gsn.org).

• GNA: Globewide Network Academy, a non-profit corporation, affiliated with Usenet University; goal is to create an fully accredited online university (http://uu-nna.mit.edu:8001/uu-gna/index.html).

• High school/Internet: a case study of an east central Florida high school, by Michael Anthony Gallos to the internet (ftp://sci-ed.fit.edu/pub/Internet/study/).

• HUB: mathematics and science education (http://hub.terc.edu).

• IAT archive: Institute for Academic technology ftp site (ftp://gandalf.iat.unc.edu/user/home/anonftp/guides/).

• iCDL: International Centre for Distance Learning, in the United Kingdom on the campus of the Open University, disseminating distance education information worldwide (http://acs-info.open.ac.uk/info/other/ICDL/ICDL-Facts.html).

• IKE: IBM Kiosk for Education (gopher://ike.engr.washington.edu).

• IKE Web: IBM Kiosk for Education (http://ike.engr.washington.edu/ike.html).

• ILT: Institute for Learning Technologies as part of the Columbia University Virtual Information Initiative (http://www.ilt.tc.columbia.edu).

• ILC-Southampton: Interactive Learning Centre, University of Southampton, United Kingdom (http://ilc.ecs.soton.ac.uk/welcome.html).

• Internet and Ed: by Noel Estabrook (ftp://ftp.msu.edu/pub/education/).

• IRD/Educators: Internet Resource Directory for Educators (ftp://tcet.unt.edu/pub/telecomputing-info/IRD).

• IT Connections: Instructional Technology Connections, by Martin Ryder (gopher://ccnucd.denver.colorado.edu/h0/UCD/dept/edu/IT/itcon.html).

• JASON: introduce teachers to the world of telecommunicating (telnet://topcat.bsc.mass.edu).

• JASON Project web: A journey to the rain forest, caverns, Mayan ruins and coral reef of Belize (http://seawifs.gsfc.nasa.gov/JASON/JASON.html).

• Jones Ed Nets: Jones Education Networks (Mind Extension University, Jones Computer Network) (http://www.meu.edu).

• Kenyon workbook: Workbook for Kenyon’s Summer Institute on Academic Information Resources, emphasis on faculty and student exploration of information resources, communication, and collaboration, from Kenyon College, Gambier, Ohio (ftp://ftp.kenyon.edu/pub/pub/e-pubs/workbook/).

• KidLink: gopher aimed at 10-15 year olds (gopher://kids.ccit.duq.edu).

• K-12 List: list of K-12 Internet School Sites, by Gleason Sackman (http://toons.cc.ndsu.nodak.edu/∼sackmann/k12.html).

• K-12 Registry: WWW Schools Registry, from Web66/Hillside Elementary School (http://hillside.coled.umn.edu/others.html).

• K-12/Armadillo: annotated directory of K-12 resources from Rice University (http://chico.rice.edu/armadillo/Rice/K12resources.html).

• K-12 Gopher: CICnet’s K-12 Internet gopher (gopher://gopher.cic.net/11/cicnet-gophers/k12-gopher).

• K-12 Info/CNIDR: includes global schoolhouse project and Janice’s K12 Cyberspace Outpost (http://k12.cnidr.org).

• K-12 Briarwood: collection of K-12 resources of the Internet, including a Curriculum Database (http://www.briarwood.com).

• K-12 NASA/Langley: NASA Langley Research Center’s HPCC K-12 Program (http://k12mac.larc.nasa.gov/hpcck12home.html).

• K-12 NASA/NAS: pointers to online Educational Resources, by Chris Beaumont (http://www.nas.nasa.gov/HPCC/K12/edures.html).

• K-12 Web 66: resources and information to facilitate the introduction of Web technology into K12 schools (http://web66.coled.umn.edu).

• Latitude 28: Latitude28 Schoolhouse, a privately sponsored project, designed to make educational materials accessible to students of all ages (http://www.packet.net/schoolhouse/Welcome.html).

• Learning V: Learning Village, Internet access to schools, by National Public Telecomputing Network and Ameritech (mailto:tmg@nptn.org).

• LRDC: Learning Research and Development Center at the University of Pittsburgh–thinking, knowing, and understanding in and beyond school (http://www.lrdc.pitt.edu).

• NCSA education: National Center for Supercomputing Applications Education Program Home Page (http://www.ncsa.uiuc.edu/Edu/EduHome.html).

• Maricopa: Center for Learning and Instruction, Maricopa Community Colleges, Arizona (http://www.mcli.dist.maricopa.edu).

• MEU BBS: Mind Extension University Bulletin Board System (telnet://bbs.meu.edu).

• Networking: Networking on the Network, an essay by Phil Agre (mailto:comserve@vm.its.rpi.edu Body: send Profess Network).

• Networking: Networking on the Network, an essay by Phil Agre (http://communication.ucsd.edu/pagre/network.html).

• NCET: National Council for Educational Technology Information Service (UK)

(http://datasun.ncet.org.uk).

• OISE Gopher: Ontario Institute for Studies in Education (gopher://porpoise.oise.on.ca:70).

• Online LC: Mount Allison University’s Online Learning Centre–tele-education (gopher://pringle.mta.ca).

• Peterson’s: Educational Center for education and lifelong learning (http://http://www.petersons.com/).

• Plugged In: a non-profit group dedicated to bringing the educational opportunities created by new technologies to children and families from low-income communities, located in East Palo Alto, California (http://www.pluggedin.org).

• Primary/Sec: Answers to Commonly Asked Primary and Secondary School Internet User Questions (ftp://nic.merit.edu/documents/fyi/fyi 22.txt).

• Singapore-Ed: Ministry of Education (MOE), Singapore (http://www.moe.ac.sg).

• Teacher Education: Society for Technology and Teacher Education (STATE), the University of Virginia, and the University of Houston have collaborated to establish a Teacher Education Server on the Internet (http://curry.edschool.virginia.edu/teis/).

• TECFA: an academic team active in the field of educational technology at the School of Education and Psychology of the University of Geneva (http://tecfa.unige.ch/tecfa.html).

• US Dept of Ed gopher: United States Department of Education (gopher://gopher.ed.gov).

• US Dept of Ed Web: United States Department of Education (http://www.ed.gov).

• USENET University: an online a society of people interested in learning, teaching or tutoring (ftp://nic.funet.fi/pub/doc/uu/FAQ).

• VOU: Virtual Online University, an online liberal arts university online, using a virtual environment called a VEE, Virtual Education Environment (http://www.drag.net/∼VOU/).

• WWW teaching: WWW for instructional use, a collection of course materials, from the University of Texas at Austin (http://www.utexas.edu/world/instruction/index.html).

3.4 Entertainment

—– Major Collections

• Entertainment-Yahoo: Entertainment section from Yahoo (http://www.yahoo.com/Entertainment/).

—– Selected Instances

• Coke Machines: bsy’s List of Internet Accessible Coke Machines (http://www.cs.cmu.edu:8001/afs/cs.cmu.edu/user/bsy/www/coke.html).

• The Dominion: Sci-Fi Channel (http://www.scifi.com).

• Games: Games FAQS and other entertainment resources (http://wcl-rs.bham.ac.uk/GamesDomain).

• Interactive games: Zarf’s List of Interactive Games on the Web (http://www.cs.cmu.edu:8001/afs/cs.cmu.edu/user/zarf/www/games.html).

• WWW/Sports: World Wide Web of Sports, Spanning the globe to bring you a constant variety of sports information (http://tns-www.lcs.mit.edu/cgi-bin/sports).

3.5 Government

3.5.1 Initiatives for information infrastructures

These government-sponsored initiatives and organizations to implement or support national or state information infrastructures.

• Canada: Industry Canada Ministry, Communications Development Directorate (http://debra.dgbt.doc.ca/.dtp/dtp.html).

• Europe: European Council report on specific measures to be taken by the European Community and the Members States for information infrastructures (http://www.earn.net/EC/bangemann.html).

• Europe/Info Market: I’M (Info Market)- Europe Home Page, information about Europe and the European electronic information market (http://www.echo.lu).

• France: Reseau National de Telecommunications pour la Technologie, l’Enseignement et la Recherche (http://web.urec.fr/docs/renater/renater.html).

• GII: Agenda for Cooperation, by Al Gore and Ron Brown (http://ntiaunix1.ntia.doc.gov:70/0/papers/documents/giiagend.html).

• GII Info: Global Information Infrastructure information from the Computer Professionals for Social Responsibility (http://cpsr.org/dox/global.html).

• Govt Policy: Government Policy and the Information Superhighway, a survey of information about many government initiatives, by the National Library of Australia (http://www.nla.gov.au/lis/govnii.html).

• Japan-MITI: Japan’s Ministry of International Trade and Industry, program for advanced information infrastructure (http://www.miti.go.jp/index-e.html).

• Japan-MPT: Reforms toward the Intellectually Creative Society of the 21st Century (http://www.glocom.ac.jp/WhatsNew/MPT.html).

• Singapore-IT200: Information from the National Computing Board (NCB) of Singapore’s Digital Media Centre (http://www.ncb.gov.sg).

• Singapore-ITI: Information Technology Institute (ITI), the applied R and D arm of the National Computer Board (NCB) of Singapore (http://www.iti.gov.sg).

• Singapore-NII SCAN: a publication of major NII policy trends around the world (http://www.ncb.gov.sg/nii/contents.html).

• USA-HPCC Web: the National Coordination Office (NCO) for High Performance Computing and Communications (http://www.hpcc.gov).

• USA-HPCC Gopher: the National Coordination Office (NCO) for High Performance Computing and Communications (gopher://gopher.hpcc.gov:70).

• USA-Info highway: pointers to information about efforts to create information networks at government and private levels (http://ai.iit.nrc.ca/superhighway.html).

• USA-NII FTP: National Information Infrastructure (ftp://ftp.ntia.doc.gov/pub/niiagenda.asc).

• USA-NII BBS: The White House Information Infrastructure Task Force BBS (telnet://gopher@iitf.doc.gov).

• USA-NII Gopher: The White House Information Infrastructure Task Force BBS (gopher://iitf.doc.gov).

• USA-NII Dialup: The White House Information Infrastructure Task Force BBS (phone:1-202-501-1920).

• USA-NII DIIG: The Digital Information Infrastructure Guide (DIIG) is a resource to facilitate the development of the National Information Infrastructure (NII) (http://farnsworth.mit.edu/diig.html).

• USA-NII Testbed: National Information Infrastructure Testbed (http://www.esi.com/niit top.html).

• USA-NII Web: The White House Information Infrastructure Task Force Web (http://iitf.doc.gov).

• USA-NII Related: Related Efforts in the National Information Infrastructure (http://www.acl.lanl.gov/sunrise/RelatedInfo/OtherProjects.html).

• USA-NII-XIWT: USA’s National Information Infrastructure Cross-Industry Working Team, a coalition of industry to define requirements for a national information infrastructure (http://www.cnri.reston.va.us:3000/XIWT/public.html).

• USA-North Carolina: North Carolina Information Highway (NCIH) home page (http://www.ncih.net).

3.5.2 Information supplied by governments

Government-sponsored or government-related information, particularly dealing with CMC issues. —— Major Collections

• Government-Yahoo: a collection of links about government information, from Yahoo (http://www.yahoo.com/Government/).

—— Countries

• Canada: Open Government – information about Canadian Federal government (http://debra.dgbt.doc.ca/opengov).

• Federal Info Resources: documents from the Information Policy Branch, Office of Information and Regulatory Affairs, Office of Management and Budget of the U.S. government (ftp://nic.merit.edu/omb/INDEX.omb).

• Federal Govt: Federal Government Information on the Internet, by Maggie Parhamovich, Head, Government Publications, Dickinson Library, University of Nevada Las Vegas (gopher://una.hh.lib.umich.edu/00/inetdirsstacks/usfedgov%3aparhamovich).

• Fedix: Federal Information Exchange, Inc., access to a wide variety of government data (http://www.fie.com).

• Fedworld: a system for government information (telnet://fedworld.doc.gov).

• Fedworld web: a gateway to U.S. Government information operated by the Department of Commerce’s National Technical Information Service (NTIS) (http://www.fedworld.gov).

• Government-Galaxy: page from EINet’s Galaxy showing a large collection of government information (http://www.einet.net/galaxy/Government.html).

• GILS: Government Information Locator Service (ftp://ftp.cni.org/pub/gils/).

• Govt/Citizenship: information about federal agencies, guides, policy, and other government information (gopher://eryx.syr.edu).

• Govt Information: Internet Sources of Government Information by Blake Gumprecht (ftp://ftp.nwnet.net/user-docs/government/gumprecht-guide.txt).

• Japan PM: Japan Prime Minister’s official residence Web server (http://www.kantei.go.jp).

• UK CCTA: United Kingdom Government Centre for Information Systems (http://www.open.gov.uk).

• UK HM Treasury: United Kingdom Her Majesty’s Treasury (http://www.hm-treasury.gov.uk).

• US Govt Gophers: a collection of government gophers (gopher://peg.cwis.uci.edu:7000/11/gopher.welcome/peg/GOPHERS/gov).

• US Govt Webs: List of WWW Servers (USA - Federal Government), from Federal Information Exchange (http://www.fie.com/www/us gov.htm).

• US NPR: US National Performance Review (http://www.npr.gov).

• US Patents and Trademarks: includes Intellectual Property and the National Information Infrastructure issues (http://www.uspto.gov).

• US House: US House of Representatives gopher (gopher://gopher.house.gov:70).

• US House: US House of Representatives (http://www.house.gov).

• US Senate: (gopher://gopher.senate.gov:70).

• US Thomas: US Legislative Information (http://thomas.loc.gov).

• US White House: (http://www.whitehouse.gov).

3.5.3 Policy

These pointers relate to public policy issues or government-citizen interaction, or legal issues. • ACE: Americans Communicating Electronically (gopher://cyfer.esusda.gov/11/ace).

• CapWeb/Policy.Net: Political issues, campaigns, publications, and a guide to the United States Congress (CapWeb) (http://policy.net).

• Computer Tech Legal Net: Computer Technology Legal information resources (http://www.com/mmm/mmmhome.html).

• ICTSP: The George Washington University School of Engineering and Applied Science Institute for Computer and Telecommunications Systems Policy (http://www.seas.gwu.edu/seas/instctsp/ictsp.html).

• Information Policy: Information Policy for Electronic Information Resources, by Mary Lou Goodyear (mailto:listserv@uhupvm1.uh.edu Body: GET GOODYEAR PRV4N6).

• Intellectual Property: World Intellectual Property Organization (WIPO), includes Working Group on Intellectual Property Rights, a subgroup of the Information Infrastructure Task Force (http://www.uspto.gov/wipo.html).

• International Trade: network resources for legal research and education (http://ananse.irv.uit.no/trade law/misc/objective.html).

• Internet Economics Collection: a collection of documents that have to do with the economics of the Internet, information goods, and related issues, by Hal Varian (http://gopher.econ.lsa.umich.edu/EconInternet.html).

• Internet Policy: What Should We Plan Given the Dilemma of the Network?, by G. Cook (ftp://nic.merit.edu/documents/rfc/rfc1527.txt).

• MIT-RPCP: MIT’s Research Program on Communications Policy (http://farnsworth.mit.edu).

• Open Platform: Public Policy for the Information Age, EFF (ftp://ftp.eff.org/pub/EFF/papers/Open Platform/).

• Politics: a collection of etexts on miscellaneous political topics (gopher://fir.cic.net/11/Politics).

• PPP: Political Participation Project, a research project at the MIT Artificial Intelligence Lab designed to explore how interactive media can be used to facilitate political participation; web includes collection of political sources, a bibliography, directory of grassroots organizations, and publications (http://www.ai.mit.edu/projects/ppp/home.html).

• Tap Info: reports on activities relating to federal information policy (ftp://ftp.cpsr.org/taxpayer assets).

• Telecom legislation: (ftp://ftp.govt.washington.edu/legislation.telecom/).

• Telecom policy: (http://www.ba.com/sites.html).

• Voters Telecom Watch: monitoring important bills and alerting the public at crucial times in the life of a bill (gopher://gopher.panix.com/11/vtw/vtwinfo).

• Warren Gopher: a collection of electronic newsletters distributed by Jim Warren on the subjects of political action and government access through the use of computer communications (gopher://gopher.path.net:8102/1).

3.6 Information

3.6.1 Dissemination and Retrieval

• ACM/SIGIR: Association of Computing Machinery (ACM) Special Interest Group on Information Retrieval (http://info.sigir.acm.org/sigir/).

• Agents, Intelligent: a collection of information about software information agents (http://retriever.cs.umbc.edu/agents/).

• Addressing: Uniform Resource Identifiers/Locators/Names (URI, URL, URN)

(http://www.w3.org/hypertext/WWW/Addressing/Addressing.html).

• ArD: database of published articles in the field of information systems (http://www.zib-berlin.de/People/mueller/GNA/ISG/Services/ArD/).

• ALMANAC: A service for multi-media document and information delivery and database functions (mailto:almanac@oes.orst.edu Body: send guide).

• Bibl-Mode: an Emacs Lisp package to keep track of information on the net (ftp://ftp.maths.tcd.ie/pub/bosullvn/elisp/).

• CAIT-WUSTL: Center for the Application of Information Technology, part of the School of Engineering and Applied Science at Washington University in St. Louis (http://www.cait.wustl.edu/cait/intro.html).

• Carl/Uncover: tables of contents and article level access to over 16,000 unique multidisciplinary journals, reflecting the collections of selected CARL System and other libraries (telnet://database.carl.org).

• Cataloging Internet Res: explores how to catalog electronic information, by Priscilla Caplan (mailto:listserv@uhupvm1.uh.edu Body: get caplan prv4n2).

• CATRIONA: CATaloguing and Retrieval of Information Over Networks Applications (http://www.bubl.bath.ac.uk/BUBL/catriona.html).

• CGI: Common Gateway Interface, a standard for external gateway programs to interface with information servers such as HTTP servers (http://hoohoo.ncsa.uiuc.edu/cgi/overview.html).

• CHAT: Conversational Hypertext Access Technology, natural-language interface (ftp://debra.dgbt.doc.ca/pub/chat/).

• CHAT: natural language interface to information (telnet://debra.doc.ca).

• CIT: Center for Information Technology, a laboratory operated by Stanford University, the encoding, storage, communication, manipulation, and use of information in digital form (http://logic.stanford.edu/cit/cit.html).

• CIIR-UMass: Center for Intelligent Information Retrieval, University of Massachusetts, Amherst (http://ciir.cs.umass.edu).

• CNI FTP: Coalition for Networked Information (ftp://ftp.cni.org/CNI/).

• CNI Gopher: Coalition for Networked Information, promotes creation of and access to information resources in networked environments (gopher://gopher.cni.org).

• CNI Web: (http://www.cni.org/CNI.homepage.html).

• CNI search: Coalition for Networked Information search server (telnet://brsuser@gopher.cni.org).

• CNI TopNode Project: part of CNI’s Directories and Resource Information Services Working Group (ftp://ftp.cni.org/CNI/projects/topnode/).

• CWI: Centrum voor Wiskunde en Informatica, Centre for Mathematics and Computer Science, Amsterdam, Netherlands (http://www.cwi.nl/default.html).

• CNIDR Web Page: Coalition for Networked Information Discovery and Retrieval Home Page (http://cnidr.org/welcome.html).

• CyberWeb: useful information for information providers (http://www.charm.net/∼web/).

• Data Research: databases and information, some free (don’t abuse free services)

(http://www.dra.com).

• DIMUND: Document Understanding Information and Resources (gopher://dimund.umd.edu).

• DIMUND FTP: Document Understanding FTP Archives (ftp://dimund.umd.edu/pub/).

• Doc Center: Document Center, a hard copy document delivery service for government and industry specifications and standards (http://doccenter.com/doccenter/home.html).

• Electronic pub. info: Electronic Publishing information from the U.S. Library of Congress (gopher://marvel.loc.gov/11/research/e).

• Electronic pub. samples: samples of electronic publishing (http://www.elpress.com/samples/samples.html).

• IAFA: Internet Anonymous FTP Archives Working Group (ftp://archive.cc.mcgill.ca/pub/Network/iafa/charter).

• Info Provider: discussion of WWW and other servers from an information provider standpoint (http://www.w3.org/hypertext/WWW/Provider/Overview.html).

• Integrated Information: A Vision of an Integrated Internet Information Service, by Weider and Deutsch (ftp://venera.isi.edu/internet-drafts/draft-ietf-iiir-vision-01.txt).

• IIP: Intelligent Information Infrastructure Project, MIT Artificial Intelligence Laboratory (http://www.ai.mit.edu/projects/iiip/home-page.html).

• Interpedia: mission is to be a primary source of information for Internet users, and a guide to many of the online resources available (http://www.hmc.edu/www/interpedia/index.html).

• IRLP: Internet Resource Location Project (http://www.cs.colorado.edu/home/gc/cs/genbbb wwww.html).

• IRIS-Brown: Brown University’s Institute for Information Research and Scholarship (http://www.iris.brown.edu/iris).

• IRTF: Internet Research Task Force Research Group on Resource Discovery (IRTF-RD) (http://rd.cs.colorado.edu/∼schwartz/IRTF.html).

• ISI-USC: University of Southern California (USC) Information Sciences Institute (ISI) (http://www.isi.edu).

• ISRI-UNLV: Information Science Research Institute, University of Nevada, Las Vegas (http://www.isri.unlv.edu).

• ISWorld Net: information systems web, sponsored by the Edwin L. Cox School Management at Southern Methodist University (http://www.cox.smu.edu/mis/iswnet/home.html).

• Library Special: Library special collections on the internet (ftp://vm1.nodak.edu/nnews/nnews.1993-10).

• Lycos: a research program in providing information retrieval and discovery in the WWW, using a finite memory model of the web to guide intelligent, directed searches for specific information needs (http://lycos.cs.cmu.edu).

• Mailbase: an enhanced electronic mailing list service (http://mailbase.ac.uk/welcome.html).

• NIR Archives: Networked Information Retrieval (ftp://mailbase.ac.uk/pub/lists/nir/).

• NIR Gopher: Networked Information Retrieval Gopher (gopher://mailbase.ac.uk).

• OCLC Research: Toward Providing Library Services for CMC (ftp://ftp.rsch.oclc.org/pub/internet resources project/report/).

• OSS: Open Source Solutions, open sources, methods, and products; competitive in the Age of Information (gopher://gopher.oss.net).

• PCP: Principia Cybernetica Project (PCP)–the computer-supported collaborative development of an evolutionary-systemic philosophy (http://pespmc1.vub.ac.be/RELATED.html).

• Publishing: Advanced Publishing in the Web, by Terje Norderhaug (http://www.ifi.uio.no/∼terjen/WWWauthoring/abstract.html).

• Quality: Quality and Standards of the Internet Information Resources (http://coombs.anu.edu.au/SpecialProj/QLTY/QltyHome.html).

• ResInfo: working toward a basic operational framework necessary to implement a globally distributed information system (http://www.hmc.edu/∼jared/professional/resinfo/resinfo.html).

• Resource Transponders: by Weider and Deutsch (ftp://venera.isi.edu/internet-drafts/draft-ietf-iiir-transponders-01.txt).

• Retrieval success: successful stories of using the Internet for reference, by Karen Schneider (ftp://mailbase.ac.uk/pub/lists/unite/files/internet-stories.txt).

• Riddle: Rapid Information Display and Dissemination in a Library Environment (ftp://ftp.cwi.nl/pub/RIDDLE/).

• SIFTER-UAB: Department of Computer and Information Sciences at the University of Alabama at Birmingham (UAB), Studies of Information Filtering Technology for Electronic Resources (SIFTER) Research Group (http://twinbrook.cis.uab.edu:70/AdPage.80).

• Scout project: approximate matching and probabilistic retrieval for multimedia and hypermedia applications (http://fuzine.mt.cs.cmu.edu/scout/home.html).

• SGML Open: a non-profit, international consortium of providers of products and services, dedicated to accelerating the further adoption, application, and implementation of the Standard Generalized Markup Language (http://www.sgmlopen.org).

• SGML Review: A biased review of SGML, by Tim Berners-Lee (http://www.w3.org/hypertext/WWW/MarkUp/SGML.html).

• SIGNIDR: Special Interest Group on Networked Information Discovery and Retrieval (http://www.wais.com/SIGNIDR/).

• System comparison: comparing issues in world-wide networked hypertext systems, by Arthur Smith (ftp://snorri.chem.washington.edu/comparison table).

• UNITE Archive: User Network Interface To Everything (ftp://mailbase.ac.uk/pub/lists/unite/).

• UWI: UnderWorld Industries, nodes that wish to share information with other UWI nodes (http://zapruder.pds.med.umich.edu/uwi/uwi-info.html).

• Web Publishing: The Web Communications Comprehensive Guide to Publishing on the Web (http://www.webcom.com/html/).

• Z39.50: Network Information Dissemination Standards (http://www.research.att.com/∼wald/z3950.html).

3.6.2 Library

• Access: catalogs and databases (ftp://ftp.unt.edu/library/libraries.txt).

• ALA: American Library Association (gopher://gopher.uic.edu/11/library/ala/).

• ALAWON archive: files from American Library Association (mailto:listserv@uicvm.uic.edu Body: send ala-wo filelist).

• ALIX: Automated Library Information Xchange; Advice, opinion, and software by and for librarians; a service of the Federal Library and Information Center Committee, Federal Library Network (telnet://alix.loc.gov:3001).

• ARL Web: Association of Research Libraries Web server (http://arl.cni.org).

• BUBL gopher: The Bulletin Board for Libraries gopher (gopher://bubl.bath.ac.uk).

• BUBL Web: The Bulletin Board for Libraries Web server (http://bubl.bath.ac.uk/BUBLHOME.html).

• CARL: access to UnCover and other indexes (http://www.carl.org/carl.html).

• Dialog: a service of Knight-Ridder Information, Inc. (http://www.dialog.com).

• Internet libraries: various internet material for librarians (ftp://nic.funet.fi/pub/doc/library/).

• IPL: Internet Public Library, information and services to enhance the value of the Internet to its users (http://ipl.sils.umich.edu).

• LaUNCpad: access various infomration services, from the University of North Carolina (telnet://launch@launchpad.unc.edu).

• LIS: Library and Information Science resources, by Eric Lease Morgan (http://www.lib.ncsu.edu/disciplines/library.html).

• LIS-Galaxy: Library and Information Science listing from Galaxy (http://www.einet.net/galaxy/Social-Sciences/Library-and-Information-Science.html).

• Library Resources: Library Resources on the Internet–Strategies for Selection and use, Ed. Laine Farley (ftp://dla.ucop.edu/pub/internet/libcat-guide).

• Lib resources/Northwestern: Library resources on the Internet, from Northwestern University (http://www.library.nwu.edu/DOCS/LibResources.html).

• Lib webs: Library Information Servers via WWW, by Thomas Dowling (http://www.lib.washington.edu/∼tdowling/libweb.html).

• LEXIS-NEXIS: Legal and news information (http://www.meaddata.com).

• OCLC: Online Computer Library Center, a nonprofit computer library service and research organization (http://www.oclc.org).

• Public Libraries: St. Joseph County (Indiana, USA) list of public libraries with Internet services (http://sjcpl.lib.in.us/homepage/PublicLibraries/PublicLibraryServers.html).

• RLG: Research Libraries Group, Inc., a not-for-profit membership corporation of universities, archives, historical societies, museums, and other institutions devoted to improving access to information that supports research and learning; includes access to Eureka (search system), RLIN (library support system), and Zephyr (Z39.50 services), database, citation files, Ariel document delivery (http://www-rlg.stanford.edu/welcome.html).

• US LOC Marvel: United States Library of Congress, Marvel (Machine-Assisted Realization of the Virtual Electronic Library) (telnet://marvel@marvel.loc.gov).

• US LOC telnet: United States Library of Congress telnet access (telnet://locis.loc.gov).

• US LOC web: United States Library of Congress web (http://lcweb.loc.gov/homepage/lchp.html).

• WWW/Mosaic: The World-Wide Web and Mosaic-An Overview for Librarians, by Eric Lease Morgan (http://www.lib.ncsu.edu/staff/morgan/www-and-libraries.html).

3.7 Multiple

• 101 stories: Reprinted from 101 Success Stories of Information Technology in Higher Education; The Joe Wyatt Challenge, Edited by Judith Boettcher. A Project of EDUCOM’s Educational Uses of Information Technology (EUIT) (gopher://ivory.educom.edu/11/stories.101).

• FARNET stories: Federation of American Research Networks (FARNet) descriptions of 51 Reasons for a National Information Infrastructure (NII) (gopher://gopher.cni.org/11/cniftp/miscdocs/farnet).

• Success stories: Internet Reference Success Stories, collected by Karen G. Schneider (http://www.intac.com/∼kgs/success/successcover.html).

3.8 Scholarship

• CIOS: Communication Institute for Online Scholarship, a non-profit organization for online communication scholarship; gopher server (gopher://cios.llc.rpi.edu).

• CNI: Coalition for Networked Information, a project of the Association of Research Libraries, CAUSE, and EDUCOM to promote the creation of and access to information resources in networked environments in order to enrich scholarship and to enhance intellectual productivity (http://www.cni.org/CNI.homepage.html).

• IATH: Institute for Advanced Technology in the Humanities at the University of Virginia in Charlottesville, an effort to bring Jefferson’s educational ideas of exchange across disciplines and integrated living and learning (http://jefferson.village.virginia.edu).

• IRIS-Brown: Brown University’s Institute for Information Research and Scholarship, explores ways technology can be used for research, teaching, and learning (http://www.iris.brown.edu/iris).

• Scholarly Communications Project: of University Libraries, Virginia Polytechnic Institute and State University, a project to pioneer electronic communication of scholarly materials (http://borg.lib.vt.edu).

• Scholarly Communication Reports: Quarterly Technical Reports (ftp://borg.lib.vt.edu/pub/vpiej-l/reports).

• Scholarly Societies gopher: University of Waterloo Library maintains links to gophers and other servers of scholarly societies (gopher://watserv2.uwaterloo.ca/11/servers/campus/scholars).

• Scholarly Societies web: University of Waterloo Library maintains links to gophers and other servers of scholarly societies (http://www.lib.uwaterloo.ca/society/overview.html).

• Scholarly Comm paper/Kahin: Scholarly Communication in the Network Environment Issues of Principle, Policy, and Practice, by Brian Kahin (ftp://ftp.cni.org/CNI/projects/Harvard.scp/kahin.txt).

• Scholarly Comm/Libraries: Published by The Association of Research Libraries for The Andrew W. Mellon Foundation (ftp://ftp.cni.org/ARL/mellon/).

• Scholarly Communication study: University Libraries and Scholarly Communication, A Study Prepared for The Andrew W. Mellon Foundation, by Anthony M. Cummings, Marcia L. Witte, William G. Bowen, Laura O. Lazarus, and Richard H. Ekman, November 1992 (http://www.lib.virginia.edu/mellon/mellon.html).

• Scholarly Publishing: Centre for Networked Access to Scholarly Information at Australian National University Library (http://info.anu.edu.au).

• Scholarly Publishing models: Model University Policy Regarding Faculty Publication in Scientific and Technical Scholarly Journals A Background Paper and Review of the Issues (ftp://sunsite.unc.edu/pub/docs/about-the-net/trln-copyright-paper).

4 TECHNOLOGY 36

4 TECHNOLOGY

4.1 Computing

• ACM Gopher: Association for Computing Machinery’s gopher information server (gopher://acm.org).

• ACM Web: Association for Computing Machinery (http://info.acm.org).

• CACS-U of S LA: Center for Advanced Computer Studies, University of Southwestern Louisiana (http://www.cacs.usl.edu/Departments/CACS/).

• CPU Info Center: central processing unit (CPU) information, includes press announcements, papers, machine information, by Tom Burd (http://www.ncsa.uiuc.edu/General/MetaCenter/MetaCenterHome.html).

• HPCC-NSE: High Performance Computing and Communication (USA) National Coordinating office, National Software Exchange (http://www.netlib.org/nse/home.html).

• HPC Archive: London and South East Centre for High Performance Computing archive on high performance computing, includes articles and facility to add articles (http://www.lpac.qmw.ac.uk/SEL-HPC/Articles/index.html).

• HPC-Southampton: University of Southampton High Performance Computing Centre (http://cs1.soton.ac.uk).

• HPCWire: The High-Performance Computing news and information service, covering workstations through supercomputers (telnet://hpcwire.ans.net).

• IEEE: Institute of Electrical and Electronics Engineers (http://www.ieee.org).

• Ohio Supercomputer: a state-funded computing resource, provides high-performance computing to scientists and engineers at Ohio colleges, universities, and companies (http://www.osc.edu/welcome.html).

• NCSA-USA: National (USA) Center for Supercomputing Applications (http://www.ncsa.uiuc.edu/General/NCSAHome.html).

• NMCCSE-USA: United States National MetaCenter for Computational Science and Engineering (http://www.ncsa.uiuc.edu/General/MetaCenter/MetaCenterHome.html).

• NPAC-Syracuse: Northeast Parallel Architectures Center Home Page, at Syracuse University, New York (http://minerva.npac.syr.edu/home.html).

• Pitt SCC: Pittsburgh Supercomputing Center home page (http://pscinfo.psc.edu).

• RICIS-Houston: Research Institute for Computing and Information Systems (http://rbse.jsc.nasa.gov).

• SDSC: San Diego Supercomputer Center (http://gopher.sdsc.edu/Home.html).

• SEI-CMU: Software Engineering Institute at Carnegie Mellon University (http://www.sei.cmu.edu).

4.2 Developing

• IEEE Tech: Portfolio of Emerging Technologies (mailto:info.new.tech@ieee.org Body: anything).

• NIMT-Ireland: National Institute for Management Technology, Ireland (http://www.nimt.rtc-cork.ie/nimt.htm).

• NIST-USA Web: National Institute of Standards and Technology WWW Home Page (http://www.nist.gov/welcome.html).

• NIST-USA Gopher: National Institute of Standards and Technology (USA) Gopher (gopher://gopher-server.nist.gov).

• NSF-USA Gopher: National Science Foundation (USA) Gopher (gopher://stis.nsf.gov:70/11).

• NTTC-USA: National Technology Transfer Center (http://iridium.nttc.edu/nttc.html).

4.3 Human Interaction

• ACM/SIGCHI: Association of Computing Machinery (ACM) Special Interest Group on Computers and Human Interaction (gopher://gopher.acm.org/11[the files.sig forums.sigchi]).

• HCIBIB: A mail-based retrieval system interface to a database related to Human-Computer Interaction (HCI). (mailto:hcibib@bellcore.com Body: query:).

• HCIBIB web: references to literature on human-computer interaction, including hypertext/hypermedia (http://www.tu-graz.ac.at/CHCIbib).

• HCI-Galaxy: Human-computer Interaction page from EINET Galaxy (http://galaxy.einet.net/galaxy/Engineering-and-Technology/Computer-Technology/Human-computer-Interaction.html).

• HCI ftp site: Human-Computer Interaction (HCI) bibliography repository (ftp://archive.cis.ohio-state.edu/pub/hcibib/).

• HCI Index/deGraaff: Human-Computer Interaction Index web (http://www.twi.tudelft.nl/Local/HCI/HCI-Index.html).

• HCI Launching Pad: Human-Computer Interaction resources and pointers, by Keith Instone (http://www.cs.bgsu.edu/HCI/).

• HCS: Center for Human-Computer studies, at Uppsala University, Sweden (http://www.cmd.uu.se).

• HITL: Human Interface Technology Laboratory, the University of Washington, includes information on the Virtual Worlds Society and the Virtual Worlds Consortium (http://www.hitl.washington.edu).

• Ubicomp: Ubiquitous Computing–computing and communications available everywhere to help people communicate and get information, information from Mark Weiser of XEROX PARC (http://www.ubiq.com/hypertext/weiser/UbiHome.html).

4.4 Multimedia

4.4.1 Audio

• CDM-NYU: New York University (NYU) Center for Digital Multimedia, a New York State Center for Advanced Technology, sponsored by the New York State Science and Technology Foundation (http://found.cs.nyu.edu).

• CERL: The CERL Sound Group (U of IL) (http://datura.cerl.uiuc.edu).

• Clips: Sites with audio clips (http://www.eecs.nwu.edu/∼jmyers/other-sounds.html).

• Internet Sound: various documents and programs having to do with sound (ftp://ftp.cwi.nl/pub/audio/INDEX).

• Internet Talk Radio: by Carl Malamud (ftp://sunsite.unc.edu/pub/talk-radio/ITRintro.readme).

• Internet Multicasting FAQ: FAQ for the Internet Multicasting Service (mailto:info@radio.com Body: send FAQ).

• Internet Multicasting WWW: Home page for the Internet Multicasting Service (http://www.cmf.nrl.navy.mil/radio/radio.html).

• Internet Talk Radio sites: (mailto:sites@radio.com Body: send SITES).

• IUMA: Internet Underground Music Archive (http://sunsite.unc.edu/ianc/index.html).

• MCRL-Ottawa: Multimedia Communications Research Laboratory at the University of Ottawa (http://mango.genie.uottawa.ca).

• MIDI: Musical Instrument Digital Interface (http://www.eeb.ele.tue.nl/midi/index.html).

• Multicast Backbone: live audio and video multicast virtual network on top of Internet (ftp://venera.isi.edu/mbone/faq.txt).

• Mbone FAQ Web: Frequently Asked Questions (FAQ) on the Multicast Backbone (MBONE) (http://www.research.att.com/mbone-faq.html).

• Music Resources: (http://www.music.indiana.edu/misc/music resources.html).

• NeXT sounds: a collection of sound files for NeXT machines (ftp://wuarchive.wustl.edu/pub/NeXT-Music/).

• Say: Text to Audio (translate text to sound) (http://www tios.cs.utwente.nl/say/form/).

• Sound Site: PC sound (ftp://oak.oakland.edu/pub/misc/sound).

• UnderWorld: Major source of sound information, repositories, music, and voice, by Jennifer Myers (http://www.nd.edu/StudentLinks/jkeating/links/sound.html).

4.4.2 Graphics

• ACM/SIGGRAPH: Association of Computing Machinery (ACM) Special Interest Group on Graphics (gopher://siggraph.org).

• AIG-Manchester: Advanced Interfaces Group (AIG) at the Computer Science Department at the University of Manchester (http://www.cs.man.ac.uk/aig/aig.html).

• CGU-Manchester: The Computer Graphics Unit - Research (http://info.mcc.ac.uk/CGU/CGU-research.html).

• CVU-GA Tech: Georgia Institute of Technology’s Graphics, Visualization, + Usability Center (http://www.gatech.edu/gvu/gvutop.html).

• Scientific Visualization: Annotated Scientific Visualization URL Bibliography (http://www.nas.nasa.gov/RNR/Visualization/annotatedURLs.html).

• Thant’s: animation index (http://mambo.ucsc.edu/psl/thant/thant.html).

• Video: demonstration of vsbrowser, a video file browser (http://tns-www.lcs.mit.edu/cgi-bin/vs/vsbrowser).

• ZGDV: Zentrum f—r Graphische Datenverarbeitung e.V., Computer Graphics Center (http://zgdv.igd.fhg.de).

4.4.3 Multi

• Bell Atl-CNM: Bell-Atlantic Center for Networked Multimedia, interactive multimedia applications over networks (http://www.cnm.bell-atl.com).

• File formats: formats of graphics and sound files (ftp://wuarchive.wustl.edu/pub/doc/graphic-formats/).

• File formats: graphics and sound file formats (ftp://ftp.ncsa.uiuc.edu/misc/file.formats/).

• GNN I-Media: Global Network Navigator’s Interactive Media Center (http://gnn.com/gnn/meta/imedia/index.html).

• ICME-RPI: The International Center for Multimedia in Education at Rensselaer Polytechnic Institute, Troy, New York (http://www.ciue.rpi.edu/index.htm).

• Media Lab: MIT Media Lab Home Page (http://debussy.media.mit.edu).

• MICE: Multimedia Integrated Conferencing for European Researchers (http://www.cs.ucl.ac.uk/mice/).

• Multimedia Index: Multimedia Information Sources, by Simon Gibbs (http://viswiz.gmd.de/MultimediaInfo/).

• Multimedia Survey: A Survey of Distributed Multimedia Research, Standards and Products, by RARE (ftp://ftp.ed.ac.uk/pub/mmsurvey/mmsurvey.txt).

• Multimedia Lab BU: Multimedia Laboratory at Boston University (http://spiderman.bu.edu).

• NYU-Digital: New York University Center for Digital Multimedia (http://found.cs.nyu.edu).

• Rob’s Multimedia Lab: (http://www.acm.uiuc.edu:80/rml).

• Sunsite Multimedia: multimedia presentations based on SunSITE (http://sunsite.unc.edu/exhibits/exex.html).

• TNS Tech demo: Technology Demonstrations–multimedia (http://tns-www.lcs.mit.edu/vs/demos.html).

4.4.4 Hypermedia

• ACM/SIGLINK: Association of Computing Machinery (ACM) Special Interest Group on Hypertext/Hypermedia (gopher://gopher.acm.org/11[the files.sig forums.siglink]).

• Beyond Gutenberg: Hypertext and the Future of the Humanities, Yale University, New Haven, Connecticut, Thursday-Friday May 12-13, 1994 (http://www.cis.yale.edu/htxt-conf/index.html).

• Bush, Vannevar: As We May Think, article from July 1945 issue of The Atlantic Monthly about hypertext (http://www.csi.uottawa.ca/∼dduchier/misc/vbush/as-we-may-think.html).

• H Hyperbook: a simple hypertext markup language (http://siva.cshl.org/h/h.body.html).

• Hypermedia Lab-TAMU: Texas A and M Hypermedia Research Lab (ftp://bush.cs.tamu.edu/pub/home.html).

• Hypermedia/Internet: Hypermedia and the Internet (http://life.anu.edu.au/education/hypermedia.html).

• Hypertext resources: lists of articles, systems, organizations and resources, by Volker Zink (http://www.uni-konstanz.de/FuF/Inf-Wiss/IW/hypertext e.html).

• Hypertext systems: An Overview of Hypertext and IR systems and applications (http://www.w3.org/hypertext/Products/Overview.html).

• HTML Analyzer: assist the maintenance of HyperText MarkUp Language (HTML) databases (http://www.gatech.edu/pitkow/html analyzer/README.html).

• HTML DTD: 1986, Document Type Definition for the HyperText Markup Language as used by the World Wide Web application (HTML DTD) (http://www.w3.org/hypertext/WWW/MarkUp/HTML.dtd.html).

• HTML Specs: hypertext markup language specs, by Tim Berners-Lee and Daniel Connolly (ftp://www.w3.org/pub/www/doc/html-spec.txt.Z).

• HTML Web: a collection of top-level information, from CERN, about hypertext markup language (http://www.w3.org/hypertext/WWW/MarkUp/MarkUp.html).

• Hypermedia Review: State of the Art Review on Hypermedia Issues And Applications, by V. Balasubramanian, Graduate School of Management, Rutgers University, Newark, New Jersey (http://www.csi.uottawa.ca/∼dduchier/misc/hypertext review/).

• Hypertext/rhetoric: Hypertext and the Rhetorical Contract, by Wm. Dennis Horn (http://fire.clarkson.edu/horn/proposal-mla.html).

• Hypertext terms: glossary of terms from the WWW project, from CERN (http://www.w3.org/hypertext/WWW/Terms.html).

• MapMarker: a tool for creating clickable maps for HTML (http://www.dl.ac.uk/CBMT/mapmarker/HOME.html).

4.5 Virtual

• DIS: Distributed Interactive Simulation (ftp://ftp.netcom.com/pub/frankc/dis.html).

• Meme: virtual world development system (http://remarque.berkeley.edu/∼marc/home.html).

• Meta VE: Meta Virtual Environments (http://www.gatech.edu/gvu/people/Masters/Rob.Kooper/Meta.VR.html).

• MIT Media Lab: access to Massachusetts Institute of Technology’s Media Lab (ftp://media-lab.media.mit.edu/access/).

• MIT Media Lab–how to: How To Do Research In the MIT AI Lab (http://www.cs.indiana.edu/docproject/mit.research.how.to/mit.research.how.to.html).

• MIT TNS: MIT’s Telemedia, Networks, and Systems Group (http://tns-www.lcs.mit.edu/tns-www-home.html).

• MSDL: Manchester Scene Description Language (ftp://ftp.mcc.ac.uk/pub/cgu/MSD/scene-survey.tar.Z).

• SUD: Single User VR Resources (gopher://actlab.rtf.utexas.edu/11/SUD).

• VR collection/Texas: (gopher://ftp.cc.utexas.edu:3003/11/pub/output/vr).

• VR Page/Cardiff: (http://www.cm.cf.ac.uk/User/Andrew.Wilson/VR/).

• VR Page/Chris Hand: Research, Papers, Archives, Events, User Groups and so on (http://www.cms.dmu.ac.uk:9999/People/cph/vrstuff.html).

• VR Page/Luke Sheneman: (http://www.cs.uidaho.edu/lal/cyberspace/VR/VR.html).

• VR Web: Virtual Reality Web Page (http://guinan.gsfc.nasa.gov/W3/VR.html).

• VR Archive: Sunsite Virtual Reality archive–papers, information, maintained by David Barberi (http://sunsite.unc.edu/dbarberi/vr.html).

• VR Testbed: Open Virtual Reality Testbed Home Page (http://nemo.ncsl.nist.gov/∼sressler/OVRThome.html).

• VRML: Virtual Reality Modeling Language (http://vrml.wired.com).

• VSR: Virtual Shared Reality Project (http://nfhsg3.rus.uni-stuttgart.de/virtual/index.html).

4.6 Networks

4.6.1 Access and connectivity

These documents will help with gaining or finding out about access to networks.

• BBS Internet List: Zamfield’s Wonderfully Incomplete, Complete Internet BBS List (ftp://sunsite.unc.edu/pub/docs/about-the-net/libsoft/internet bbs.txt).

• Connecting to Internet: What Connecting Institutions Should Anticipate (ftp://nic.merit.edu/documents/fyi/fyi 16.txt).

• Connectivity-Yahoo: Computers-Internet-Connectivity (http://www.yahoo.com/Computers/Internet/Connectivity/).

• DLIST: a list of dedicated line Internet providers, by Susan Estrada (mailto:dlist@ora.com Body: Please send DLIST).

• Freenet papers: papers about network public access (ftp://alfred.carleton.ca/pub/freenet/working.papers/).

• FSLIST: the Forgotten Sites List of public access Internet sites, by Louis Raphael (ftp://freedom.nmsu.edu/pub/docs/fslist/).

• GNET Archive: bring the net to lesser-developed nations and poorer parts of the world (ftp://dhvx20.csudh.edu/global net/).

• Inet-Access FAQ: how to become an Internet service provider, an extremely detailed guide to procedures, equipment, hooking to the Net, agreements, software, fees, technical issues, marketing, legal issues, resources, by David H Dennis (david@amazing.cinenet.net) (http://amazing.cinenet.net/faq.html).

• Internet Access Guide: Access Guide to introducing.the.internet, by Ellen Hoffman (ftp://nic.merit.edu/introducing.the.internet/access.guide).

• Internet Access: Individual access to Internet, by James Milles (ftp://sluaxa.slu.edu/pub/millesjg/internet.access).

• Internet Modem: Internet service providers in the US accessible through dial-up connections from a personal computer, by Genevieve Engel (ftp://dla.ucop.edu/pub/internet/dial-access).

• Internet Providers: All Providers Alphabetically (ftp://sri.com/netinfo/internet-access-providers-alphabetical-listing.txt).

• Internet Providers Non-US: Non-US Internet Providers (ftp://sri.com/netinfo/internet-access-providers-non-us.txt).

• K12 Access: documents of California Education Network (CENet) Technical Planning Committee (ftp://ftp.cc.berkeley.edu/k12/README).

• NIXPUB: Public/Open Access UNIX, by Bux Technical Services (ftp://rtfm.mit.edu/pub/usenet/alt.bbs/Nixpub Posting (Long)).

• Network Startup: NSF-sponsored Network Startup Resource Center (ftp://ftp.psg.com/README).

• PDIAL: The Public Dialup Internet Access List, by Peter Kaminski (ftp://rtfm.mit.edu/pub/usenet/news.answers/pdial).

• PDIAL search: directory provides information on service providers in Northern California, Southern California, and the United States. It was compiled from the Internet pdial listing (http://www.commerce.net/directories/news/inet.prov.dir.html).

• PSGnet/RAINet: networking in the developing world, low-cost networking tools, computer networking in general (gopher://gopher.psg.com).

• RAIN: Regional Alliance for Information Networking (http://www.rain.org).

• Registering on the Net: Transition and Modernization of the Internet Registration Service, by S. Williamson (ftp://nic.merit.edu/documents/rfc/rfc1400.txt).

• Rural Nets/GAIN report: (ftp://nysernet.org/pub/gain/final report).

• Rural Nets: Rural Datafication Project (mailto:rjacot@cic.net).

• Rural Datafication gopher: information about the rural datafication project– Extending Information Highways for Education, Research, and Economic Development in the Great Lakes States (gopher://gopher.cic.net/11/cicnet-gophers/ruraldata-project).

• Rural Datafication Web: bring the power of the Internet to rural and otherwise underserved communities (http://www.cic.net/rd-home.html).

• Service Providers: Network Service Provider WWW Servers (http://www.eit.com/web/www.servers/networkservice.html).

• Slip Starter Kit: help get started with making a slip connection with Windows 3.1, by Jean van Eeden (ftp://aztec.co.za/pub/win3/Slip Starter Kit).

• Winsock: Windows and TCP/IP for Internet access, by Harry M. Kriz (ftp://nebula.lib.vt.edu/pub/windows/winsock/wtcpip05.asc).

4.6.2 Administration

• Domain Administration: (ftp://nic.merit.edu/documents/rfc/rfc1033.txt).

• Domain Name surveys: (gopher://is.internic.net/11/infoguide/about-internet/domain-surveys/).

• Domain Name Survey: an attempt to discover every host on the Internet by doing a complete search of the Domain Name System (http://www.nw.com/zone/WWW/top.html).

• Domain Names: (ftp://nic.merit.edu/documents/rfc/rfc1034.txt).

• Host managers: (ftp://nic.merit.edu/documents/rfc/rfc1173.txt).

• Internet Servers: Building Internet Servers, a collection of information and links from Charm Net (http://www.charm.net/∼cyber/).

• SNMP: simple network management protocol project group, at the University of Twente, the Netherlands (http://snmp.cs.utwente.nl).

4.6.3 Networking

• Amateur Radio Packet: connects between Amateur Radio Packet (digital data stream) network and the Internet (ftp://ftp.std.com/pub/hamradio/faq/packet.faq).

• Andrew Consortium: a portable set of applications that runs under X11 (http://www.cs.cmu.edu:8001/afs/cs.cmu.edu/project/atk-ftp/web/andrew-home.html).

• ATM forum: worldwide organization, aimed at promoting ATM within the industry and the end user community (http://www.atmforum.com).

• ATM Research: Asynchronous Transfer Mode (ATM) Research at Naval Research Lab (http://netlab.itd.nrl.navy.mil/ATM.html).

• Bitnet Address: get the Internet address of a Bitnet host that is also on the internet (send email to host with SHOW Alias) (mailto:listserv@ubvm.bitnet Body: SHOW ALIAS UBVM).

• Bitnet info: a large collection of documentation about Bitnet and EARN (ftp://lilac.berkeley.edu/netinfo/bitnet/).

• BITNET intro: (mailto:listserv@bitnic.educom.edu Body: send BITNET INTRO).

• BITNET nodes: those Bitnet nodes that have internet addresses (mailto:listserv@bitnic.educom.edu Body: get internet listing).

• Bitnet Nodes: a listing of BITNET (Because It’s Time Network) nodes (gopher://nak.berkeley.edu:4303/11/bitnet).

• Cell Relay: cell-relay or broadband technologies (ATM/DQDB/SONET, etc.) including research papers, standards, product information, mailing list archives, and events (gopher://cell-relay.indiana.edu/1).

• Concise: database about networks, networking tools and projects (http://www.w3.org:80/hypertext/DataSources/CONCISE/UserGuide/Overview.html).

• Data communication–Yahoo: Computers-Networks and Data Communications (http://www.yahoo.com/Computers/Networks and Data Communication/).

• Ethernet page: resources related to the Ethernet (IEEE 802.3) local area network system, by Charles Spurgeon (http://wwwhost.ots.utexas.edu/ethernet/ethernet-home.html).

• FidoNet News: archives of the newsgroup comp.org.fidonet (ftp://rtfm.mit.edu/pub/usenet/comp.org.fidonet/).

• FidoNet Nodes: a list of systems with FidoNet (ftp://genome.wi.mit.edu/wais-sources/fidonet-nodelist.src).

• FidoNet Gateway: How to use the UUCP/Fido-Net Gateway, by Lee Damon, Dale Weber, assisted by Lisa Gronke (ftp://ftp.csn.org/pub/mail/internet.fidonet).

• GOSIP: Government Open Systems Interconnection Profile (ftp://%FILE/rfc1169.txt).

• IBM’s collection: networking information, protocols, standards (ftp://networking.raleigh.ibm.com/pub).

• INET 93: Proceedings of the INET ’93 conference (ftp://mordor.stanford.edu/pub/inet93/).

• International Connect: International Connectivity Table, by Larry Landweber (ftp://ftp.cs.wisc.edu/connectivity table/).

• Internet Country Codes: FAQ about country codes (ftp://rtfm.mit.edu/pub/usenet/news.answers/mail/country-codes).

• Internet Domain Names: Relationship of Telex Answerback Codes to Internet Domains (ftp://nic.merit.edu/documents/rfc/rfc1394.txt).

• Internet + Networking-Galaxy: list of resources related to Internet and Networking from EINet’s Galaxy (http://galaxy.einet.net/Reference-and-Interdisciplinary-Information/Internet-and-Networking.html).

• Internet Protocols: listings of working groups and information about protocols–applications, internet, next generation, network management, operational requirements, routing, security, and much more (http://netlab.itd.nrl.navy.mil/Internet.html).

• Internet Root Domain: lists of Internet hosts (ftp://ftp.rs.internic.net/domain/).

• InterNetwork Mail: methods of sending mail from one network to another, by John Chew and Scott Yanoff (ftp://ftp.csd.uwm.edu/pub/internetwork-mail-guide).

• Intro TCP/IP: describes the Internet protocols (ftp://nic.merit.edu/introducing.the.internet/intro.to.ip).

• IP address resolver: (mailto:resolve@widener.edu Body: site SITE NAME).

• ISDN: the Combinet, Inc. Integrated Services Digital Network (ISDN) deployment database (telnet://isdn@bbs.combinet.com).

• ISDN info/Bellcore: collection of Integrated Services Digital Network (ISDN)information, from Bellcore (ftp://info.bellcore.com/pub/ISDN/).

• ISDN info/Kegel: ISDN Page, a collection of pointers resources about Integrated Services Digital Network (ISDN), including standards and discussions, providers, vendors, products, by Dan Kegel (http://alumni.caltech.edu/∼dank/isdn/).

• ISDN info/Pac Bell+ATT: Integrated Services Digital Network (ISDN) from Pacific Bell and American Telephone and Telegraph (http://www.pacbell.com/isdn/isdn home.html).

• Matrix info: bitnet, fidonet, fredmail, arpanet, ddn (gopher://is.internic.net/11/infoguide/about-internet/other-networks).

• Matrix: information about connected email systems (Quarterman’s Matrix)

(gopher://nkosi.well.sf.ca.us/11/matrix).

• Minitel: French Videotex terminal(Minitel) emulator, requires X11 terminal or simulator (http://www.enst.fr/∼meunier/english/minitel/).

• Network Research sites: a list of network researching sites (http://netlab.itd.nrl.navy.mil/onr.html).

• Networking Overview: Overview of information available (http://web.doc.ic.ac.uk:80/bySubject/Networking.html).

• Networking EINet: list of resources related to networking by EINet’s Galaxy (http://galaxy.einet.net/galaxy/Engineering-and-Technology/Computer-Technology/Networking.html).

• PSGnet/RAINet info: networking in developing world, low-cost tools, networking in general (gopher://rain.psg.com).

• NREN Information: Merit’s directory of National Research and Education Network information (ftp://nic.merit.edu/nren/INDEX.nren).

• NREN Recompetition: contains information about the recompetition of the NSFNET backbone project (ftp://nic.merit.edu/cise/recompete/INDEX.recompete).

• OneNet: a global network of Macintosh computers (phone:415-948-4775).

• Personal IP: PPP, MS-Windows and other information and links about connecting with internet protocols, from Charm Net (http://www.charm.net/ppp.html).

• PCLT: PC Lube and Tune; informative introductory material on PC hardware, networks, and newer operating systems (http://pclt.cis.yale.edu/pclt/default.htm).

• RSA info: information on many cryptographic related topics (ftp://rsa.com).

• SDSC Appl Net Res Group: activities of the San Diego Supercomputer Center Applied Network Research group. (ftp://ftp.sdsc.edu/pub/sdsc/anr/README).

• Sprintlink: FTP site for Sprint’s networking activities (ftp://ftp.sprintlink.net).

• Sprintlink Gopher: Sprint’s internetworking activities and networking information (gopher://ftp.sprintlink.net).

• Wireless: The Wireless Opportunities Coalition, a group of organizations and companies seeking to expand wireless communications development, manufacturing, and use (http://wireless.policy.net/wireless/wireless.html).

4.6.4 Security

• Business Security: Security for Businesses on the Internet, by Marianne Mueller (http://www.catalog.com/mrm/security.html).

• CERT FTP: Computer Emergency Response Team at Carnegie Mellon Univ.

(ftp://cert.org/pub/).

• CERT-DFN: Computer Emergency Response Team for the German Research Network (http://www.cert.dfn.de/eng/).

• Cryptorebel/Cypherpunk: Vince Cate’s Cryptorebel and Cypherpunk page (ftp://furmint.nectar.cs.cmu.edu/security/README.html).

• CSC: Computer Systems Consulting, system security issues information (http://www.spy.org).

• DoD Security: DoD Goal Security Architecture (DGSA) (ftp://asc.dtic.dla.mil/pub/tafim/).

• First: Forum of Incident Response and Security Teams (http://first.org).

• Hack/phreak: resources, happenings, connections, from Randy King (http://www.phantom.com/∼king/).

• Internet Security: GAO report on Computer Security (June 1989)

(ftp://nic.merit.edu/cise/gao8957.txt).

• Internet Worm: a collection of papers about internet security compromises (ftp://nic.funet.fi/pub/doc/security/worm/).

• LOT: Los Altos Technologies; offers Gabriel to combat Satan (http://www.lat.com).

• NIST Security: US National Institute of Standards and Technology (NIST) Computer Security Resource Clearinghouse (http://csrc.ncsl.nist.gov).

• SAIC: Science Applications International Corp, computer security (http://mls.saic.com/mls.security.text.html).

• Security index: Computer and Network Security Reference Index, by Rodney Campbell (http://www.tansu.com.au/Info/security.html).

• SHEN: A Security Scheme for the World Wide Web (http://www.w3.org/hypertext/WWW/Shen/ref/shen.html).

• Site Security: Site Security Handbook, FYI 8, guidance on how to deal with security issues in the Internet, eds. Holbrook, Reynolds (ftp://nic.merit.edu/documents/fyi/fyi 08.txt).

4.6.5 Statistics

• IBC Stats: Internet Business Center’s collection of Net statistics–shows lists of Net cities, states, Net Presence by industry, from The Internet Group (http://tig.com/IBC/Statistics.html).

• Internet Charts/ISOC: charts of traffic, connectivity, hosts, etc., from the Internet Society (ftp://ftp.isoc.org/isoc/charts/).

• Internet Growth: charts showing the Internet’s past and projected growth, by Texas Internet Consulting (ftp://tic.com/matrix/growth/internet/).

• Internet Growth/Lottor: Internet Growth (1981-1991)

(ftp://nic.merit.edu/documents/rfc/rfc1296.txt).

• Internet Stats/Demographics-Yahoo:

(http://www.yahoo.com/Computers/Internet/Statistics and Demographics/).

• NSFnet stats: NSF Statistics about Internet use, from Merit (ftp://nic.merit.edu/nsfnet/statistics/).

• NSFnet stats/GVU Center: Georgia Tech’s Graphics, Visualization, and Usability Center NSFNET Backbone Statistics Page, includes graphs of statistics (http://www.cc.gatech.edu/gvu/stats/NSF/merit.html).

4.6.6 Maps

• ARPAnet Map: an index of Interface Message Processors on the ARPAnet (circa 1986) (http://web.kaleida.com/u/hopkins/arpanet/arpanet.html).

• Internet Maps (Many): (ftp://ftp.uu.net/inet/maps/).

• Internet Maps (NSFNET): (ftp://nic.merit.edu/maps/).

• Internet Maps (SuraNet): (ftp://ftp.sura.net/pub/maps/).

• Internet/Matrix: maps from MIDS (Matrix Information and Directory Services)

(gopher://gopher.tic.com/11/matrix/maps/matrix).

• Internet Topology-Yahoo: (http://www.yahoo.com/Computers/Internet/Network Topology/).

• USENET Maps: Maps of Usenet news feeds/backbones (ftp://gatekeeper.dec.com/pub/maps/).

• UUCP Maps: Unix-Unix Copy Protocol Map Data (gopher://agate.berkeley.edu:4324/1uumaps).

• WWW Resource Maps: The Virtual Tourist, a collection of maps from all over the world to help you locate Internet sites and resources (http://wings.buffalo.edu/world).

4.7 Telecommunications

• ATP-LLNL: Advanced Telecommunications Program at Lawrence Livermore National Laboratory (http://www-atp.llnl.gov/atp/).

• Computer + Communications: InfoBahn, Global Information Infrastructure, Telecommunications, large resource collection of companies, media, organizations, programs and projects, standards, and usenet groups and FAQs (http://www-atp.llnl.gov/atp/telecom.html).

• Computing + Telecom: World-Wide Web Virtual Library entry for Communications and Telecommunications (http://www.analysys.co.uk/commslib.htm).

• CTR-Columbia U Web: Columbia University Center for Telecommunications Research (CTR) (http://www.ctr.columbia.edu/CUCTR Home.html).

• Data Comm/Networking: Data Communications and Networking Links, by Don Joslyn (http://www.racal.com/networking.html).

• INT-France: Institut National des Telecommunications, France (http://arctique.int-evry.fr).

• ITC: International Telecommunications Center–telecommunications, data communications and networking; includes archives, information, software, product and employment information, sponsored by telematrix.com (http://www.telematrix.com).

• Tele/Communications: Information Sources about Communications and Telecommunications (http://www.tansu.com.au/Info/communications.html).

• Telecomm Archives: files about telecommunications, from the Usenet group comp.dcom.telecom (ftp://lcs.mit.edu/telecom-archives/).

• TelecomInfo: from New York State Department of Education gopher (gopher://unix5.nysed.gov/11/TelecommInfo).

• Telecom Information Resources: technical, economic, public policy, and social aspects of telecommunications including voice, data, video, wired, wireless, cable TV, and satellite, are included. (http://www.ipps.lsa.umich.edu/telecom-info.html).

• Telephone industry: Telephone Industry Information Page, a service of The Telephone Customer’s Corner (http://www.teleport.com/∼mw/cc/tii.html).

• TIS-Kansas: Telecommunications and Information Sciences Laboratory, University of Kansas (http://www.tisl.ukans.edu).

• US-FCC FTP: Federal Communications Commission (USA) (ftp://ftp.fcc.gov).

• US-FCC Gopher: Federal Communications Commission (USA) (gopher://gopher.fcc.gov).

• US-FCC web: Federal Communications Commission (USA) (http://www.fcc.gov).

• US-ITS: Institute for Telecommunication Sciences, USA government research and engineering laboratory (http://www.its.bldrdoc.gov/its.html).

• US-NTIA: National Telecommunications and Information Administration (USA)

(telnet://ntiabbs.ntia.doc.gov).

• US-NTIA web: National Telecommunications and Information Administration (USA) (http://www.ntia.doc.gov).

• WilTel Library: Telecommunications Library, telecom business and technology, sponsored by WilTel Network Services (http://www.wiltel.com/library/library.html).

5 CULTURE 45

5 CULTURE

5.1 Art

• Art-Yahoo: (http://www.yahoo.com/Art/).

• ANIMA: Arts Network for Integrated Media Applications (http://wimsey.com/anima/ANIMAhomeF.html).

• Art/images: a collection of art and images in several formats (ASCII, TIFF, GIF, JPEG) (gopher://cs4sun.cs.ttu.edu/11/Art%20and%20Images).

• Art/Net: Art on the Net, artists share and create works together on the Internet (http://www.art.net).

• ArtSource: a gathering point for networked resources on Art and Architecture (http://www.uky.edu/Artsource/artsourcehome.html).

• Artwork: OTIS project to collect online art (ftp://sunsite.unc.edu/pub/multimedia/pictures/OTIS/).

• ASCII Art: A FAQ about ASCII art, by Jorn Barger (ftp://ftp.mcs.com/mcsnet.users/jorn/asciifaq.txt).

• ASCII Art Bazaar: collection of ASCII art and images, 12 megabytes of information covering an estimated 24,000 art works from more than 3300 contributions classified under 759 subject titles (gopher://twinbrook.cis.uab.edu/1asciiarc.70).

• ASCII Art collection: a collection of files showing images, fonts, and other ASCII art (ftp://ftp.cs.ttu.edu/pub/asciiart/).

• Cirque de la Mama: to bring works of art to people and to bring people to works of art (http://lancet.mit.edu/cirque/cirque.html).

• CIS-AH: Center for Integrative Studies - Arts and Humanities (http://web.cal.msu.edu).

• Electric Gallery: presents naive and primitive art that is unique and famous thoroughout the world (http://www.egallery.com/egallery/).

• FineArt Forum: List of Art Related Web Resources (http://www.msstate.edu/Fineart Online/art-resources.html).

• Free Art: pages and the graphics, by Harlan Wallach (http://www.mcs.net/∼wallach/freeart/buttons.html).

• NWHQ: New World Headquarters–free expression and the distribution of artistic ideas, independent artists supporting independent artists (http://www.knosso.com/NWHQ/index.html).

5.2 Community

5.2.1 Information

• Civic Networking: WWW Guide to Community Networking, by Catherine Kummer (http://http2.sils.umich.edu/ILS/community.html).

• Community Nets/McGee: information about community networks, by Arthur R. McGee (ftp://ftp.netcom.com/pub/am/amcgee/community/).

• CPSR Community Net info: (http://www.cpsr.org/dox/community.nets.html).

• Free-Nets Home Page: Free-net information, presented as a public service by Peter Scott, at the University of Saskatchewan Libraries (http://herald.usask.ca/∼scottp/free.html).

• Free-Net List/Gopher: (gopher://gopher.tamu.edu/11/.dir/freenet.dir).

• Free-Net List/Hytelnet: (http://www.cc.ukans.edu/hytelnet html/FRE000.html).

• NPTN: National Public Telecomputing Network (ftp://nptn.org/pub/).

5.2.2 Virtual

—– Some instances of virtual communities.

• Blacksburg electronic village: (http://www.bev.net).

• CIAO!: British Columbia, Canada (telnet://ciao.trail.bc.ca).

• Cleveland Free-Net: the world’s first community Fee-Net (telnet://visitor@freenet-in-a.cwru.edu).

• Digital City: de Digitale Stad, Amsterdam, Netherlands (telnet://dds.hacktic.nl).

• EnviroFreenet: an on-line community of people who are concerned about the earth (telnet://envirolink.org).

• Internet Town Hall: a service of the Internet Multicasting Service (http://www.town.hall.org).

• Oceania: The Atlantis Project; new country in development–you may get Constitution and Laws, passport info, free subscription to newsletter Oceania Oracle, and more (http://oceania.org).

• Silicon Valley: Silicon Valley Public Access Link Community Page (http://www.svpal.org).

• Trailer Park: home pages for the homepagesless (http://www.cybernetics.net/users/phrantic/marty.htm).

• Virtual City: Virtual City Network Project (http://virtual.net/VirtualCity/).

5.3 Language

• ABU: L’Association des bibliophiles Universels (http://www.cnam.fr/ABU/principal/ABU.v2.html).

• ACW web: Alliance for Computers and Writing (http://prairie island.ttu.edu/acw/acw.html).

• ALEX: find and retrieve the full-text of documents on the Internet from such archives as Project Gutenberg, Wiretap, the On-line Book Initiative, the Eris system at Virginia Tech, the English Server at Carnegie Mellon University, and the online Oxford Text Archive (gopher://www.lib.ncsu.edu/stacks/alex-index.html).

• Babel: A glossary of computer-oriented abbreviations and acronyms (ftp://ftp.temple.edu/pub/info/help-net/babel95a.txt).

• Baylor etexts: (ftp://ftp.byu.edu/pub/next/Literature/).

• Book collection: electronic books, reference, special collections (gopher://psulias.psu.edu/11[ shelves]).

• Citation: Bibliographic References for Computer Files in the Social Sciences, A Discussion Paper, by Sue A. Dodd (ftp://ftp.msstate.edu/pub/docs/history/netuse/electronic.biblio.cite).

• Computer Writing: Computer Generated Writing resources, by Marius Watz (http://www.uio.no/∼mwatz/c-g.writing/).

• Computer Jargon search: search the jargon file on WWW (http://web.cnam.fr/bin.html/By Searchable Index?Jargon File.html).

• Computer Jargon: The Jargon File (the definitive compendium of hacker slang)

(ftp://aeneas.mit.edu/pub/gnu/jargon-README).

• Computing Dictionary (gopher): The Free On-line Dictionary of Computing (gopher://wombat.doc.ic.ac.uk).

• Computing Dictionary (web): The Free On-line Dictionary of Computing (http://wombat.doc.ic.ac.uk).

• CMC Glossary: Compiled by Collins and Berge, from St. John’s University (gopher://sjumusic.stjohns.edu:1070/11/%40uni%3acmc.glossary).

• CMT: The Center for Machine Translation (CMT) at the School of Computer Science at Carnegie Mellon University; advanced research and development in natural language processing, with a focus on multi-lingual machine translation (http://www.mt.cs.cmu.edu/cmt/CMT-home.html).

• CPET: Catalogue of Projects in Electronic Text (ftp://guvax.georgetown.edu/cpet projects in electronic text/).

• CSLI-Stanford: Center for the Study of Language and Information (CSLI), Stanford University (http://csli-www.stanford.edu).

• CTI-Oxford: Centre for Textual Studies, Oxford (mailto:ctitext@vax.ox.ac.uk).

• eText-Caltech: the eText (electronic hypermedia textbooks) group at Caltech (http://www.etext.caltech.edu).

• Electronic Text archive: copy of archive services on etext.archive.umich.edu (gopher://fir.cic.net/00/0-README).

• Electronic Text: Catalogue of projects in electronic text, Center for Text and Technology (ftp://guvax.georgetown.edu/cpet projects in electronic text/).

• ETC-UV: Electronic Text Center – University of Virginia (http://www.lib.virginia.edu/etext/ETC.html).

• Electronic Word: Democracy, Technology, and the Arts, by Richard A. Lanham (ftp://press-gopher.uchicago.edu/pub/Excerpts/lanham.txt).

• Gutenberg Project: a project to give away online texts, hundreds of titles (http://jg.cso.uiuc.edu/PG/welcome.html).

• Hacker’s Dictionary: a searchable index of Hacker’s Jargon (http://iicm.tu-graz.ac.at/Cjargon).

• Human: human languages page, cataloging human-language resources and making those resources available to the Web community through a concise index (http://www.willamette.edu/∼tjones/Language-Page.html).

• IBIC: Internet Book Information Center (http://sunsite.unc.edu/ibic/IBIC-homepage.html).

• Internet Glossary: search a collection of Internet terms (wais://pinus.slu.se:210/Internet-user-glossary?).

• Internet Glossary: Internet-specific terms defined (ftp://nic.merit.edu/documents/fyi/fyi 18.txt).

• Internet Wiretap: electronic books and information (ftp://wiretap.spies.com/About/).

• Internet Wiretap Gopher: electronic books and information (gopher://wiretap.Spies.COM).

• ITK: Instituut voor Taal- en Kennistechnologie, Institute for Language Technology and Artificial Intelligence (http://itkwww.kub.nl:2080/itk/itkhome.html).

• Jargon: Jargon File Resources, browse or download the file in different forms (http://www.ccil.org/jargon/jargon.html).

• LETRS: Library Electronic Text Resource Service (gopher://gopher.indiana.edu:1067/11/letrs/gopher).

• Natural Language: Natural Language Software Registry, summary of the capabilities and sources of language processing software available to researchers (http://cl-www.dfki.uni-sb.de/cl/registry/ed note.html).

• OBI: Online Book Initiative (ftp://ftp.std.com/obi/README).

• Online books: an index of hundreds of on-line books, collections of on-line books and documents (http://www.cs.cmu.edu:8001/Web/books.html).

• OWL: Purdue Writing Labs’ Online Writing Lab (OWL) (mailto:owl@sage.cc.purdue.edu Subject: owl-request).

• Oxford Archive: Oxford Text Archive (ftp://ota.ox.ac.uk/ota/).

• Post-Gutenberg Galaxy: The Fourth Revolution in the Means of Production of Knowledge, by Steven Harnad (ftp://infolib.murdoch.edu.au/pub/jnl/harnad.jnl).

• Scholar: Natural Language Processing On Line (gopher://jhuniverse.hcf.jhu.edu/11/.HAC/Journals/.SCHOLAR).

• Smileys (all): all the smileys in the known universe (gopher://gopher.ora.com/00/feature articles/universe.smiley).

• Smiley Catalog: (telnet://help@twinbrook.cis.uab.edu:3399).

• Smileys: a smiley server by David W. Sanderson (ftp://ftp.uu.net/usenet/comp.sources.misc/volume23/smiley/part01.Z).

• Writer’s Resources: Internet Writer Resources, Compiled/Edited/Maintained by L. Detweiler (ftp://rtfm.mit.edu/pub/usenet/news.answers/writing/resources).

5.4 People

5.4.1 Aspects

• APA: American Psychological Association (http://www.apa.org).

• Cognitive/Psychological sources: information sources on the Internet about academic programs, periodicals, network resources, and many other online resources, by Scott Mainwaring (http://matia.stanford.edu/cogsci.html).

• GA Tech–Cognitive: Cognitive Science at Georgia Tech (http://www.cc.gatech.edu/cogsci/cogsci.html).

• HCRL-Open U: Human Cognition Research Laboratory, The Open University, Milton Keynes, United Kingdom (http://hcrl.open.ac.uk).

• Webaholics: resources that will get you hooked to the Web; includes a Forms interface to add entries to a Webaholics support group (http://www.ohiou.edu/∼rbarrett/webaholics/ver2/).

5.4.2 Lists – directories, home pages

—– Major general collections

• Who’s Who: Who’s Who on the Internet, from CityLive! Magazine, part of the WWW Virtual Library (http://web.city.ac.uk/citylive/pages.html).

• GNN’s Netizen’s: Global Network Navigator’s Internet Center Netizen’s project, a directory of home pages written by GNN users (http://nearnet.gnn.com/gnn/netizens/index.html).

• People-Yahoo: Entertainment–People, personal home pages collection, from Yahoo (http://www.yahoo.com/Entertainment/People/).

• Personal Pages World Wide: a meta collection of personal pages worldwide, from the University of Texas at Austin (http://www.utexas.edu/world/personal/index.html).

• Who’s Online: a collective database of a non-commercial biographies of people on the Net (http://www.ictp.trieste.it/Canessa/whoiswho.html).

• Galaxy’s Net Citizens: (http://www.einet.net/galaxy/Community/Net-Citizens.html).

• DA-CLOD People Page: Distributedly Administered Categorical List Of Documents collection of

personal home pages (http://schiller.wustl.edu/DACLOD/daclod?id=00008.dcl). —– Specialized collections

• CMC People: people interested in the study of CMC, from the CMC studies center (http://www.rpi.edu/∼decemj/cmc/people.html).

• COS: Community of Science, identify and locate researchers, inventions, and facilities at U.S. and Canadian universities by interest and expertise (http://cos.gdb.org).

• Communications Scholars: (http://alnilam.ucs.indiana.edu:1027/sources/dirpage.html).

• Four11: Directory Services, a free and easy-to-use directory of online users and their e-mail addresses, from Four11 Directory Services (SLED) (http://www.four11.com).

• HCI People: people interested in human-computer interaction (http://www.cs.bgsu.edu/HCI/bin/people).

• Houh’s People: Henry Houh’s List of People on the Web (http://tns-www.lcs.mit.edu/people/hhh/people.html).

• Home Page Publisher: create/edit your own home page with its own URL, includes a collection of home pages (http://www-bprc.mps.ohio-state.edu/HomePage/).

• Internet People: Who’s Who in the Internet, Biographies of IAB, IESG and IRSG Members (ftp://nic.merit.edu/documents/fyi/fyi 09.txt).

• Netpages: a phone-book style directory for the Internet, from Aldea Communications, Inc. (ftp://ftp.sco.com/NetPages/).

• WBW: World Birthday Web, you can record your birthday and link to your home page, by Tom Boutell (http://sunsite.unc.edu/btbin/birthday).

• WWPR: World-Wide Profile Registry (commercial) (http://snark.wizard.com/wwpr.html).

5.5 Society

• ACM SIGCAS: Association for Computing Machinery (ACM) Special Interest Group on Computers and Society (gopher://gopher.acm.org/11[the files.sig forums.sigcas]).

• Anonymity FAQ: Anonymity on the Internet (ftp://rtfm.mit.edu/pub/usenet/news.answers/net-anonymity/).

• CCH: Centre for Computing in the Humanities (gopher://alpha.epas.utoronto.ca/11/cch).

• Computer Underground: The Social Organization of the Computer Underground, an MS Thesis by Gordon R. Meyer (ftp://ftp.eff.org/pub/Publications/CuD/Papers/meyer).

• Coombs Papers: science and humanities papers, bibliographies, directories, theses abstracts and other high-grade research material (gopher://coombs.anu.edu.au).

• Culture/Tech: a collection of texts exploring relationship of technology and culture (http://english-server.hss.cmu.edu/Cyber.html).

• Cyber-Culture-Yahoo: Society and Culture-Cyber-Culture (http://www.yahoo.com/Society and Culture/Cyber Culture/).

• Cyber papers/EFF: collection of cyberspace-related papers, from the Electronic Frontier Foundation (ftp://ftp.eff.org/pub/Publications/CuD/Papers/).

• Cyberspace: The New Frontier, by The Laboratory for Applied Logic at the University of Idaho (http://www.cs.uidaho.edu/lal/cyberspace/cyberspace.html).

• Cyberpunk FAQ: from the Usenet newsgroup alt.cyberpunk (ftp://rtfm.mit.edu/pub/usenet/news.answers/cyberpunk-faq).

• Cypherpunk Topics: information pulled off of many sources, mostly sci.crypt and the cypherpunks mailing list (ftp://ftp.u.washington.edu/public/phantom/cpunk/README.html).

• Cypherpunks gopher: clipper, DC Nets, digital cash, protocols, other info (gopher://chaos.bsu.edu).

• Cypherpunks home page: PGP, remailers, rants, various crypto-tools, newspaper clippings, and a good deal of other things (ftp://ftp.csua.berkeley.edu/cypherpunks/Home.html).

• MetaNet Gopher: management, organizational change and development, education, the arts and the humanities, the impact of technology on society, the future, law, health, the environment, public policy, reinventing government (gopher://gopher.tmn.com).

• English Server: examine the possibilities of collaborative, community-run communications (http://english-server.hss.cmu.edu).

• Fourth World: Indigenous Peoples’ Information for the Online Community (http://www.halcyon.com/FWDP/fwdp.html).

• Future Culture: FutureCulture FAQ, or Cyberography, is maintained by Andy Hawks (mailto:future-request@nyx.cs.du.edu Subject: send faq).

• Future Culture: information about Net culture, media, virtual communities, cyberpunk, memetics (http://www.uio.no/∼mwatz/futurec/).

• Gender Issues: Gender Issues in Computer Networking, by Leslie Regan Shade (ftp://alfred.carleton.ca/pub/freenet/93conference/papers/leslie regan shade.txt).

• Gender/Spertus: Ellen Spertus’ Writings on Gender and Science/Engineering (http://www.ai.mit.edu/people/ellens/gender.html).

• Global/Women: Global Fund for Women, an international grantmaking organization (http://www.ai.mit.edu/people/ellens/gfw.html).

• Humanities: Ad Hoc Steering Committee of the National Initiative for Humanities and Arts Computing (mailto:rre-request@weber.ucsd.edu Subject: archive send humanities).

• IEEE SSIT: The Society on Social Implications of Technology (SSIT) of the Institute of Electrical and Electronics Engineers (IEEE) (http://www4.ncsu.edu/unity/users/j/jherkert/index.html).

• Internet demographics: a survey by Texas Internet Consulting (ftp://ftp.tic.com/survey/).

• McGee collection: Art McGee’s collection on culture and society issues–Activism, African, community, development, gender, indigenous, Latin (ftp://ftp.netcom.com/pub/am/amcgee/).

• Net Ethics: Ethics and the Internet (ftp://nic.merit.edu/documents/rfc/rfc1087.txt).

• Net Etiquette Guide: by Arlene H. Rinaldi (ftp://ftp.lib.berkeley.edu/pub/net.training/FAU/).

• Net Rights/NRC: Rights and Responsibilities of Participants in Networked Communities, from the US National Research Council, National Academy Press, National Academy of Sciences (http://www.nas.edu:70/1/nap/online/rights).

• Netizen Anthology: The Netizens and the Wonderful World of the Net–an anthology, Ronda Hauben, Michael Hauben (http://www.columbia.edu/∼hauben/project book.html).

• Networking: Networking on the Network, by Phil Agre (mailto:rre-request@weber.ucsd.edu Subject: archive send network).

• Privacy: archive about network privacy, from usenet group alt.privacy (ftp://rtfm.mit.edu/pub/usenet/alt.privacy/).

• Privacy forum: materials for the Privacy forum, including all Privacy Forum Digest issues and materials (gopher://vortex.com/11/privacy).

• SeniorNet: senior citizen’s network on America On Line (voice) (phone:415-750-5030).

• SeniorNet Profile: (http://nearnet.gnn.com/gnn/bus/senior/index.html).

• Togethernet: on-line information and communication network for sustainable planetary future (gopher://gopher.together.uvm.edu).

• Usenet Oracle: (mailto:oracle@cs.indiana.edu Subject: help).

• Women: online writings and resources by/about/for women (http://www.mit.edu:8001/people/sorokin/women/index.html).

• UWI Cultural Play: UWI’s Web’s Edge/UnderWorld Industries’ Cultural Playground (http://kzsu.stanford.edu/uwi.html).

6 FORUMS 50

6 FORUMS

These are forums for interaction and information dissemination related to CMC and networking.

6.1 Academic resources in communication and related fields

• ACM SIGCOMM: Association for Computing Machinery (ACM) Special Interest Group on Communications (http://gatekeeper.dec.com/pub/doc/sigcomm/ccr/overview.html).

• CCS-MIT: The Center for Coordination Science, Massachusetts Institute of Technology (http://www-sloan.mit.edu/ccs/ccsmain.html).

• CHC-Oxford: Centre for Humanities Computing (CHC) based at Oxford University Computing Services (OUCS) (http://www.ox.ac.uk/depts/humanities/).

• Communication and Education: networks, communications, and educational applications of networking, by Chandrasekhar Vallath (http://alnilam.ucs.indiana.edu:1027).

• Communications-Galaxy: Galaxy’s page on communications (http://www.einet.net/galaxy/Artsand-Humanities/Language-and-Literature/Communications.html).

• Complex systems page: Complex (Adaptive) Systems Information, by Alex Mallet (http://www.seas.upenn.edu/∼ale/cplxsys.html).

• Complex systems resources: Complex Systems Resources, by David Green and Terry Bossomaier (http://life.anu.edu.au/complex systems/complex.html).

• CMC Studies Center: a resource sharing service for people interested in Computer-Mediated Communication (http://www.rpi.edu/∼decemj/cmc/center.html).

• Comserve: online service for human communication research (mailto:comserve@vm.its.rpi.edu Body: send Comserve Helpfile).

• Comserve gopher: online service for human communication research (gopher://cios.llc.rpi.edu).

• Coombs Web: information for researchers in the field of social sciences and humanities (http://coombs.anu.edu.au/CoombsHome.html).

• Cybernetics/systems theory: (http://pespmc1.vub.ac.be/CYBSYSTH.html).

• Internet studies: social, political and cultural issues of the Internet, from Coombs Computing Unit (http://coombs.anu.edu.au/SpecialProj/QLTY/NetStudies.html).

• Iowa Comm gopher: Gopher server at the University of Iowa featuring many communications-related resources, including advertisiting, CMC, diversity, journalism, media studies, multimedia, and telecommunications (gopher://iam41.arcade.uiowa.edu:2270).

• ISYS: A full text searching/retrieval facility of Comserve’s databases or hotline discussions (mailto:comserve@vm.its.rpi.edu Body: help topics isys).

• KMI: Knowledge Media Institute, learning and cognitive sciences melded with computing and telecommunications; includes Knowledge Systems Group of the Human Cognition Research Laboratory, Electronic Media Research Group, Multimedia Enabling Technologies Group (http://kmi.open.ac.uk).

• McLuhan Probes: from the Herbert Marshall McLuhan Foundation (http://www.mcluhan.ca/mcluhan/).

• Poynter: The Poynter Institute for Media Studies (http://www.nando.net/prof/poynter/home.html).

• Social Science: Social Science Information Gateway (http://sosig.esrc.bris.ac.uk).

• UCSD-comm: U.C. San Diego listing of Resources in Communication (http://gort.ucsd.edu/ek/comm/comm.html).

6.2 Meetings and Activities

• ACM: Association for Computing Machinery conference calendar (gopher://gopher.acm.org/11[the files.calendar]).

• ACM/SIGCHI: Association of Computing Machinery (ACM) Special Interest Group on Computers and Human Interaction conferences (gopher://gopher.acm.org/11[the files.sig forums.sigchi]).

• AusWeb95: the first Australian conference on the WorldWideWeb, Ballina Beach Resort, Ballina, Far North Coast of New South Wales, Australia. 29 April to 2 May 1995 (http://www.scu.edu.au/ausweb95/index.html).

• CMC Activities: Activities page from the Computer-Mediated Communication Studies Center (http://www.rpi.edu/∼decemj/cmc/activities.html).

• Conferences-Yahoo: Computers-Internet-Conferences and Events (http://www.yahoo.com/Computers/Internet/Conferences and Events/).

• CPSR: Computer Professionals for Social Responsibility annual meetings (http://www.cpsr.org/dox/confs.html).

• DIAC: Directions and Implications in Advanced Computing, a CPSR (Computer Professionals for Social Responsibility) conference (http://www.cpsr.org/dox/diac.html).

• Digital Media: (http://programs.digitalworld.com).

• EDUCOM: conferences and events (http://educom.edu/conf.semi/.index.html).

• Hypertext/Networking: conferences relevant to hypertext and networking, from CERN virtual library (http://www.w3.org/hypertext/Conferences/Overview.html).

• GNN Calendar: (http://nearnet.gnn.com/gnn/news/calendar/index.html).

• Internic conferences: (gopher://ds1.internic.net/11/pub/conf.announce).

• Internet Conferences: a list of activities related to the Internet, the Internet Conference Calendar (ICONC) from Automatrix, Inc. (http://www.automatrix.com/conferences/).

• Internet User Groups: a list of Internet-related groups which hold regular meetings, by Will Clurman and Dan Murray (http://www.rmiug.org/rmiug/usergroups.html).

• Networld/Interop: current NetWorld+Interop exposition and conference information (http://programs.interop.com).

• Online: Online, Inc.’s conferences for information professionals; Online/CD-ROM, Multimedia Schools conferences (gopher://online.lib.uic.edu).

• Telecom events: conferences, exhibitions, and trade shows worldwide in the fields of telecommunications, electronics, fiber optics, computers, opto-electronics, and end user applications (http://www.telematrix.com/world/).

• US NII: (http://iitf.doc.gov:70/1s/calendar).

• WWW Conferences: International conferences on the Web, past and future (http://www.w3.org/hypertext/Conferences/Overview-WWW.html).

• WWW Conferences-Yahoo:

(http://www.yahoo.com/Computers/World Wide Web/Conferences/).

6.3 Discussion

There are many discussion forums related to CMC and Internet topics. Use the following specialized directories to locate a forum, or use the PAML Guide listed under Internet-Directories, above. —– Some Lists of Discussion lists related to Internet/CMC

• Comm studies: communications related groups from Directory of Scholarly Electronic Conferences (gopher://una.hh.lib.umich.edu/00/inetdirsstacks/commun%3afehrmann).

• Computers: Listserv lists related to Computers, from InfoMagnent (http://www.clark.net/pub/listserv/lscmp1.html).

• Education: education-related discussion groups and electronic journals (http://k12.cnidr.org:90/lists.html).

• Internet and Bitnet: Listserv lists related to Internet and Bitnet, from InfoMagnent (http://www.clark.net/pub/listserv/lslsbn1.html).

• Library: Library-Oriented Lists and Electronic Serials, compiled by Steve Bonario, SBonario@uh.edu and Ann Thornton, AThornton@uh.edu, Systems Librarians, University of Houston (gopher://info.lib.uh.edu/00/tools/netinfo/library).

• WWW List of Lists: lists of mailing lists and Usenet groups on topics related to the World Wide Web, by Brian Kelly (http://www.leeds.ac.uk/ucs/WWW/WWW mailing lists.html).

—– Selected Discussion/information lists

• ANN-LOTS: annotating lists of things (mailto:listserv@vm1.nodak.edu Body: subscribe ann-lots YOUR NAME).

• AUDIOGRAPHICS: audiographics in the context of distance education (mailto:LISTSERV@cln.etc.bc.ca Body: Sub Audiographics-L YOUR NAME).

• CBNVEE: educational environments using networked virtual reality software (mailto:listserv@mcmuse.mc.maricopa.edu Body: sub CBNVEE YOUR NAME).

• CMC: Computer-Mediated Communication Comserve Hotline (mailto:comserve@vm.its.rpi.edu Body: join CMC YOUR NAME).

• Commnet research: social scientific research on community networks by providing a forum for the discussion and sharing of research methods, resources, and data (mailto:listserv@calvin.dgbt.doc.ca Body: subscribe commnet-research YOUR NAME).

• Cybermind: the philosophical and psychological implications of subjectivity in cyberspace (mailto:majordomo@world.std.com Body: subscribe cybermind YOUR EMAIL ADDRESS).

• Cypherwonks: general cutting-edge ‘cyberspatial development’ projects such as in cryptography, digital cash, and ‘electronic democracy’ (mailto:Majordomo@Lists.EUnet.fi Body: sub cypherwonks YOUR NAME).

• EduTel: Comserve Hotline use of computers in education (mailto:comserve@vm.its.rpi.edu Body: join edutel YOUR NAME).

• EDTECH: Educational Technology (mailto:listserv@msu.bitnet Body: subscribe EDTECH YOUR NAME).

• Interpedia: a public-domain Internet encyclopedia (mailto:interpedia-request@telerama.lm.com Body: subscribe YOUR EMAIL ADDRESS).

• Interpedia Web: (http://www.hmc.edu/www/interpedia).

• IAMCRNet: Comserve Hotline International Association for Mass Communication Research Network (mailto:comserve@vm.its.rpi.edu Body: join IAMCRNet YOUR NAME).

• CNI-DIRECTORIES: Coalition for Networked Information Directories. (mailto:listproc@cni.org Body: sub CNI-DIRECTORIES YOUR NAME).

• FUTUREC: FutureCulture e-list (mailto:futurec@uafsysb.uark.edu Body: sub FUTUREC YOUR NAME).

• GOPHERJEWELS: sharing interesting gopher finds (mailto:Listproc@EINet.net Body: SUBSCRIBE GOPHERJEWELS YOUR NAME).

• HELP-NET: For people new to the Internet (mailto:listserv@templevm.bitnet Body: sub HELP-NET YOUR NAME).

• HYPEREDU: HYPERTEXT in EDUCATION, of hypertexts and hypermedia in education–advantages versus disadvantages (mailto:LISTSERV@ITOCSIVM.CSI.IT Body: SUB HYPEREDU YOUR NAME).

• HYTEL-L: announcements of new versions of the popular HYTELNET program (mailto:listserv@kentvm.kent.edu Body: sub HYTEL-L YOUR NAME).

• IECC: Intercultural Email Classroom Connections Lists (http://www.stolaf.edu/network/iecc.html).

• InterNIC Announce: announcements list for Internet Network Information Center (mailto:listserv@is.internic.net Body: sub announce YOUR NAME).

• InterNIC Net Resources: resources only, low volume (mailto:listserv@is.internic.net Body: sub net-resources YOUR NAME).

• InterNIC NICS: Internet Network Information Center Network Information Centers list (mailto:listserv@is.internic.net Body: sub YOUR EMAIL ADR nics).

• InterNIC Net Happenings: conferences, publications, political news, resources, newsletters, etc. (mailto:listserv@is.internic.net Body: sub net-happenings YOUR NAME).

• Interpedia: an online Internet encyclopedia (mailto:interpedia-request@telerama.lm.com Body: sub YOUR EMAIL ADDRESS).

• IPCT-L: The Interpersonal Computing and Technology List, issues in education, especially networking and computing. (mailto:listserv@guvm.georgetown.edu Body: sub IPCT-L YOUR NAME).

• IPCT Gopher: (gopher://guvm.ccf.georgetown.edu/11/LISTSERV).

• IRVC-L: foster, encourage, advance, and communicate research on virtual culture (mailto:listserv@byrd.mu.wvnet.edu Body: sub IRVC-L YOUR NAME).

• LM NET: A World-Wide Discussion Group For School Library Media People (mailto:listserv@suvm.syr.edu Body: sub LM NET YOUR NAME).

• MediaMOO: (telnet://purple-crayon.media.mit.edu:8888).

• MEDIA-L: Media in Education (mailto:listserv@bingvmb.bitnet Body: subscribe MEDIA-L YOUR NAME).

• MediaMOO: Netoric–Tuesday 8PM EST, Electronic Conferences, CMC Workshops (telnet://purple-crayon.media.mit.edu:8888).

• MBU: Megabyte University (Computers and Writing) (mailto:listserv@ttuvm1.ttu.edu Body: sub MBU-L YOUR NAME).

• MBU FAQ: Frequently Asked Questions about Megabyte University (http://daedalus.com/MBU/MBU.intro.html).

• Net-info: a discussion list for net resources (mailto:majordomo@dorsai.org Body: subscribe Net-info YOUR NAME).

• NETTRAIN: a discussion list about training/support of others in using the resources available on Bitnet and Internet (mailto:listserv@ubvm.cc.buffalo.edu Body: sub NETTRAIN YOUR NAME).

• Net Happenings Archive: search net-happenings mailing list archive (telnet://brsuser@gopher.cni.org).

• NEW-LIST: discusses new LISTSERV mailing lists. (mailto:listserv@ndsuvm1.bitnet Body: sub NEW-LIST YOUR NAME).

• NewJour-L: new electronic journals/newsletters (mailto:listserv@e-math.ams.org Body: sub NewJour-L YOUR NAME).

• NewNIR-L: announcements of new Network Information Retrieval and Online Public Access Catalogue Services (mailto:listserv@itocsivm.csi.it Body: sub NEWNIR-L YOUR NAME).

• NEWSLTR: newsletters currently posted to net-happenings, available only in digest form (mailto:listserv@vm1.nodak.edu Body: subscribe newsltr YOUR NAME).

• NII Regional: National Information Infrastructure (USA) Regional Project Development (mailto:majordomo@rain.org Body: subscribe niiregional-l).

• NII Teach: National Information Infrastructure (USA) and its role in education (mailto:NII-Teach-request@scholastic.com Body: subscribe NII-Teach YOUR NAME).

• NIR: Network Information Retrieval. List dedicated to studying networked information retrieval tools and processes. (mailto:mailbase@mailbase.ac.uk Body: sub NIR YOUR NAME).

• OCLC-JOURNALS: OCLC electronic journal publishing (mailto:listserv@oclc.org Body: sub OCLC-JOURNALS YOUR NAME).

• Online-news: online newspapers and magazines (mailto:Majordomo@marketplace.com Body: subscribe online-news YOUR BIO STATEMENT).

• UNITE: interface to all network services.ng. theme (mailto:mailbase@mailbase.ac.uk Body: join UNITE YOUR NAME).

• RESODLAA: distance education and open learning research (mailto:Listserver@usq.edu.au Body: sub RESODLAA YOUR NAME).

• RISKS: Forum on risks to the public in computers and related systems, ACM Committee on Computers and Public Policy (http://catless.ncl.ac.uk/Risks).

• RITIM-L: Research Institute for Telecommunications and Information Marketing discussion list (mailto:listserv@uriacc.uri.edu Body: sub RITIM-L YOUR NAME).

• Security: the public disclosure of bugs, system problems, viruses (mailto:problems-request@tdr.com).

• SCHOLAR: text analysis and natural language processing, funded by the Andrew W. Mellon Foundation (mailto:listserv@cunyvm.cuny.edu Body: sub SCHOLAR YOUR NAME).

• TECHNET: Technical Support for Education and Research (mailto:listserv@acadvm1.uottawa.ca Body: subscribe TECHNET YOUR NAME).

• Technomads: wireless, mobile data comm networks (mailto:technomads-request@ucsd.edu Body: sub technomads YOUR NAME).

• ThesisNet: criticism of works-in-progress related to all aspects of human interaction with and within cyberspace (mailto:thesisnet-request@seas.upenn.edu Body: sub THESISNET YOUR NAME).

• TIPSHEET: (mailto:listserv@wsuvm1.csc.wsu.edu Body: subscribe TIPSHEET).

• VPIEJ-L: Electronic publishing issues, especially Scholarly Electronic Journals (mailto:listserv@vtvm1.cc.vt.edu Body: sub VPIEJ-L YOUR NAME).

• Web4Lib: library-based World-Wide Web developers and managers (mailto:listserv@library.berkeley.edu Body: sub WEB4LIB YOUR NAME).

• WWWEDU: the use of the World Wide Web in education (http://k12.cnidr.org:90/wwwedu/9412/maillist.html).

• WWWORDER: creating World-Wide Web-based item or catalog order systems (mailto:listproc@einet.net Body: subscribe WWWORDER YOUR NAME).

• WWW talk archives: (http://gummo.stanford.edu/html/hypermail/archives.html).

• WWW courseware: new or existing educational applications of the World-Wide Web (mailto:www-courseware-request@eit.com Body: sub YOUR EMAIL ADDRESS).

• WWW literature: new or existing literary resources/works on the World-Wide Web (mailto:www-literature-request@eit.com Body: sub YOUR EMAIL ADDRESS).

6.4 Periodicals

These are periodicals with some coverage related to CMC, computer communications, networking or related issues.

• Communications Periodicals List: social science and communications periodicals, by Chandrasekhar Vallath (http://alnilam.ucs.indiana.edu:1027/sources/journals.html).

• Computer Publications: (http://www.webcom.com/∼ecola/tradepub.html).

• Computer/Tech Zines-Yahoo: Entertainment-Magazines- Computer and Technology (http://www.yahoo.com/Entertainment/Magazines/Computer and Technology/).

• Cyber Zines-Yahoo: Entertainment-Magazines-Cyberspace and Internet Culture (http://www.yahoo.com/Entertainment/Magazines/Cyberspace and Internet Culture/).

• Internet Magazines-Yahoo: (http://www.yahoo.com/Computers/Internet/Magazines/).

• Internet Newsletters-Yahoo: (http://www.yahoo.com/Computers/Internet/Newsletters/).

• Internet Press: A guide to electronic journals about the Internet (ftp://rtfm.mit.edu/pub/usenet/news.answers/internet-services/internet-press).

• ACN: Amateur Computerist Newsletter (http://www.cc.columbia.edu/∼hauben/home.html).

• ALAWON: American Library Association (ALA) Online Newsletter (mailto:listserv@uicvm.uic.edu Body: sub ALA-WO YOUR NAME).

• Bits and Bytes: a weekly electronic newsletter about miscellaneous high-tech issues (mailto:listserv@acad1.dana.edu Body: SUBSCRIBE bits-n-bytes).

• Boardwatch: Boardwatch Magazine BBS, voice line for information (phone:1-303-973-6038).

• Boardwatch email: Boardwatch Magazine BBS (mailto:jack.rickard@boardwatch.com).

• Byte: back issues (ftp://oak.oakland.edu/pub/misc/byte/).

• CJC: Canadian Journal of Communication (http://edie.cprost.sfu.ca/cjc/cjc-info.html).

• CETH Newsletter: Center for Electronic Texts in the Humanities, Princeton and Rutgers (mailto:ceth@zodiac.rutgers.edu).

• Chaos Corner: Dr. Chaos’s random and interesting things about computers, networks, and other things (mailto:chaos-request@pelican.cit.cornell.edu Body: sub YOUR NAME).

• Chaos Corner Archive: (ftp://pelican.cit.cornell.edu/pub/).

• Chaos Digest: an e-journal edited by Jean-Bernard Condat (mailto:condat@altern.com).

• CMC Magazine: Computer-Mediated Communication Magazine (ISSN 1076-027X) reports on people, events, applications, and research related to computer-mediated communication (CMC); publishes news, features, essays, and research reports about the phenomenon of human communication and information retrieval on global networks (http://www.rpi.edu/∼decemj/cmc/mag/current/toc.html).

• CMP Publications: a variety of information technology publications (http://techweb.cmp.com/techweb/default.html).

• Communicator: publication of The University of Michigan School of Information and Library Studies, Information and Library Studies Student Association (http://http2.sils.umich.edu/ILSSA/HomePage.html).

• Connect: A Telecomputing Magazine for All Modem Users, covers online services, Internet/Usenet (mailto:pegasus@cyberspace.org).

• Cook Report($): a monthly newsletter on NREN, National Information Infrastructure, and Internet commercialization (mailto:cook@path.net Body: Please send Cook Report info.).

• Convergence: entertainment in the age of the information super-highway (mailto:AEnigma@Netcom.com Body: Please send Convergence FAQ).

• CPU: electronic journal for computer industry workers (mailto:listserv@cpsr.org Body: sub CPSR-CPU YOUR NAME).

• CPU Archives: (ftp://etext.archive.umich.edu/pub/CPSR/work/).

• Cu Digest web: The Computer Underground Digest (http://sun.soci.niu.edu/∼cudigest/).

• Current Cites: A journal which provides citations and brief annotations for articles from 30 journals in networks and information and computer technology (ftp://ftp.lib.berkeley.edu/pub/Current.Cites/).

• Cyberkind: cyber-related art and literature (http://sunsite.unc.edu/shannon/ckind/title.html).

• CyberNews: global student-run media (print, radio, tv, etc...) connectivity, by Richard Linton Ginn (http://www.hmc.edu/www/people/teverett/cybernews/Home.html).

• Cybernetics Digest: converging information technologies ($) (http://www.pscu.com).

• Cyberwire: Brock N. Meeks’ CyberWire dispatch–news and views on many cyber-related issues (http://cyberwerks.com:70/1/cyberwire).

• DECNEWS: from Digital Equipment Corporation’s Education Business Unit for the education and research communities worldwide (mailto:listserv@ubvm.cc.buffalo.edu Body: sub DECNEWS YOUR NAME).

• DISTED: Online Chronicle of Distance Education and Communication (mailto:listserv@alpha.acast.nova.edu Body: sub DISTED YOUR NAME).

• EDUPAGE: summary of information technology news (mailto:listserv@bitnic.educom.edu Body: sub EDUPAGE YOUR NAME).

• EDUPAGE Web Archive: (http://educom.edu/edupage.old/.index.html).

• Effector Online: Electronic Frontier Foundation Newsletter (ftp://ftp.eff.org/pub/EFF/newsletters/Index).

• EJC/REC: Electronic Journal of Communication/La Revue Electronique de Communication Covers communication theory, research, practice, and policy (mailto:comserve@vm.its.rpi.edu Body: join EJCREC YOUR NAME).

• EPIN: Electronic Public Information Newsletter (mailto:epin@access.digex.com Body: Please send EPIN info.).

• EJournal: Concerned with implications of electronic networks and texts (mailto:listserv@albany.bitnet Body: sub EJRNL YOUR NAME).

• EJournal Archive: back issues of EJournal (ftp://ftp.hanover.edu/pub/ejournal).

• EJVC: Electronic Journal on Virtual Culture, communicate scholarly thought on aspects of virtual [CMC] culture. (mailto:listserv@kentvm.kent.edu Body: sub EJVC-L YOUR NAME).

• EJVC Archive: (ftp://byrd.mu.wvnet.edu/pub/ejvc/).

• Fed Comm Law J: Federal Communications Law Journal Web Page (http://www.law.indiana.edu:80/fclj/fclj.html).

• GlasNews: a quarterly publication on East-West contacts in communications - including journalism, telecommunications, advertising and public relations, by Art Pattison Communications Exchange Program, based in Seattle (http://solar.rtd.utk.edu/friends/news/glasnews/master.html).

• GNN: Global Network Navigator (mailto:info@gnn.com Body: Please send GNN info.).

• GNN NetNews: (http://nearnet.gnn.com/gnn/news/index.html).

• HOTT-LIST: (Hot Off The Tree) advances in computer, communications, and electronics technology (mailto:listserv@ucsd.edu Body: SUBSCRIBE HOTT-LIST).

• Hotwired: online magazine ’more signal less noise,’ from Wired Magazine (’imminent’ as of Oct 1994) (http://www.hotwired.com).

• HPC SELECT NEWS: The Electronic News Magazine for High-Performance Computing (mailto:sub@hpcwire.ans.net Body: Select 601).

• In Perspective: the WilTel magazine, explores a broad range of telecommunications issues (http://www.wiltel.com/perspect/homepers.html).

• Information and Computation: theoretical computer science (http://theory.lcs.mit.edu/∼iandc/).

• INFOBITS: electronic service of the Institute for Academic Technology’s Information Resources Group (mailto:listserv@gibbs.oit.unc.edu Body: sub INFOBITS YOUR NAME).

• InfoCycle: technology management (mailto:jwsmith@io.org).

• InfoCycle web: (http://www.io.org/∼jwsmith).

• Info Policy Online: newsletter of the Information Industry Association (mailto:iiaipo-request@his.com Body: subscribe YOUR NAME).

• INFOSYS: The Electronic Newsletter for Information Systems (mailto:listserv@american.edu Body: sub INFOSYS YOUR NAME).

• Interactive Age: the official newspaper of the interactive industry, covering the convergence of the telecommunications, cable, computer and entertainment industries (http://www.cmp.com/interactive-age/mission.html).

• Internaut: a guide and discussion of of online resources (http://www.zilker.net/users/internaut/index.html).

• Internet Informer: Information from the material and electronic worlds, donation-supported (mailto:Informer@Cris.Com).

• Internet Monthly Report: monthly communications to the Internet Research Group (ftp://nic.merit.edu/internet/newsletters/internet.monthly.report/).

• IRN: Internet Resources Newsletter, a FREE WWW newsletter produced by Heriot-Watt University Library (http://www.hw.ac.uk/libWWW/irn/irn.html).

• InterText: an electronically distributed magazine of fiction (ftp://network.ucsd.edu/intertext/).

• JASBITS: A newsletter of miscellaneous net information, editor James A. Squires (mailto:jsquires@cerf.net Subject: JASBITS SUBSCRIBE Body: YOUR NAME how you knew).

• JAUC: The Journal Of American Underground Computing (ISSN 1074-3111)

(ftp://etext.archive.umich.edu/pub/Zines/JAUC/).

• JCMC: the Journal of Computer-Mediated Communication (http://www.huji.ac.il/www jcmc/jcmc.html).

• JEP: the Journal of Electronic Publishing (http://www.press.umich.edu/jep/).

• LI Newswire: Learned Information’s news and views shaping the world of Online Information (http://info.learned.co.uk:70/1s/newswire).

• Media West Magazine: (http://wimsey.com/Media Wave/).

• Meta Magazine: news and editorial relevant to events and issues shaping, and being shaped by, the net community (ftp://ftp.netcom.com/pub/mlinksva/meta.html).

• Mondo2000 archive: (gopher://gopher.well.sf.ca.us/11/Publications/MONDO).

• NetComTalk: a quarterly journal of the College of Communications at Boston University (http://web.bu.edu/COM/communication.html).

• Netweaver: The Newsletter of the Electronic Networking Association (mailto:comserve@vm.its.rpi.edu Body: send Netweave Winter91).

• Nets and Comm: Networks and Community, a newsletter of FUTURE DATA (mailto:samsam@vm1.yorku.ca).

• Nets and Comm archive: back issues of Networks and Community, a newsletter of FUTURE DATA (gopher://gopher.nlc-bnc.ca/11gopher$root pub%3a%5bnetcomm%5d).

• Netsurfer Digest: news about the Web, new resources and software (http://www.netsurf.com/nsd/nsd.backissues.html).

• IBJ Gopher: Internet Business Journal Gopher (gopher://gopher.fonorola.net/11/Internet%20Business%20Journal).

• Internet Review: An ‘irregular and on-line journal of new and interesting stuff on the net’ (gopher://gopher.msen.com).

• Internet World: a paper journal from Mecklermedia (mailto:meckler@jvnc.net Body: Please send Internet World info.).

• Legion of Doom Tech J: (ftp://ftp.eff.org/pub/Publications/CuD/LOD/).

• LIBRES: Library and Information Science Electronic Journal (ISSN 1058-6768)

(ftp://cc.curtin.edu.au/LIB-RESEARCH/).

• LI Newswire: covers electronic information and online activities (http://info.learned.co.uk).

• Link Letter: Merit/NSFNET newsletter. (ftp://nic.merit.edu/newsletters/linkletter).

• Lynx Magazine: (http://www.cityscape.co.uk/lynx/).

• Matrix News: (monthly online newsletter, Covers crossnetwork issues. (mailto:mids@tic.com Body: Please send Matrix News info.).

• Matrix News some issues: (ftp://quake.think.com/pub/mids/matrix news).

• MC Journal: The Journal of Academic Media Librarianship (mailto:listserv@ubvm.cc.buffalo.edu Body: sub MCJRNL YOUR NAME).

• Meta: net community interests–cryptography, electronic publishing, free software, internet, privacy (ftp://ftp.netcom.com/pub/mlinksva).

• NETNEWS: newsletter for network resources by Metronet (St. Paul, MN)

(mailto:listserv@ndsuvm1.bitnet Body: sub NNEWS YOUR NAME).

• PAC-Hiway: regarding the development of the information highway (mailto:listprocessor@cunews.carleton.ca Body: Subscribe Pac-Hiway YOUR NAME).

• PACSR: The Public-Access Computer Systems Review (mailto:listserv@uhupvm1.uh.edu Body: sub PACS-P YOUR NAME).

• PACSR Index: The Public-Access Computer Systems Review Index (mailto:listserv@uhupvm1.uh.edu Body: get index pr f=mail).

• PACS Gopher: (gopher://info.lib.uh.edu/11/articles/e-journals/uhlibrary).

• PC Magazine: (http://www.pcmag.ziff.com).

• PC Week Labs: (http://www.pcmag.ziff.com).

• Phrack: hackers (ftp://ftp.eff.org/pub/).

• Postmodern Culture: An Electronic Journal of Interdisciplinary Criticism (http://jefferson.village.virginia.edu/pmc/contents.all.html).

• REACH: Research and Education Applications of Computers in the Humanities Newsletter (ftp://ucsbuxa.ucsb.edu/hcf/).

• Red Rock Eater: Red Rock Eater (RRE) News Service, Phil Agre’s periodic messages about Internet (mailto:rre-request@weber.ucsd.edu Subject: subscribe YOUR NAME).

• Red Rock Eater RRE web: Red Rock Eater News Service, Phil Agre’s periodic messages about Internet (http://communication.ucsd.edu/pagre/rre.html).

• SPEED: an electronic journal devoted to the critical investigation of technology, media, and society (gopher://alishaw.ucsb.edu/11/.speed).

• Surfpunk Technical Journal: ‘a dangerous multinational hacker zine originating near BARRNET in the fashionable western arm of the northern California matrix’ (http://www.acns.nwu.edu/surfpunk/).

• Telektronikk: (http://www.nta.no/telektronikk/4.93.html).

• TER: Telecommunications Electronic Reviews (gopher://info.lib.uh.edu:70/11/articles/e-journals/lita).

• TNO: Network Observer, a free on-line newsletter about networks and democracy (http://communication.ucsd.edu/pagre/tno.html).

• TNS: Telelearning Network Synthesizer, Edited by Tom Copley (gopher://gigantor.arlington.com:70/11/tns).

• TidBITS: a free weekly electronic publication that reports on interesting products and events in the computer industry (mailto:listserv@ricevm1.rice.edu Body: sub TIDBITS YOUR NAME).

• TOW Monitor: The Online World Monitor newsletter, companion to The Online World book, both by Odd de Presno (mailto:listserv@vm1.nodak.edu Body: get tow monitor).

• Tradewinds: A monthly round-up of Internet coverage in trade and industry magazines (gopher://ftp.std.com/11/periodicals/TRADEWINDS).

• Virtual Mirror: (http://mirror.wwa.com/mirror/).

• Voorhees Reports: court and regulatory developments involving telecommunications, intellectual property, and information (mailto:markvoor@phantom.com Body: Please send Voorhees Reports info.).

• The WELL: The Whole Earth Electronic Link (gopher://gopher.well.sf.ca.us).

• The WELL Web: The Whole Earth Electronic Link web (http://www.well.com).

• Voices/Net: (ftp://ftp/Library/Zines/).

• Washington/Telecom: Washington Telecom Newswire, events in Washington DC USA affecting telecommunications (http://wtn.com/wtn/wtn.html).

• Web Week: (http://www.mecklerweb.com/ww/).

• Web Word: (http://www.euro.net/innovation/Web Word Base/Web.WordHP.html).

• WEBSter: (http://www.tgc.com/webster.html).

• Wired: a magazine of cyberspace and technology (mailto:infobot@wired.com Body: get index).

• Wired Magazine: (http://www.wired.com).

6.5 Usenet

Usenet newsgroups are sometimes a rich source of information about the Internet, networks, and CMC issues. This section lists newsgroups in which topics related to networks, the Internet, or CMC are discussed. Watch for FAQ = periodic posting of Frequently Asked Questions in these groups. —– Collections

• Computer/Comm: Computer and Communications related Usenet news groups, list from Advanced Telecommunications Program at Lawrence Livermore National Laboratory, emphasis on telecommunications and technical issues, well-organized with links to FAQ as well as group (http://www-atp.llnl.gov/atp/telecom-usenet.html).

• Data comm/Networking: (http://www.racal.com/dcom/dcomgroups.html). Selected newsgroups

• Beginners: for people new to USENET (news:news.announce.newusers).

• Computer Archives: (news:comp.archives).

• Computer BBS: (news:comp.bbs.misc).

• Cu-Digest: Computer Underground Digest (news:comp.society.cu-digest).

• Cyberpunk: cyberpunk/cyberspace culture and literature (news:alt.cyberpunk).

• Email: (news:comp.mail.misc).

• FAQs: Frequently Asked Questions (and answers) (news:news.answers).

• Gopher: (news:comp.infosystems.gopher).

• Groupware (CSCW): (news:comp.groupware).

• Hackers: (news:alt.hackers).

• ISDN: (news:comp.dcom.isdn).

• Information Systems: (news:comp.infosystems).

• Internet Access: (news:alt.internet.access.wanted).

• Internet BBS: (news:alt.bbs.internet).

• Internet Best: (news:alt.best.of.internet).

• Internet Culture: (news:alt.culture.internet).

• Internet Library: (news:comp.internet.library).

• Interent.media-coverage: alt.internet.media-coverage Charter and FAQ (http://www.dorsai.org/∼tristan/AIMC/).

• Internet Protocols: (news:comp.protocols.tcp-ip).

• Internet Relay Chat: (news:alt.irc).

• Internet Services: (news:alt.internet.services).

• Internet Talk Radio: (news:alt.internet.talk-radio).

• Lists of Information: (news:news.lists).

• MUDS: Multiple User Dialogues (news:rec.games.mud.announce).

• New Newsgroups: (news:news.announce.newgroups).

• Newsgroups: (news:news.groups).

• PostScript Net Maps: (news:news.lists.ps-maps).

• Technical Reports: (news:comp.doc.techreports).

• Telecomm: (news:clari.nb.telecom).

• Telecomm: (news:comp.dcom.telecom).

• Undernet: an alternative to the current EFNet IRC system (news:alt.irc.undernet).

• USENET Culture: (news:alt.culture.usenet).

• USENET Humor: (news:alt.humor.best-of-usenet).

• USENET Recovery: (news:alt.usenet.recovery).

• USENET University: (news:alt.uu.future).

• WAIS: (news:comp.infosystems.wais).

• WWW Usenet Groups: (http://www.halcyon.com/grant/Misc/web-groups.html).

7 ORGANIZATIONS 60

7 ORGANIZATIONS

These are organizations that relate to Internet, CMC, or computing, communications, telecommunications, or other related activities, through the fields of study they represent or through their activities and research.

7.1 Academic

— Lists of programs in CMC and related fields

• Communications Programs: list compiled by Don Gale of The University of Colorado, Bruce Klopfenstein of Bowling Green State University, and C. Vallath of Indiana University (http://alnilam.ucs.indiana.edu:1027/sources/programs.html).

• Journalism/Communications: by Gary Ritzenthaler and Sherwood Lawrence (http://www.jou.ufl.edu/commres/jouwww.htm).

— Departments or organizations

• ACA: American Communication Association (http://cavern.uark.edu/comminfo/www/ACA.html).

• Annenberg/USC: Annenberg School for Communication at the University of Southern California (gopher://cwis.usc.edu/11/University Information/Academic Departments/Annenberg).

• BGSU/Telecom: Bowling Green State University, Ohio, Department of Telecommunications (http://www.bgsu.edu/departments/tcom/).

• Brussels: Brussels University, Sercie Telematique et communication (http://sun4.iihe.ac.be).

• CCCC: Computerized Conferencing and Communications Center at New Jersey Institute of Technology (NJIT) (http://it.njit.edu/njIT/Department/CCCC/default.html).

• CICA: The Center for Innovative Computer Applications at Indiana University (http://www.cica.indiana.edu).

• CISM-UT Austin: Center for Information Systems Management, College and Graduate School of Business, University of Texas at Austin (http://cism.bus.utexas.edu).

• CMC Studies Center: an organization devoted to the study of computer-mediated communication research, resource-sharing, and discussion (http://www.rpi.edu/∼decemj/cmc/center.html).

• CMU INI: Information Networking Institute at Carnegie Mellon University, the nation’s first research and education center devoted to Information Networking (http://www.cs.cmu.edu:8001/afs/andrew.cmu.edu/inst/ini/www/INI-home.html).

• Columbia U–telecom research: Columbia University Center for Telecommunications Research (CTR) (http://www.ctr.columbia.edu/CUCTR Home.html).

• CUI: Centre Universitaire d’Informatique, l’Universite de Geneve (http://cuiwww.unige.ch/home.html).

• ESCA: European Speech Communication Association (http://ophale.icp.grenet.fr/esca/esca.html).

• St. Gallen-Info Mgt: Institute for Information Management, University of St.Gallen, Switzerland; information management in business and administration, including groupware, work-flow, document management, information management, process engineering and CASE, information markets and communication (http://www-iwi.unisg.ch).

• GLOCOM–Japan: Center for Global Communications, GLOCOM International University of Japan (http://www.glocom.ac.jp).

• ITI: The Information Technology Institute, University of Salford, United Kingdom (http://www.salford.ac.uk/docs/depts/iti/homepage.html).

• IWI; Institute for Information Management, University of St.Gallen, Switzerland: (http://www-iwi.unisg.ch).

• Journalism/Communications sites: journalism and mass communications (http://www.jou.ufl.edu/commres/jouwww.htm).

• KTH: Institutionen for Teleinformatik, KTH, Department of Teleinformatics, The Royal Institute of Technology, Sweden (http://www.it.kth.se).

• Linz: Johannes Kepler University of Linz Department of Information Systems (http://www.ifs.uni-linz.ac.at/tnf/inf/ifs/home.html).

• Nottingham–comm research: Communications Research Group, Nottingham, United Kingdom (http://web.cs.nott.ac.uk).

• Michigan Tech: CCLI (Center for Computer-assisted Language Instruction) at Michigan Technological University (gopher://gilligan.hu.mtu.edu/11).

• MTU HU: Michigan Technological University humanities department (http://www.hu.mtu.edu).

• MIT: MIT Program on Communications and Media Policy (http://far.mit.edu).

• Newhouse: Newhouse School of Public Communications at Syracuse University (http://spc.syr.edu/0/Newhouse/NEWHOUSE.html).

• Nova U-CCIS: Center for Computer and Information Sciences (mailto:cciscat@alpha.acast.nova.edu Body: anything).

• NJIT-CCCC: Computerized Conferencing and Communications Center at New Jersey Institute of Technology (NJIT) (http://it.njit.edu/njIT/Department/CCCC/default.html).

• Open U: Open University, British distance teaching, open access university (http://pepper.open.ack.uk).

• Open U-iCDL: International Centre for Distance Learning, in the United Kingdom on the campus of the Open University, disseminating distance education information worldwide (http://acs-info.open.ac.uk/info/other/ICDL/ICDL-Facts.html).

• Syracuse: School of Information Studies (gopher://cwis.syr.edu/11/othergophers/ist).

• Telecom Paris: part of a larger institution for graduate-level instruction in telecommunications that links the Ecole Nationale Superieure des Telecommunications and the Institut National des Telecommunications (http://www.enst.fr).

• UCI-ICS: Information and Computer Science, University of California, Irvine (http://www.ics.uci.edu/ICShome.html).

• UCSD-Comm: University of California San Diego Communication Department (http://communication.ucsd.edu).

• UHM/Comm: University of Hawaii at Manoa Department of Communication (http://www.soc.hawaii.edu/con/com/newcom.html).

• U of FL–comm research: Communication Research Center, College of Journalism + Communication at the University of Florida (http://www.jou.ufl.edu/commres/crc.htm).

• USC ISI: University of Southern California (USC) Information Sciences Institute (ISI) (http://www.isi.edu).

• ZZT: Telecommunications Department, Zagreb, capital of Croatia (http://tjev.tel.etf.hr/zzt/enstart.html).

7.2 Research

—-lists:

• Net Research sites list: emphasis on technical aspects (http://netlab.itd.nrl.navy.mil/onr.html).

—-sites:

• ARPA-USA: USA Advanced Research Projects Agency (http://ftp.arpa.mil).

• CCAE-Lisbon: Centro de Comunicagues em Ambientes Empresariais, Lisbon, Portugal (http://www.inesc.pt/inesc/centros/ccae.html).

• CCD-Brookhaven: Computing and Communications Division of Brookhaven National Laboratory (http://docserver.bnl.gov/com/www/default.html).

• CCSR-Beckman: Center for Complex Systems Research, at the Institute for Advanced Science and Technology at the University of Illinois at Urbana-Champaign (http://www.ccsr.uiuc.edu/CCSRHome.html).

• CCRC-Washington U: Washington University (St. Louis), Computer and Communications Research Center (http://dworkin.wustl.edu).

• CIT-Stanford: Center for Information Technology, a laboratory operated by Stanford University, the encoding, storage, communication, manipulation, and use of information in digital form (http://logic.stanford.edu/cit/cit.html).

• CITI-Michigan: The University of Michigan’s Center for Information Technology Integration (http://www.citi.umich.edu).

• CERN: European Laboratory for Particle Physics, Geneva, Switzerland, (and birthplace of the World-Wide Web) (http://www1.cern.ch).

• CNET-France: Centre National d’Etudes des Telecommunication, France Telecom (http://www.cnet.fr).

• CRA: Computing Research Association, an association of more than 160 North American academic departments of computer science and computer engineering, industrial laboratories engaging in basic computing research and affiliated professional societies (http://cra.org).

• CRC-Ottowa: Communications Research Centre, Ottawa, Canada (http://www.crc.doc.ca/index.html).

• CRL-Japan: Communications Research Laboratory of the Ministry of Posts and Telecommunications is the sole national institute responsible for the study of telecommunications technologies, radio science, and radio applications in Japan (http://www.crl.go.jp).

• CSL-SRI: Computer Science Laboratory (CSL) at SRI International (http://www.csl.sri.com).

• CTRG-U of CO: Collaboration Technology Research Group (group user interfaces, distributed editors, workflow systems, cooperatively grown information systems), University of Colorado (http://www.cs.colorado.edu/homes/carlosm/public html/ctrg.html).

• CSHC-U of MD: a catalyst for commercial communications services worldwide via a hybrid information infrastructure (http://wais.isr.umd.edu/CCDS/homepage.html).

• CWI: Centrum voor Wiskunde en Informatica, Centre for Mathematics and Computer Science, Amsterdam, Netherlands (http://www.cwi.nl/default.html).

• ERCIM: European Research Consortium for Informatics and Mathematics, fosters collaboration with the European research community and industry (http://www-ercim.inria.fr).

• ERG-Aberdeen: Electronics Research Group, University of Aberdeen, United Kingdom (http://www.erg.abdn.ac.uk/misc/personnel.html).

• HCRC-Edinburgh: Human Communication Research Centre at the University of Edinburgh, United Kingdom (http://www.cogsci.ed.ac.uk/hcrc/home.html).

• ICBL-Edinburgh: Institute for Computer Based Learning, Heriot-Watt University, Edinburgh, United Kingdom (http://www.icbl.hw.ac.uk).

• ICSI-Berkeley: The International Computer Science Institute (http://http.icsi.berkeley.edu).

• IHNS-Japan: Institute for Hyper Network Society, articulate the fundamental structure and dynamics of hypernetwork society (http://www.glocom.ac.jp/IHNS/intro.html).

• IIASA-Austria: International Institute for Applied Systems Analysis–Global Environmental Change; Global Economic and Technological Transitions; Systems Methods for the Analysis of Global Issues (http://www.iiasa.ac.at/welcome.html).

• INRIA-France: Institut National de Recherche en Informatique et en Automatique, a French research institute in computer-science (http://www.inria.fr).

• IRIS-Brown: Brown University’s Institute for Information Research and Scholarship (http://www.iris.brown.edu/iris).

• ISI-USC: University of Southern California (USC) Information Sciences Institute (ISI) (http://www.isi.edu).

• ISR: Institute for Systems Research, focusing on systems engineering research, at the University of Maryland and Harvard University (http://www.isr.umd.edu).

• ISRI-UNLV: Information Science Research Institute, University of Nevada, Las Vegas (http://www.isri.unlv.edu).

• ISS-Singapore: Institute of Systems Science, National University of Singapore (http://www.iss.nus.sg).

• LLNL: Lawrence Livermore National Laboratory (LLNL) (http://www-atp.llnl.gov).

• NECTEC-Thiland: National Electronics and Computer Technology Center (Thailand), a research organization operated under the National Science and Technology Development Agency (http://www.nectec.or.th/nectec.html).

• NRL-AITB: Naval Research Laboratory in Washington, D.C.’s Advanced Information Technology Branch (http://www.ait.nrl.navy.mil/home.html).

• OSF RI: Open Software Foundation Research Institute (http://riwww.osf.org:8001).

• OSI-LAB: OSI (Open Systems Interconnecttion)-Protocol Testing Laboratory of the SWISS PTT TELECOM located in Bern, Zurich and Fribourg, Switzerland (http://www.osilab.ch).

• RAND: (http://www.rand.org).

• RENATER-Paris: Reseau National de telecommunications pour la Technologie, l’Enseignement et la Recherche (http://www.urec.fr/Renater/).

• RITIM: The Research Institute for Telecommunications and Information Marketing (http://www.wiltel.com/ritim/ritim.html).

• RLG: Research Libraries Group, Inc., a not-for-profit membership corporation of universities, archives, historical societies, museums, and other institutions devoted to improving access to information that supports research and learning; includes access to Eureka (search system), RLIN (library support system), and Zephyr (Z39.50 services), database, citation files, Ariel document delivery (http://www-rlg.stanford.edu/welcome.html).

• RPCP-MIT: Research Program on Communications Policy (http://farnsworth.mit.edu).

• Sandia: Sandia National Laboratories (http://www.sandia.gov).

• SFI: Santa Fe Insitute, complex systems, including global computer networks (http://santafe.edu).

• SLS-MIT: Spoken Language Systems Group at the Laboratory for Computer Science at the Massachusetts Institute of Technology (http://sls-www.lcs.mit.edu).

• SRI: a non-profit organization, one of the world’s largest contract research firms, technological and business problems (http://www.sri.com).

• SRC-Digital: Digital Equipment Corporation’s Systems Research Center (http://www.research.digital.com/SRC/home.html).

• UTRC: United Technologies Research Center (http://utrcwww.utc.com/UTRC/General/UTRCGen.html).

7.3 Commercial

These are commercial organizations that deal with Internet, CMC, or computing, communications, telecommunications, or other related activities.

7.3.1 Lists

• Access Providers–Yahoo: Internet Access Providers list from Yahoo (http://www.yahoo.com/Business/Corporations/Internet Access Providers/).

• Applications–commerce:

(http://www.rpi.edu/Internet/Guides/decemj/icmc/applications-commerce.html).

• Business-indices: Business-Corporations-indicies, from Yahoo (http://www.yahoo.com/yahoo/Business/Corporations/indicies/).

• Computer and Comm Companies: Computer and Communication Company Sites on the Web, by James E. (Jed) Donnelley (http://www-atp.llnl.gov/atp/companies.html).

• Computer Companies: list from Web of Wonder (WOW), includes systems, consulting, hardware, software, other links (http://www.digimark.net/wow/companie/computer/index.html).

• DLIST: a list of dedicated line Internet providers, by Susan Estrada (mailto:dlist@ora.com Body: Please send DLIST).

• Gray Pages: information on business products and services through a Web interface (http://www.trinet.com/tgp/).

• Ho, Thomas: favorite Electronic Commerce WWW resources, includes information sources, links to articles, economic development, service/presence providers (http://www.engr.iupui.edu/∼ho/interests/commmenu.html).

• Indices: Business-Corporations-Indices, from Yahoo (http://www.yahoo.com/Business/Corporations/Indices/).

• IAP Catalog: Internet Access Provider Catalog (indexed by area code/country code), from Network-USA (http://www.netusa.net/ISP).

• Internet Consultants: CommerceNet’s directory of Internet consultants (products and services) (http://www.commerce.net/directories/consultants/consultants.html).

• Internet Providers: All Providers Alphabetically (ftp://sri.com/netinfo/internet-access-providers-alphabetical-listing.txt).

• Internet Providers Non-US: Non-US Internet Providers (ftp://sri.com/netinfo/internet-access-providers-non-us.txt).

• PDIAL: The Public Dialup Internet Access List, by Peter Kaminski (http://www.pdial.com).

• WWYP: World Wide Yellow Pages (http://www.yellow.com).

• WWW Service: companies providing various WWW services such as serving pages, authoring pages, etc., by Mary E. S. Morris of Finesse Liveware (ftp://ftp.einet.net/pub/INET-MARKETING/www-svc-providers).

7.3.2 Providers

See:

• Internet Access Providers-Yahoo:

(http://www.yahoo.com/Business/Corporations/Internet Access Providers/).

• National Internet Service Providers:

(http://www.commerce.net/directories/products/isp.usa.ix.html). Below is a partial list of commercial organizations that provide Internet connectivity or access.

• America Online: (phone:800-827-6364).

• BIX: (phone:800-695-4775).

• CONNECT: (phone:408-973-0110).

• CompuServe: (phone:800-848-8199).

• CompuServe: (http://www.compuserve.com).

• Delphi: (phone:800-695-4005).

• Delphi web: (http://www.delphi.com).

• Dialog: (phone:800-334-2564).

• Dow Jones: (phone:800-522-3567).

• EUnet: European Internet-related service organization (http://www.eu.net).

• GEnie: (phone:800-638-9636).

• INO: Intelligence Network Online, Inc. (mailto:info@intnet.net).

• Japan-IIJ: Internet Initiative Japan Inc. (http://www.iij.ad.jp).

• MCI Mail: (phone:800-444-6245).

• MSEN: MSEN offers Usenet and Internet Connectivity (gopher://gopher.msen.com).

• Netcom: (mailto:info@netcom.com).

• PCIX: Internet Service Provider, New England USA (http://www.pcix.com).

• PC-LINK: (phone:800-827-8532).

• PSI: Performance Systems International, dial-up Internet access (phone:703-620-6651).

• PSI: Performance Systems International (mailto:all-info@psi.com).

• PSI Web: (http://www.psi.net).

• Pipeline: NYC-based Internet gateway (mailto:infobot@pipeline.com Body: help).

• Prodigy: (phone:800-776-3449).

• Prodigy/Astranet: (http://www.astranet.com).

• Software Tool and Die: The World, online service for dialup access to the Internet (gopher://world.std.com).

• TMN: The Meta Network (mailto:info@tmn.com).

• The WELL: (mailto:support@well.sf.ca.us).

7.3.3 Consortia

• ACC: Alliance for Competitive Communications (formerly MFJ Task force) an ad-hoc coalition of the seven regional Bell operating companies (http://bell.com).

• CIX Web: Commercial Internet Exchange (http://www.cix.org/CIXhome.html).

• IFIP gopher: International Federation for Information Processing, a multinational federation of professional and technical organizations (or national groupings of such organizations) concerned with information processing (gopher://ietf.cnri.reston.va.us/00/ifip/ifip.info).

• IFIP web: International Federation for Information Processing, a multinational federation of professional and technical organizations (or national groupings of such organizations) concerned with information processing (http://www.dit.upm.es/∼cdk/ifip.html).

• IIA: Information Industry Association, 500 member companies in the generation, processing, distribution and use of information (mailto:iia.ipo@his.com).

• NIIP: National Industrial Information Infrastructure Protocol Consortium (http://www.niiip.org).

• NIIT: National Information Infrastructure Testbed (http://www.niit.org/niit/).

• NIPDE: National Initiative for Product Data Exchange (http://www.eeel.nist.gov/nipde/).

• NOMA: National Online Media Association (mailto:natbbs-request@echonyc.com).

• SGML Open: a non-profit, international consortium of providers of products and services, dedicated to accelerating the further adoption, application, and implementation of the Standard Generalized Markup Language (http://www.sgmlopen.org).

• Wireless: The Wireless Opportunities Coalition, a group of organizations and companies seeking to expand wireless communications development, manufacturing, and use (http://wireless.policy.net/wireless/wireless.html).

7.3.4 Developers

—– Lists

• Comm+Media: Business-Corporations-Communications and Media Services, from Yahoo (http://www.yahoo.com/yahoo/Business/Corporations/Communications and Media Services/).

• Internet Consulting-Yahoo: Internet-Consulting, from Yahoo (http://www.yahoo.com/Business/Corporations/Internet Consulting).

• Internet Presence-Yahoo: Internet-Presence-Providers, from Yahoo (http://www.yahoo.com/Business/Corporations/Internet Presence Providers/).

• WWW Consultants/Providers:

(http://hanksville.phast.umass.edu/∼bdelong/res/wwwexp.html). —– Selected Instances of consultants/marketers/markets/advertisers related to CMC.

• DCI: Digital Consulting-computer/telecommunications conferences (e.g. Distributed Computing World, etc) (http://www.oec.com/DCI/index.html).

• MIDS/TIC web: Matrix Information and Directory Services/Texas Internet Consulting, includes internet information (http://www.tic.com).

• Mecklermedia: publishers of Internet books and conferences (phone:800-MECKLER).

• Mecklerweb: (http://www.mecklerweb.com/home.htm).

• PIPEX: The Public IP Exchange (http://www.pipex.net).

• WAIS, Inc.: interactive on-line publishing systems and services to organizations (http://www.wais.com).

7.3.5 Products

• Cisco: manufactures and sells multiprotocol routers (http://www.cisco.com).

• Consensus: software support for collaboration, including groupware (ftp://www.consensus.com:8300).

• DEC: Digital Equipment Corporation, Marketing Home Page (ftp://gatekeeper.dec.com/info.html).

• IBM: International Business Machines Corporation (http://www.ibm.com).

• ICI: Internet Computer Index, lists the almost all of the Internet resources relating to PCs, Macintoshes, and Unix systems (http://ici.proper.com).

• Microsoft: research in fields of computer science that might be beneficial to Microsoft in the future (http://www.research.microsoft.com).

• Motorola: wireless communications, semiconductors and advanced electronic systems and services (http://www.mot.com).

• NCD: Network Computing Devices (http://www.ncd.com).

• NEC: (http://www.nec.co.jp).

• Novell: (http://www.novell.com).

• OCM: Online Computer Market, information about computer hardware, software and services (http://www.ocm.com).

• RNS: Rockwell Network Systems, high-performance, standards-based networking equipment (http://www.cmc.com).

• TMC: Thinking Machines Corporation (http://www.think.com).

• RSA: RSA Data Security, Inc., world leader in cryptography (http://www.rsa.com).

• RTZ: RTZ software, Cupertino, California, real-time multimedia conferencing products (ftp://ftp.netcom.com/pub/rtz/www/rtzhomepage.html).

• SCO Web Page: The Santa Cruz Operation (http://www.sco.com/index.html).

• Silicon Surf: Silicon Graphics’ SILICON SURF Home Page (http://www.sgi.com).

• Spry, Inc.: applications for LANs, AIR Mosaic, NetAccess(a computer industry business directory for Mosaic users) (http://www.spry.com).

• Sun Home Page: Sun Microsystems distributed computing technologies, products, and services (http://www.sun.com).

• Sun Microsystems: Archives, Newsgroups and More!

(http://sunsite.unc.edu/sun/inform/sun-info.html).

• VTLS Web Page: automation solutions for libraries and other information centers worldwide (http://www.vtls.com).

• Wildfire Communications: personal communication assistants for the mobile professional (http://www.utopia.com/companies/wildfire/home.html).

• Xerox PARC: Xerox Palo Alto Research Center (http://www.parc.xerox.com).

7.3.6 Telecom

—– Lists

• Telecom-Yahoo: Business-Corporations-Telecommunications

(http://www.yahoo.com/Business/Corporations/Telecommunications/). —– Selected instances

• Ameritech: (http://www.ameritech.com).

• AT+T: (http://www.att.com).

• AT+T Research: Web server for AT+T Bell Laboratories Research (http://www.research.att.com).

• Bellcore: (http://info.bellcore.com).

• Bell Atlantic: (http://www.ba.com).

• British Telecom: (http://www.bt.net).

• MCI: (http://www.mci.com).

• MPT-Japan: Ministry of Posts and Telecommunications, Japan (http://www.mpt.go.jp).

• NTT-Japan: Nippon Telegraph and Telephone (http://www.ntt.jp/index.html).

• Pacific Bell: (http://www.pacbell.com).

• Southwestern Bell: (http://www.sbc.com).

• Sprint: (http://www.sprintlink.net).

• Tampere: Tampere Telephone Company, Tampere, Finland (http://www.tpo.fi).

• US West: (http://www.uswest.com).

• WilTel: (http://www.wiltel.com).

7.4 Internet

—– These are organizations that work to develop the Internet.

• Internet WG: Internet working groups (http://netlab.itd.nrl.navy.mil/Internet.html).

• IEPG: Internet Engineering and Planning Group, a group of Internet service providers working together to promote a global Internet (http://info.aarnet.edu.au/iepg/).

• IETF ftp: Internet Engineering Task Force (ftp://cnri.reston.va.us/ietf/).

• IETF web: (http://www.ietf.cnri.reston.va.us/home.html).

• IS: Internet Society (http://info.isoc.org).

• Merit: NSFNet Engineering (http://rrdb.merit.edu).

• W3C: the World Wide Web Consortium, supports the stable evolution of the World Wide Web and its protocols (http://www.w3.org/hypertext/WWW/Consortium/).

• Working Groups: internet working groups (http://netlab.itd.nrl.navy.mil/Internet.html).

7.5 Network

These are consortia, collections of networks, or national or regional networking organizations or groups. • EARN: European Academic Research Network (ftp://ftp.earn.net).

• EARN Web: (http://www.earn.net/welcome.html).

• NEARnet: New England Academic and Research Network (ftp://nic.near.net/docs/).

• NYSERNet: New York State Education and Research Network (ftp://nysernet.org/pub/resources/guides/).

• NYSERNet web: New York State Education and Research Network (http://nysernet.org/nysernet.html).

• RARE: Reseaux Associes pour la Recherche Europeenne, European research networks and users (ftp://ftp.rare.nl/rare/).

• RARE docs: document store of Reseaux Associes pour la Recherche Europeenne, European research networks and users (http://www.rare.nl).

• RIPE: Reseaux IP Europeen, ensures coordination to operate a pan-European IP network. (ftp://ftp.ripe.net/ripe/).

• RIPE Gopher: Reseaux IP Europeens, collaborative organization which consists of European Internet service providers (gopher://gopher.ripe.net).

• RIPE Web: Reseaux IP Europeens (http://www.ripe.net).

• SuperJANET: Super Joint Academic Network, an advanced high speed optical fibre network linking a large number of sites within the academic community (http://gala.jnt.ac.uk).

• SURAnet: Southeastern Universities Research Association Network (ftp://ftp.sura.net/pub/README).

• TARENA: Trans-European Research and Education Networking Association, international information and telecommunications infrastructure; fromed from the merging of RARE and EARN (Oct 1994) (http://www.terena.nl).

7.6 Non-profit

These are organizations which work to foster networking or advocate for issues related to CMC, communications, or technology.

• APT: Alliance for Public Technology, ’a Washington, DC based nonprofit, tax exempt coalition of public interest groups and individuals whose goal is to foster broad access to affordable, usable information and communication services and technology’ (http://apt.org/apt.html).

• ASIS: American Society for Information Science (mailto:asis@cni.org).

• ASIS students: Web page for the Bay Area (San Francisco, USA) Student Chapter of American Society for Information Science, includes links to other chapters (http://ranga.berkeley.edu/ASIS/asis-sc.html).

• CCN: The Center for Civic Networking, applying information infrastructure to the broad public good (mailto:ccn@civicnet.org).

• CCN: Civic Network is an evolving collection of online resources dedicated to supporting civic life and civic participation (http://www.civic.net:2401).

• CCN Archive: (ftp://ftp.std.com/ftp/amo/civicnet/).

• CIOS: Communication Institute for Online Scholarship, a non-profit organization for online communication scholarship (mailto:comserve@vm.its.rpi.edu Body: help topics associates).

• CIOS gopher: Communication Institute for Online Scholarship, a non-profit organization for online communication scholarship; gopher server (gopher://cios.llc.rpi.edu).

• CITS: Center For Information, Technology, and Society (ftp://ftp.std.com/customers/nonprofits/CITES).

• CIX: Commercial Internet eXchange Association, non-profit trade association promoting communications services. (ftp://cix.org/cix/README).

• CIX Web: Commercial Internet Exchange (http://www.cix.org/CIXhome.html).

• CME: Center for Media Education, promotes democratic potential of electronic media. (mailto:cme@access.digex.net Subject: ”CME Sub”).

• CNI: Coalition for Networked Information, promotes creation of and access to information resources in networked environments (ftp://ftp.cni.org/CNI/).

• CNI Web: (http://www.cni.org/CNI.homepage.html).

• CNIDR Web Page: Coalition for Networked Information Discovery and Retrieval Home Page (http://cnidr.org/welcome.html).

• CNRI: Corporation for National Research Initiatives, non-profit organization dedicated to developing networked information technology (http://www.cnri.reston.va.us).

• CoSN: Consortium for School Networking, helps educators and students access information and communications resources (mailto:cosn@bitnic.bitnet Body: Please send CoSN info.).

• CoSN Gopher: Consortium for School Networking (gopher://cosn.org).

• CPI: Coalition for Public Information, information developers, providers and educators dedicated to ensuring public access to the evolving electronic information infrastructure in Canada (gopher://resudox.net:1994).

• CPSR: Computer Professionals for Social Responsibility, alliance of computer professionals who discuss the impact of computer technology on society. (ftp://ftp.cpsr.org/cpsr/).

• CPSR Gopher: Computer Professionals for Social Responsibility, alliance of computer professionals who discuss the impact of computer technology on society. (gopher://cpsr.org).

• CPSR Web: Computer Professionals for Social Responsibility, alliance of computer professionals who discuss the impact of computer technology on society. (http://www.cpsr.org/home).

• CREN: supports low-cost access to worldwide electronic networking for education and research (gopher://info.educom.edu).

• CSD: The Center for Software Development (http://www.center.org/csd/home.html).

• CSF: Communications for a Sustainable Future (gopher://csf.colorado.edu).

• ECHO: European Commission Host Organisation (telnet://echo@echo.lu).

• EDUCOM: dedicated to the study and access of information technology in higher education and integrating technology into learning and research (mailto:inquiry@educom.edu Body: Please send EDUCOM info.).

• EDUCOM Gopher: (gopher://educom.edu).

• EDUCOM Web: (http://educom.edu).

• EFF: public interest organization to educate public about computer and communication technologies. (ftp://ftp.eff.org/pub/EFF/about-eff).

• EFF Gopher: A gopher information server for the Electronic Frontier Foundation (gopher://gopher.eff.org).

• EFF Web Page: Electronic Frontier Foundation Web Page (http://www.eff.org).

• EMA: Electronic Messaging Association, foster the development and use of secure global electronic commerce (http://www.ema.org/ema/).

• ENA: Electronic Networking Association, promotes electronic networking (mailto:comserve@vm.its.rpi.edu Body: send Netweave Winter91).

• FARNET: non-profit corporation to advance the use of computer networks for research and education. (ftp://ftp.farnet.org).

• GNA: Globewide Network Academy, a non-profit corporation, affiliated with the Usenet University, goal is create an fully accredited online university (http://uu-gna.mit.edu:8001/uu-gna/index.html).

• HPCC Office: the National Coordination Office (NCO) for High Performance Computing and Communications (gopher://www.hpcc.gov).

• IANA: Internet Assigned Numbers Authority (mailto:iana@isi.edu).

• IATH: Institute for Advanced Technology in the Humanities (http://jefferson.village.virginia.edu).

• ICC: The International Center for Communications, link communities through technology (http://www.ifi.uio.no/∼terjen/cv/ICC/info.html).

• IGC: Institute for Global Communications, improving global network communication and information exchange (ftp://igc.org/README).

• IGC Gopher: (gopher://igc.apc.org).

• IRMAC: Information Resource Management Association of Canada (telnet://freenet.carleton.ca).

• IRVC: Institute for Research on Virtual culture (ftp://byrd.mu.wvnet.edu/pub/estepp/IRVC).

• IRTS: the Internet Roundtable Society (http://www.irsociety.com).

• ISOC: The Internet Society, supports the development of the Internet and promotes education and applications. (ftp://ftp.isoc.org/isoc/).

• ISOC Gopher: the Internet Society’s Information server (gopher://gopher.isoc.org).

• ISOC Web: the Internet Society’s web page (http://info.isoc.org/home.html).

• ISTE: International Society for Technology in Education (mailto:ISTE@Oregon.uoregon.edu).

• Kestrel Inst: a non-profit computer science research institute methods for incremental automation of the software (http://kestrel.edu).

• NIIT: National Information Infrastructure Testbed (http://www.esi.com/niit/niit top.html).

• NPTN: National Public Telecomputing Network (ftp://nptn.org/pub/info.nptn/).

• OSF: Open Software Foundation (http://www.osf.org:8001/index.html).

• PD: Public Domain, Inc., a 501(c)3 non-profit organization whose stated mission is to explore the interface between art, technology, and theory (http://noel.pd.org).

• WAENA: WideArea Educational Network (mailto:jwmilton@waena.portal.com).

• SEA: Society for Electronic Access (http://www.panix.com/sea/).

• UniForum: the largest vendor-independent association for open systems professionals (http://www.uniforum.org).

• USENIX: a forum for innovation and research in UNIX and modern open systems (http://usenix.org).

7.7 Standards

These are organizations that play some part in setting technical standards for telecommunications and networking.

• ACM-TSC: Association for Computing Machinery Technical Standards Committee (http://www.cs.purdue.edu/homes/spaf/acm/acm-tsc.html).

• ANSI: American National Standards Institute, a US clearinghouse for standards, member of ISO (http://www.ansi.org).

• ATIS: Alliance for Telecommunications Industry Solutions (phone:(202)434-8845).

• CCITT: International Telegraph and Telephone Consultative Committee, a member of the ITU (phone:+41-22-995111).

• CCITT Doc Retrieval: (mailto:teledoc@itu.arcom.ch Body: help).

• Committee T1: ANSI (American National Standards Institute) accredited ATIS (Alliance for Telecommunications Industry Solutions) sponsored Committee T1 (Telecommunications); the focal point for developing U.S. positions for the ITU-T (formerly CCITT) (http://www.t1.org).

• DISA: Data Interchange Standards Association, Inc. (http://www.disa.org).

• ECMA: European Computer Manufacturers Association, computer and communications technologies (phone:+41-22-353634).

• EIA: Electronic Industries Association, national trade association (phone:(202)457-4966).

• IEEE: Institute of Electrical and Electronics Engineers (http://www.ieee.org).

• IEEE Standards: (http://stdsbbs.ieee.org:70/0/pub/ieeestds.htm).

• ISO: International Organization for Standardization, a voluntary body of national standardization organizations (phone:+41-22-341240).

• ISO: International Organization for Standardization (http://www.iso.ch).

• ISO info: International Organization for Standardization information, by Dr. Hiroaki Ikeda (http://www.hike.te.chiba-u.ac.jp/ikeda/ISO/home.html).

• ITU: International Telecommunication Union, a United Nations agency which coordinates

telecommunications (http://www.itu.ch). NCS: National Communications System, consortium of US federal agencies with large telecom capabilities

• NIST-USA: National Institute of Standards and Technology WWW Home Page (http://www.nist.gov/welcome.html).

8 BIBLIOGRAPHY 70

8 BIBLIOGRAPHY

This section lists useful lists of books articles, and bibliographies. For the sake of compactness, the book list here is short. For longer lists, see the online or paper bibliographies available. I’ve prepared a list of many books and article readings in CMC and computer networking. That file is the CMC/December bibliography, listed below.

8.1 Online

Online bibliographies and Syllabii

• AT+T: computer networking and related technical topics (http://www.research.att.com/biblio.html).

• Barron’s Bib: a list of links to articles by Billy Barron about Internet information services, library networking, Internet technical, BBSes, and reviews (http://www.utdallas.edu/acc/billy-bib.html).

• Bib List: list of bibliographies related to computing literature (gopher://ukoln.bath.ac.uk:7070/11/Link/Tree/Computing/Literature).

• CMC/December: bibliography of selected readings in CMC and networking, by John December (ftp://ftp.rpi.edu/pub/communications/internet-cmc.bib).

• CMC/Hacker: CMC bibliography, by Kenneth L. Hacker (mailto:comserve@vm.its.rpi.edu Body: get compumed biblio).

• CMC/Parker: (mailto:listserv@cmuvm.csv.cmich.edu Body: send netstart info).

• CMC/ProjectH: a bibliography developed collaboratively by participants in ProjectH (ftp://ftp.arch.su.edu.au/pub/projectH/).

• CMC/Shedletsky: (ftp://tcet.unt.edu/pub/telecomputing-info/cmc-bibliography.txt).

• CMC: communication (mailto:comserve@vm.its.rpi.edu Body: get compunet biblio).

• CMC course/Harrison: graduate CMC theory, Rensselaer Polytechnic Institute, Teresa Harrison, John December, Emilie Gould, Laura Gurak (http://www.rpi.edu/Internet/Guides/decemj/course/cmc.html).

• CMC course/Stephen: description of a CMC course (mailto:comserve@vm.its.rpi.edu Body: get compumed syllabus).

• CMC ED Higgins: (mailto:comserve@vm.its.rpi.edu Body: get rescmc syllabus).

• CMC ED Paulsen: (mailto:comserve@vm.its.rpi.edu Body: get disted2 biblio).

• CMC ED Rice: (mailto:comserve@vm.its.rpi.edu Body: get cmced biblio).

• CMC Study Review: Ed CMC, a field study of recent research (ftp://tcet.unt.edu/pub/telecomputing-info/cmc-study-review.txt).

• Comp Sci: A Collection of Computer Science Bibliographies available on the Internet (ftp://ftp.cs.umanitoba.ca/pub/bibliographies/index.html).

• CSCW Bibliography: (ftp://ftp.cpsc.ucalgary.ca/pub/projects/CSCWbibliography/).

• Ed Tech/England: for a course, Introduction to Educational Telecommunications (mailto:comserve@vm.its.rpi.edu Body: get EdTel biblio).

• Electronic Serials: (ftp://infolib.murdoch.edu.au/pub/bib/bailey.bib).

• Email books: A survey of books about electronic mail (http://www.dsv.su.se/∼jpalme/e-mail-books-survey.html).

• Email 1/Thelen: email in social contexts (A-H) (mailto:comserve@vm.its.rpi.edu Body: get email 1 biblio).

• Email 2/Thelen: email in social contexts (I-Z) (mailto:comserve@vm.its.rpi.edu Body: get email 2 biblio).

• FYI Bibliography: Where to Start - A Bibliography of Internetworking Information (ftp://nic.merit.edu/documents/fyi/fyi 03.txt).

• IAT Internet Bib: Institute for Academic Technology, Internet Bibliography, compiled by Carolyn Kotlas (ftp://gandalf.iat.unc.edu/user/home/anonftp/guides/irg-14.txt).

• Info Nets/Newby: course outline and bibliography for a course ‘Information Networking’ (mailto:comserve@vm.its.rpi.edu Body: get InfoNet Syllabus).

• Information Soc: a special issue of the Computer Underground Digest, bibliography by Rob Kling (ftp://ftp.eff.org/pub/Publications/CuD/CuD/cud5.68).

• Internet Bks Review: reviews of Internet books, from NETTRAIN list (mailto:listserv@ubvm.cc.buffalo.edu Body: GET NETTRAIN REVIEWS NETTRAIN).

• Internet Books Reviews: reviews of Internet books, from NETTRAIN list (ftp://ubvm.cc.buffalo.edu/nettrain/nettrain.reviews).

• Internet Books: Unofficial Internet Book List, by Kevin Savetz (ftp://rtfm.mit.edu/pub/usenet/news.answers/internet-services/book-list).

• Internet Bks Quarterman: (ftp://nic.merit.edu/documents/rfc/rfc1432.txt).

• Internet/NREN Pugh: an annotated bibliography of the Internet and NREN by Pugh (mailto:comserve@vm.its.rpi.edu Body: get internet biblio).

• Intro Internetworking: (ftp://nic.merit.edu/documents/fyi/fyi 19.txt).

• JCMC Cite Site: Journal of Computer-Mediated Communication Citation Site, list of online bibliographies and references (http://shum.huji.ac.il/jcmc/jcmc-citesite.html).

• Network Reading List: (ftp://ftp.utexas.edu/pub/netinfo/reading-list/net-read.txt).

• Networked Info Bib: (ftp://infolib.murdoch.edu.au/pub/bib/stanton.bib).

• New User’s Bib: a short bibliography for new users, from NETTRAIN list (mailto:listserv@ubvm.cc.buffalo.edu Body: GET NEWUSERS FAQ NETTRAIN).

• New User’s Bib (ftp): (ftp://sluaxa.slu.edu/pub/millesjg/newusers.faq).

• Organizational CMC: Bibliography of Organizational Computer-Mediated Communication, from contributors and editors at Journal of CMC (http://shum.cc.huji.ac.il/jcmc/rudybib.html).

• Telecom/Newby: course outline and bibliography for a course entitled ‘Telecommunications’ (mailto:comserve@vm.its.rpi.edu Body: get Telecomm Syllabus).

• WAIS Bibliography: (ftp://infolib.murdoch.edu.au/pub/bib/lincoln.bib).

• WWW Bib: Bibliography for the World Wide Web (http://www.w3.org/hypertext/WWW/Bibliography.html).

8.2 Special

Special issues/journals on networks/CMC

• Communications of the ACM (1991). Collaborative computing issue. Volume 34, number 12, December.

• Communication Research (1987). Special issue on Theories of Organizations and Information Technology. Volume 14, number 5, October.

• Communication Yearbook (1989). Issues and Concepts in Research on Computer-Mediated Communication Systems. Volume 12, chapter 8.

• Educom Review (1994). Special Issue on Networking. Volume 29, number 2, March/April.

• EJC/REC V3N2: Electronic Journal of Communication special issue on CMC (mailto:comserve@vm.its.rpi.edu Body: send EJCTOC V3N293).

• Internet Business Journal (1993). Special Issue on Internet Advertising. Volume 1, numbers 3/4, September/October 1993.

• Internet Society News (1994). Special Issue: A Year in the Life of the Internet. Volume 2, number 4, winter.

• Journal of Communication (1989). Issue on computer communication affecting social power distribution. Volume 39, number 3.

• Online Access (1993). Issue on the Internet. October.

• The Serials Librarian (1991). Issue on technology and information for serials. Volume 21, numbers 2/3.

• Serials Review (1992). Issue on Economic Models for Networked Information. Volume 18, numbers 1-2.

• Scientific American (1991). Issue on computer networks. Volume 265, number 3, September 1991.

• The South Atlantic Quarterly (1993). Flame Wars: The Discourse of Cyberculture. Volume 92, number 4, Fall 1993.

• Wall Street Journal (1993). A Report on Technology. Volume CCXXII, number 96, November 15, 1993. Section R.

• Written Communication (1991). Issue on computers, language, and writing. Volume 8, number 1, January.

8.3 Books

A selected listing of particularly useful paper books.

8.3.1 Bibliographies

• The American Center for the Study of Distance Education (1992). Computer-mediated communication for distance education. University Park, PA: The Pennsylvania State University.

• Burge, Elizabeth (1992). Computer Mediated Communication and Education: A Selected Bibliography. Toronto, Ontario: Distance Learning Office, Ontario Institute for Studies in Education, 252 Bloor St W, Toronto, Ontario M5S 1V6.

• Romiszowski, A. J. (1992). Computer-mediated communication: a selected bibliography. Englewood Cliffs, NJ: Educational Technology Publications.

8.3.2 Internet

There are many books about the Internet. See the Savetz list for a comprehensive list of books available. —— Lists of books

• Internet Books-Yahoo: lists Internet books with Web support sites (http://www.yahoo.com/Business/Products and Services/Books/Internet/).

• Internet/Savetz: Unofficial Internet Book List, by Kevin Savetz (ftp://rtfm.mit.edu/pub/usenet/news.answers/internet-services/book-list).

• Internet business books: (http://arganet.tenagra.com/Tenagra/books.html). —— Individual books

• Bang, Steve, Martin Moore, Rick Gates, et al. (1994). The Internet Unleashed. Indianapolis, IN: Sams Publishing.

• Carroll, Jim and Rick Broadhead. (1994). The Canadian Internet Handbook. Canada: Prentice-Hall.

• Canadian/Internet: support site for Canadian Internet Handbook, by Rick Broadhead and Jim Carroll (http://www.csi.nb.ca/handbook/handbook.html).

• Cronin, Mary J. (1993). Doing business on the Internet: how the electronic highway is transforming American companies. New York: Van Nostrand Reinhold.

• Braun, Eric. (1994). The Internet Directory. New York: Fawcett Columbine.

• December, John, Neil Randall, and others. (1994). The World Wide Web Unleashed. Indianapolis, IN: Sams Publishing.

• December/WWWU support: support site for the book The World Wide Web Unleashed, by John December and Neil Randall (http://www.rpi.edu/∼decemj/works/wwwu.html).

• De Presono, Odd and Mike Weaver. (1994). The Online World. Ontario: Productive Publications.

• Dern, Daniel P. (1994). The Internet Guide for New Users. New York: McGraw-Hill.

• Ellsworth, Jill H. and Matthew V. Ellsworth. (1994). . The Internet Business Book. Wiley.

• Ellsworth/Business: The Internet Business Book (ftp://ftp.std.com/pub/ibb/ibb.html).

• Estrada, Susan (1993). Connecting to the Internet: An O’Reilly Buyer’s Guide. Sebastopol, CA: O’Reilly and Associates.

• Falk, Bennett. (1994). The Internet Roadmap. San Francisco: Sybex.

• Fisher, Sharon (1993). Riding the Internet Highway. Carmel, IN: New Riders.

• Glister, Paul A. (1993). The Internet navigator: a new user’s guide to network exploration. New York: Wiley.

• Glossbrenner, Alfred and Emily. (1994). Internet Slick Tricks. New York: Random House Electronic Publishing.

• Glossbrenner, Alfred. (1995). Internet 101. New York: Windcrest/McGraw-Hill.

• Hahn, Harley and Rick Stout (1994). The Internet Complete Reference. New York: Osborne McGraw-Hill.

• Hahn, Harley and Rick Stout (1994). The Internet Yellow Pages. New York: Osborne McGraw-Hill.

• Hahn, Harley (1993). A Student’s Guide to Unix. New York: McGraw-Hill.

• Hardie, Edward T. L. and Vivian Neou (1993). Internet: mailing lists. Englewood Cliffs, NJ: Prentice Hall.

• Kehoe, Brendan (1994). Zen and the Art of the Internet: A Beginner’s Guide. 3rd ed. Englewood Cliffs, NJ: Prentice Hall.

• Krol, Ed (1994). The Whole Internet User’s Guide and Catalog. 2nd ed. Sebastopol, CA: O’Reilly and Associates.

• Lambert, Steve and Howe, Walt. (1994). Internet Basics: Your Online Access to the Global Electronic Superhighway. New York: Random House.

• Lane, Elizabeth S. and Craig Summerhill (1992). An Internet primer for information professionals: a basic guide to Internet networking technology. Westport, CT: Meckler.

• LaQuey, Tracy with Jeanne C. Ryer. Foreword by Al Gore. (1992). The Internet companion: a beginner’s guide to global networking. Reading, MA: Addison-Wesley.

• Lynch, Daniel C. and Marshall T. Rose., Eds. (1993). Internet System Handbook. Reading, MA: Addison-Wesley.

• Malamud, Carl. (1993). Exploring the Internet: a Technical Travelogue. Englewood Cliffs, NJ: Prentice Hall.

• Randall, Neil. (1994). Teach Yourself the Internet: Around the World in 21 Days. Indianapolis, IN: Sams Publishing.

• Randall/TYI support: Teach Yourself the Internet by Neil Randall, support site (http://watarts.uwaterloo.ca/ENGL/nrandall/tyi.html).

• Rose, Marshall T. (1993). The Internet Message: Closing the Book with Electronic Mail. Englewood Cliffs, NJ: PTR Prentice Hall.

• Savetz, Kevin. (1994). Your Internet Consultant–the FAQs of Life Online. Indianapolis: Sams Publishing.

• Smith, Richard and Mark Gibbs (1993). Navigating the Internet. Indianapolis: Sams Publishing.

• Strangelove, Michael and Aneurin Bosley. (1994). How to Advertise on the Internet. Strangelove Press.

• Tennant, Roy, John Ober, and Anne G. Lipow (1993). Crossing the Internet Threshold: An Instructional Handbook. Berkeley, CA: Library Solutions Press.

• Thomas, Brian J. (1995). The Internet for Scientists and Engineers. www@spie.org: The International Society for Optical Engineering.

• Thomas/scientists: The Internet for Scientists and Engineers book information (http://www.spie.org/book.html).

• Tolhurst, William A., Mary Ann Pike and Keith A. Blanton (1994). Using the Internet. Indianapolis, IN: Que.

• Quarterman, John (1994). The Internet connection: system connectivity and configuration. Reading, MA: Addison-Wesley.

8.3.3 Networks

• Aboba, Bernard. (1994). The Online User’s Encyclopedia. Reading, MA: Addison-Wesley.

• Adams, Rick and Frey, Donnalyn. (1993). !%@:: A Directory of Mail Addressing and Networks, 3rd Ed. Sebastopol, CA: O’Reilly and Associates.

• Black, Uyless. (1993). Computer Networks: protocols, standards, and interfaces, 2nd end. Englewood Cliffs, NJ: PTR Prentice-Hall.

• Kessler, Gary C. (1990). ISDN: concepts, facilities, and services. New York: McGraw-Hill.

• LaQuey, Tracy L., ed. (1990). The User’s Directory of Computer Networks. Bedford, MA: Digital Press.

• Li, Xia and Nancy B. Crane. (1993). Electronic style: a guide to citing electronic information. Westport, CT: Meckler.

• McClure, Charles R., Ann P. Bishop, Philip Doty, and Howard Rosenbaum. (1991). The National Research and Education Network (NREN): Research and Policy Perspectives. Norwood, NJ: Ablex.

• Motorola Codex. (1992). The Basics Book of Information Networking. Reading, MA: Addison-Wesley.

• Quarterman, John S. (1990). The Matrix: Computer Networks and Conferencing Systems Worldwide. Bedford, MA: Digital Press.

• Staple, Gregory C., Ed. (1993+). TeleGeography. tgi@cais.com: TeleGeography, Inc.

8.3.4 CMC

• Berge, Zane L. and Mauri P. Collins. (1994). Computer-Mediated Communication and the Online Classroom. 3 volumes. Cresskill, NJ: Hampton Press.

• Chesebro, James W. and Donald G. Bonsall. (1989). Computer-mediated communication: human relationships in a computerized world. Tuscaloosa: University of Alabama Press.

• Dunlop, Charles and Rob Kling, eds. (1991). Computerization and Controversy: Value Conflicts and Social Choices. Academic Press.

• Dutton, William H, Jay G. Blumler, Kenneth L. Kraaemer, Eds. (1987). Wired Cities: Shaping the Future of Communications. Boston, MA: G.K. Hall and Company.

• Ess, Chares, ed. (1995). CyberThoughts: Philosophical Perspectives on Computer-Mediated Communication. Albany, NY: State University of New York Press.

• Gattiker, U. E., ed. (1992). Technology-mediated communication. Berlin; New York: W. de Gruyter.

• Gilder, George. (1994). Telecosm. New York: Simon and Schuster.

• Harasim, Linda M. (1993). Global networks: computers and international communication. Cambridge, MA: MIT Press.

• Harasim, L. M., Hiltz, S. R., Teles, L., and Turoff. M. (1994). Learning networks: a field guide. Cambridge, MA: MIT Press.

• Hiltz, Starr Roxanne and Murray Turoff. (1993). The Networked Nation: Human Communication via Computer. Rev. Ed. Cambridge, MA: MIT Press.

• Jones, Steven G., ed. (1995). Cybersociety: Computer-Mediated Communication and Community. Thousand Oaks: Sage.

• Lea, Martin, ed. (1992). Contexts of computer-mediated communication. New York: Harvester Wheatsheaf.

• Mason, Robin and Anthony Kaye., eds. (1989). Mindweave: Communication, Computers, and Distance Education. New York: Pergamon Press.

• Rapaport, Matthew. (1991). Computer mediated communications : bulletin boards, computer conferencing, electronic mail, and information retrieval. New York: Wiley.

• Rheingold, Howard. (1993). The virtual community: homesteading on the electronic frontier. Reading, MA: Addison-Wesley.

• Rheingold/VC book web: the full text of the book ’The Virtual Community’ by Howard Rheingold (http://www.well.com/www/hlr/vcbook/).

• Shea, Virginia (1994). Netiquette. Info@albion.com: Albion Books.

• Sproull, Lee and Sara Kiesler. (1991). Connections: New Ways of Working in the Networked Organization. MIT Press.